### v1

# Towards User-Focused Research in Training Data Attribution for Human-Centered Explainable AI

Abstract.
While Explainable AI (XAI) aims to make AI understandable and useful to humans, it has been criticised for relying too much on formalism and solutionism, focusing more on mathematical soundness than user needs. We propose an alternative to this bottom-up approach inspired by design thinking: the XAI research community should adopt a top-down, user-focused perspective to ensure user relevance. We illustrate this with a relatively young subfield of XAI, Training Data Attribution (TDA). With the surge in TDA research and growing competition, the field risks repeating the same patterns of solutionism. We conducted a needfinding study with a diverse group of AI practitioners to identify potential user needs related to TDA. Through interviews (N=10) and a systematic survey (N=31), we uncovered new TDA tasks that are currently largely overlooked. We invite the TDA and XAI communities to consider these novel tasks and improve the user relevance of their research outcomes.
1. Introduction
Understanding AI model behaviour is critical for using AI models in practice, especially in high-stakes domains such as medicine, law, or finance (Parliament, 2023). End-users require the understanding of model behaviours to make informed decisions, while developers need this to build reliable models with limited biases (Dwivedi et al., 2023). Explainable AI (XAI) aims to provide insights into AI systems’ functioning for humans. It has produced a variety of explanation methods that address different model architectures and data modalities (Guidotti et al., 2018; Dwivedi et al., 2023).
Despite the progress, the XAI community has faced criticism for its predominant techno-centric focus on solutions as opposed to the problem of helping users understand model behaviour (Ehsan et al., 2021). Historically, XAI research progressed bottom-up: starting with methods providing qualitative observations, then structuring the field with quantitative evaluations, and eventually considering application in real-world environments including users (see Figure 1). Consequently, XAI was mainly driven by “[solutionism (seeking technical solutions) and] formalism (seeking abstract, mathematical solutions)” (Ehsan et al., 2021, p.2). An example of this bottom-up approach is the development of the subfield of feature attribution. Feature attribution explanations quantify each input feature’s contribution to the model prediction, highlighting key inputs. One of the first feature attribution methods in the context of deep learning was presented in 2013 introducing the visualisation of neural network input gradients as saliency map explanations and providing 15 qualitative examples as evidence that saliency maps explain (Simonyan et al., 2013). Several other methods followed, either building on this work and proposing changes to satisfy specific formal axioms (e.g., (Smilkov et al., 2017; Sundararajan et al., 2017)), or proposing model-agnostic methods in a quest for a one-fits-all solution (e.g., (Lundberg and Lee, 2017; Ribeiro et al., 2016)). Around 2017, XAI reached an inflection point where researchers started to re-assess the practicality of explanation methods and proposed standard evaluation protocols to measure how well explanation methods achieve their goal of explaining (Doshi-Velez and Kim, 2017; Adebayo et al., 2018; Bilodeau et al., 2024). Subsequently, recommendations from the social science view of explanations emphasise that explanations are part of a social process and shifted the focus of the field in 2019 (Miller, 2019). XAI was criticised for its strong solutionism with an increasing amount of works studying XAI as a sociotechnical system focusing on the user (Wang et al., 2019; Ehsan et al., 2021; Brennen, 2020; Rutjes et al., 2019; Kim et al., 2022, 2023; Fok and Weld, 2023; Kaur et al., 2020; Mei et al., 2023; Ehsan et al., 2024; Lakkaraju et al., 2022) and naming this subfield human-centred XAI (HCXAI) in 2020 (Ehsan and Riedl, 2020).
Schematic showing the bottom-up and the proposed top-down research approaches. There are three boxes vertically stacked. The bottom box reads “Methods: Studies on new algorithms. XAI example: Saliency maps (Simonyan et al., 2013)”. The middle box reads “Evaluations: Studies on protocols for comparing methods. XAI example: Sanity checks for saliency maps (Adebayo et al., 2017)”. The top box reads “Applications: Studies on methods in user-facing contexts. XAI example: Human-centred explainable AI (Ehsan and Riedl, 2020)”. There is an arrow pointing upwards to the left of the boxes, annotated with “Bottom-Up Approach -¿ Driven by techno-centric solutions” and an arrow pointing downwards to the right of the boxes, annotated with “Proposed Top-Down Approach -¿ Driven by user needs”.
Despite the efforts to rectify the solutionism and formalism in XAI communities, we observe repeating patterns in sub-fields of XAI. For example, such patterns are emerging in the relatively young field of training data attribution explanations (TDA), where the model behaviour is explained based on the training data (Hammoudeh and Lowd, 2024). TDA was introduced to understand deep learning models in 2017 (Koh and Liang, 2017) and has produced several technical approaches to provide TDA explanations since then (Charpiat et al., 2019; Guo et al., 2021; Pruthi et al., 2020; Schioppa et al., 2022; Park et al., 2023; Bae et al., 2024). This development demonstrates a strong focus on seeking technical solutions in the TDA field similar to how feature attribution research started. Currently, TDA research is gaining momentum due to studies showing that TDA can be applied to address issues like data valuation, debiasing, memorisation and copyright in foundational models (Feldman and Zhang, 2020; Grosse et al., 2023; Choe et al., 2024; Jain et al., 2024; Zheng et al., 2023; Kwon et al., 2023; Lin et al., 2024; Wang et al., 2024b). Yet, the current TDA landscapes shows solutionism and formalism, with few user-focused studies, leading to a possible rift between TDA methods and user needs.
In this work, we propose future research directions for the TDA community that are entirely sourced and motivated by the potential users of TDA. We strongly believe that the actual use cases and user needs should inform and guide XAI research, echoing the call for HCXAI (Ehsan and Riedl, 2020). We break down our research question into three parts:
RQ1: What do users need TDA explanations for?
RQ2: What do users need from TDA explanations?
RQ3: To what extent do existing TDA approaches align with user needs?
We conduct a two-stage needfinding study with potential users of TDA technology to answer our research questions. Through a qualitative interview study with AI practitioners (N=10), we identify model developers as a key user group for TDA explanations as opposed to model users. We derive a design space for TDA explanations representing different TDA tasks (e.g., identifying the training samples with the largest effect on model behaviour when removed in contrast to when given a larger weight during training). We study which tasks are needed in model development in a subsequent mixed-methods survey study (N=31) with model developers. By considering the user and their usage scenarios, we identify tasks relevant to users but largely overlooked by current TDA research. We invite the research community to address these gaps and shift the focus of TDA research from solutionism to user-serving questions, i.e., a top-down approach. We advocate for adopting this approach in other areas of XAI and AI as a whole to avoid focusing on technically elegant solutions that may overlook user needs. Our contributions are:
We propose a top-down approach to XAI research driven by user needs and demonstrate this approach on the emerging TDA subfield.
Through a needfinding study, we identify potential use cases, where model developers emerged as key users of TDA. We find that they require TDA explanations to be (1) flexible and adapt to the user’s intuitions in debugging and (2) reliable and stable. Generally, we observe a preference for training data group attribution over individual attribution.
As a result of studying the user needs, we identify tasks and research directions for user-focused TDA that are currently understudied.
2. Related Work
Our work proposes a top-down approach to XAI research, grounded in user needs, which positions it in human-centered explainable AI (HCXAI). We apply this approach to the relatively young subfield of training data attribution (TDA) to provide early insights into users and their needs. To this end, we utilise methods from needfinding. In the following, we relate our work to HCXAI, TDA and needfinding.
2.1. Human-centred explainable AI
The aim of explainable AI (XAI) research is to study methods that explain the behaviour of AI models (Guidotti et al., 2018). The field has developed various explanation methods over the years to explain different types of tasks and models. Yet, explanations are only meaningful in front of an audience. This human aspect was often neglected in XAI studies (Rong et al., 2023; Nauta et al., 2023). To overcome this gap, Human-centred XAI (HCXAI) emphasises the human in the explanation process and focuses on user perspectives in XAI (Ehsan and Riedl, 2020). In HCXAI, explanations of black-box models are studied holistically as sociotechnical systems centred around the user (Ehsan et al., 2024). HCXAI explores how to design AI technology and explanations that focus on user understanding and practicality (Ehsan and Riedl, 2020). Work in this area covers multiple aspects that affect the user, e.g. human-AI-collaboration (Kim et al., 2023; Shen et al., 2023; Taesiri et al., 2022; Lakkaraju et al., 2022), design guidelines (Liao et al., 2020; Hadash et al., 2022; Wang et al., 2019), evaluation of human understanding of explanations (Rong et al., 2023; Kim et al., 2022; Hadash et al., 2022; Rutjes et al., 2019) and understanding user needs and requirements (Juneja et al., 2024; Kaur et al., 2020; Brennen, 2020; Jin et al., 2023). Our work contributes to the HCXAI field by proposing an early consideration of user needs to inform ongoing research in XAI methods. We focus particularly on TDA explanations which have not been explored widely in HCXAI before.
2.2. Training data attribution
Training data attribution (TDA) explains model predictions by pointing to training data samples relevant to the model predictions (Hammoudeh and Lowd, 2024). TDA differs from the mainstream XAI approach known as feature attribution (FA) where the model predictions are attributed to the features of an input given to the model, disregarding the impact of training data. Interest in TDA has recently accelerated with the paradigm shift towards data-centric AI (DCAI) (Jin et al., 2024), where the importance of using the right training data is emphasised over other factors like model architecture and optimisation algorithms. Under the context of DCAI, TDA has been utilised to enhance data quality by cleaning faulty data labels (Teso et al., 2021) and detecting model biases (Pezeshkpour et al., 2022; Brunet et al., 2019; Wang et al., 2024a; Jain et al., 2024). The recent rise of foundational models, such as large language models (LLMs) and diffusion-based image-generation models, further contributed to the interest in TDA, as it has proven effective for studying the inner mechanisms as well as understanding the associated copyright and privacy leakage issues therein (Feldman and Zhang, 2020; Choe et al., 2024; Zheng et al., 2023; Grosse et al., 2023; Lin et al., 2024). With the accelerated growth, we observe again a general focus on technical and mathematical solutions in the TDA community (Koh and Liang, 2017; Charpiat et al., 2019; Guo et al., 2021; Schioppa et al., 2022; Pruthi et al., 2020; Park et al., 2023; Bae et al., 2024; Choe et al., 2024; Kwon et al., 2023), without a deep understanding of user needs, even in studies applying TDA to a downstream tasks (Feldman and Zhang, 2020; Jain et al., 2024; Zheng et al., 2023; Akyürek et al., 2022). Our work addresses this gap by studying practical needs and providing concrete recommendations for the field.
2.3. Needfinding
Needfinding is an exploratory and qualitative approach to user research to understand users’ needs beyond what simply asking users may tell (Patnaik and Becker, 1999). It aims at studying the true user needs, regardless of solutions, which may not be apparent to the users themselves. Needfinding stems from the field of design thinking (von Thienen et al., 2019), which is an iterative design approach driven by user needs. Needfinding is an iterative process involving users where the subsequent studies extract greater information about their problems and needs through field studies, interviews, and lab studies. In XAI, needfinding has been used to study general needs such as user expectations for XAI (Brennen, 2020) or end-user transparency needs in AI decision-support systems (Turri et al., 2024; Lin et al., 2024) by using semi-structured interviews, scenario-based study designs and qualitative analyses of case-study related text files. Needfinding has also been applied to more narrow questions, like extracting user needs for web-search explanations through surveys and semi-structured interviews (Juneja et al., 2024) or understanding explainability needs in human-AI collaborations (Kim et al., 2023). Our work extends this overall research direction by studying the user needs of TDA explanations through a two-stage study involving semi-structured interviews and a systematic survey study.
3. Training data attribution
Our work studies user needs for training data attribution (TDA) explanations. This section provides background on TDA. TDA is links specific model behaviour to the training data, treating the data as the root cause for learned behaviours (Hammoudeh and Lowd, 2024). By providing training data relevant to what the model has learnt, TDA gives insights into the model. TDA has been applied to enhancing data quality and reducing model biases, as well as answering questions around data valuation, memorisation and copyright issues in foundational models in the recent years (Choe et al., 2024; Zheng et al., 2023; Grosse et al., 2023; Lin et al., 2024), attracting unprecedented attention.
Formal definition.
The TDA community has focused on a particular instance of attributing a model prediction on a single test sample on a single training sample. More formally, let us consider a training sample and a test sample , where indicates the input and indicates the true answer the model is supposed to predict. We define the attribution score for and as the change in the correctness of the model prediction on , measured via the model loss , before and after removing the sample from the training procedure (Hampel, 1974):
Here, we indicate the model trained with all training samples as and the one trained without as .
Main focus of TDA research.
Computing directly by re-training a model without each training sample is computationally prohibitive. To identify the most influential training sample, the re-training has to be repeated as many times as the number of training samples. The main focus of the TDA community has been on efficient approximations of Equation 1. In 2017, Koh & Liang (Koh and Liang, 2017) introduced a gradient-based approximation called influence functions (IF). Since IF was still computationally prohibitive due to the need to compute and invert a massive Hessian matrix, subsequent works in 2020 and 2021 have focused on speeding up the algorithm by trading off precision and computational costs (Pruthi et al., 2020; Charpiat et al., 2019; Schioppa et al., 2022; Guo et al., 2021). After the advent of ChatGPT in late 2022 (OpenAI, 2022) and the boom in large language models in 2023 (Naveed et al., 2023), the community has developed further enhancements to apply the TDA methods on billion-scale LLMs (Bae et al., 2024; Choe et al., 2024; Grosse et al., 2023).
Critique.
We observe hints of solutionism and formalism in the development of TDA research, as the majority of the community effort is dedicated to addressing computational inefficiencies of the method itself, without questioning the practical relevance and user values of TDA as defined in Equation 1. This particular formulation of TDA has not gone without criticism within the community: a prior work (Nguyen et al., 2023) discussed the statistical insignificance of removing a single training sample in modern neural network training and (Ilyas et al., 2022) recognised the limitation and argued for examining the removal of multiple training samples. Some TDA works have explored possible applications that are potentially interesting to users such as fixing mislabelled data (Koh and Liang, 2017) or identifying brittle classes (Ilyas et al., 2022). However, these applications are still not driven by user needs. There is no known work in the TDA domain involving actual users to identify their needs to assess the relevance of the current TDA tasks and identify research directions grounded in user needs that the community should focus on.
4. Understanding user needs of TDA explanations through needfinding study
Our work studies user needs of training data attribution (TDA) explanations to facilitate user need-driven TDA research. To this end, we conduct a two-stage needfinding study. First, we identify use cases for TDA explanations using semi-structured interviews. The experimental setup is described in § 4.1. § 4.2 presents the interview findings. We derive a design space representing different types of information that may be needed from TDA in § 4.3. Second, we conduct a systematic survey study targeted at extracting user needs from this space (§ 4.4). IRB approval was obtained from one of the authors’ institutions.
4.1. Interview study
In the interview study, we conduct semi-structured interviews with users of machine learning (ML) applications outside of ML research. This serves as an exploratory study to gain an impression of potential usage scenarios and user needs for TDA RQ1 – What do users need TDA explanations for?
4.1.1. Participants
To get a realistic impression of user needs for TDA explanations in practice, we target participants who work with ML applications outside of academia. Our inclusion criteria are: Participants should (1) have at least one month of experience in working with ML systems and (2) work in a high-risk application area according to the EU AI Act (Parliament, 2023) (e.g., health care, law enforcement, complete list in Appendix A). This criterion serves to find users who are likely to use explanations, as these areas are subject to further regulations (Guidotti et al., 2018; Doshi-Velez and Kim, 2017). Recruiting participants poses a challenge, especially in high-risk application areas. Hence, we use purposive sampling (Guest et al., 2023) and approach participants from the authors’ network. We recruit 10 participants from various domains and degrees of experience (see Table 1).
| ID | Location | Domain | Type | Job experience / with ML | Type of ML |
| P1 | DE | HR | User | 3 years / 1 months | Chatbot |
| P2 | US | AV | Developer | 2 years / 7 years | Prediction model |
| P3 | NL | TC | Developer | 3 years / 5 years | Prediction model |
| P4 | FI | CV | Developer | 4 years / 6 years | Prediction model |
| P6 | CH | Health | User | 2 years / 2 years | Prediction model |
| P7 | NL | Health | Developer | 1 year / 3 years | Prediction model |
| P8 | BE | Health | Developer | 2 years / 6 years | Prediction model |
| P9 | PK | Health | Developer | 5 years / 2 years | Prediction model |
| P10 | DE | HR | User | 3 years / 1 year | Chatbot |
| P11 | DE | Health | User | 10 years / 6 years | Clustering, Chatbot |
4.1.2. Interview process
The interviews were conducted during June – September 2023, either in person or remotely through video call using Zoom 111https://zoom.us/. All interviews are one-on-one conversations in English, except for P10 in German. Participants were first briefed about the purpose of the study and data processing. Upon receiving informed consent, we began the interview recording. In general, the interviews lasted between 30 and 60 minutes. Each interview addresses the following topics to answer RQ1 - What do users need TDA explanations for?:
Job-related information. Perspectives may vary between different domains, levels of seniority and experience with the ML tool. This information provides user context and will help in differentiating user groups.
Interviewee’s workflow with ML systems. By asking about the workflow with the ML tool, we wish to understand the patterns of usage and challenges participants encounter. We collect this information to identify challenges that could potentially addressed by TDA.
Perspectives on training data. Since we investigate TDA explanations, we explicitly ask participants about the role training data plays in their tasks. This indicates what type of data-centric information is relevant to the participant’s work and for which tasks.
Perspectives on XAI and TDA. We address the participant’s perspectives on XAI and particularly on TDA to understand current explanation usage scenarios and explore participant’s opinions about TDA, a data-centric approach to XAI.
4.1.3. Data analysis
We analyse the interview data and extract relevant information for answering RQ1. The interviews are transcribed automatically using Whisper (Radford et al., 2023) and cleaned up manually by the authors. The transcript is then pseudonymised. We translated P10’s German transcript to English using DeepL 222https://www.deepl.com/translator. We then analyse the transcripts through an inductive thematic analysis (Braun and Clarke, 2012) by two coders, where the codes are directly annotated in the transcript. The analysis is iterative: The interview transcript of P1 is first jointly analysed in an initial coding workshop that resulted in an initial set of themes. Afterwards, the coders independently code five transcripts, expanding on the themes found in the initial analysis. During an intermediate coding workshop, agreements and disagreements between the coder’s themes are discussed. The workshop resulted in a merged definition of the themes that are used for the remaining transcripts. At the intermediate coding workshop, the interrater agreement is 77.3% measured by the percentage of agreement participants coded to themes. The final coding workshop takes place after both coders reviewed the remaining transcripts. The final interrater agreement is 80.3%.
4.2. Interview findings
The figure depicts the resulting themes from the coding analysis of the interview transcripts. There are 28 themes in total, spread over six theme areas. The theme areas are larger boxes with rounded edges, encompassing the themes denoted in square boxes (intended to mimic sticky notes). Themes related to training data and TDA are marked with a star. The themes areas with themes in parentheses are: Role of ML systems (Work assistant, decision support, automation), workflow with ML systems (Collab with domain experts, standard model choices, data-focused debugging process*, data-quality checks*), Role of training data (Most important variable in a model*, interesting but no capacity*, not interesting*, increasing training data as a bug fix*, training data artifacts as main bug causes*), challenges in working with ML systems (trust calibration, resource constraints, misuse due to lacking know-how, validation and evaluation*, data quality issues*, stochastic nature of ML models*), use of feature attribution XAI in practice (XAI is not used, XAI is little used, understand real-world phenomena, get customer buy-in, sanity check for model reasoning, model development*), user perspectives on data-centric XAI (usefulness of XAI is use-case dependent*, XAI is useful*, TDA complementary to feature attribution*, different notions of interesting training samples*).
The interview study aims to provide insights into the potential usefulness of TDA, answering RQ1 - What do users need TDA explanations for? The analysis yields six theme areas, which are common across all participants (see light grey boxes in Figure 2): the role of the ML system as an application, the workflow of participants with ML systems, the role of training data for participant’s work, the main challenges in working with ML systems, the use and role of explanation tools, and perspectives and opinions about data-centric explanations like TDA. We elaborate on each area below and provide a full overview in Appendix B:
4.2.1. Role of ML systems.
Our participants work with different types of ML models across several domains and in different positions (see Table 1). We find that ML systems are used as work assistants (P1,P10,P11), decision support systems (P3,5,6,7,8,9), and for automation in systems like autonomous vehicles or machinery (P2, P4). These roles provide context for our analysis, as the needs and usage scenarios of XAI may differ based on the level of user interaction. For example, P4 develops automated computer vision models for animal population monitoring with minimal human involvement, while P10 uses a chatbot to handle smaller tasks like answering employee questions. Our analysis is set against the backdrop of these use cases.
4.2.2. Workflow with ML systems.
As ML systems take different roles, the workflows differ. We notice that ML systems are not an essential part of the workflow for system users, as opposed to developers. System users only integrate the ML system in their work when it delivers helpful suggestions and report bugs or ignore it otherwise (P6, P10). We recognise a separation of domain knowledge (collaboration with domain experts): “Because personally, I cannot know if the model is doing the correct thing […] business have to tell me” (P3). For developers whose job revolves around the ML system, we identify themes that are relevant for understanding usage scenarios of XAI and TDA. We find that the development workflow is heavily centred around data: “[What] drives your model is your data. […]” (P8) (data-focused debugging process, data quality checks). Often, standard model architectures are employed with standard hyperparameters (standard model choices) and the main work lies in data curation (P2, P4, P3, P7, P8, P9). This theme area shows that there are two distinct user groups, system users who utilise ML systems and model developers who build ML systems. They have different workflows and potential XAI use cases.
4.2.3. Role of training data.
The training data plays different roles to system users and model developers. While it is the most important variable in a model for developers (P3, P4, P8), training data is generally not interesting to system users (P1, P10). System users often do not have capacity to interact much with the ML system to verify model behaviour, even if they an interest in the underlying mechanisms exists, because they are preoccupied with their own tasks (interesting but no capacity) (P6). However, for developers, training data is a lever for model performance (increasing training data as a bug fix, training data artifacts as main bug causes) (P2, P4, P3, P7, P8, P9, P11). These insights underline that XAI needs to be intuitive, and TDA is particularly suited for model development and less for system users.
4.2.4. Challenges in working with ML systems.
In the interviews, we ask participants about challenges in their ML workflow to explore if XAI can address them. We identified six themes in this theme area. We confirm Kim et al.’s (Kim et al., 2023) observation that system users’ main challenge for working with ML systems is trust calibration (P1, P6, P8, P11). In addition, we find that system users fear they may misuse the ML system due to lacking know-how (P1, P10, P11), highlighting the need for user training to make effective use of ML systems in practice. For model developers, we identify several challenges: Data quality issues are often the root cause of model malfunction (data quality issues) (P2, P4, P3, P7, P8, P9, P11), in turn impacting model validation and evaluation. For instance, participants encounter difficulties due to missing data or absent labels (P2, P4, P3, P7, P8, P9). Additionally, resource constraints (P2, P4, P11) and the stochastic nature of ML models (P2) are pain points for the development and evaluation of ML systems. Our analysis shows that the challenges faced by system users require sociotechnical consideration, whereas challenges faced by model developers are more technical with data quality and evaluation being most pressing. We believe that the latter is a suitable usage scenario for TDA.
4.2.5. Use of XAI in practice.
This theme area groups six themes about the use or lack of XAI in practice. Our analysis reveals that XAI is usually not a standard part of an ML system in practice (XAI is not used, XAI is little used) (P1, P4, P6, P7, P9, P10, P11). In fact, the uses of XAI, in particular feature attribution, among our participants are limited to model development (P2, P3, P8). XAI offers explanations for per-example debugging or act as a sanity check for model reasoning (P8, P9). Unexpectedly, developer participants also use XAI to understand real-world phenomena modelled by their ML system (P3) and as a communication means to get customer buy-in for their ML systems (P8). While explanations have different purposes, we note that participants use XAI tools mainly as an out-of-the-box functionality (e.g., the SHAP library (Lundberg and Lee, 2017)). We find that implementation thresholds must be low for the adoption of XAI in practice.
4.2.6. User perspectives on TDA
None of the participants were familiar with TDA, highlighting a gap between research and practice. Our analysis, based on discussions about the concept of TDA, identifies four themes reflecting participants’ views on XAI and TDA. Overall, participants find XAI useful for debugging and communication, though they noted its usefulness is use-case dependent and not always guaranteed. They expect TDA not to be an exception (P3, P6, P8, P11), but are optimistic about its potential, especially for model development. For instance, P2 mentioned that TDA could save time by identifying faulty training data. One participant suggested combining TDA with feature attribution for debugging (P3). We observed that participants have different notions of interesting training samples, with some focused on finding faulty data (P2, P11) and others on identifying samples similar to a specific one (P6, P7, P8).
In conclusion, we identify that TDA could be valuable for model debugging, leading us to focus on this user group and use case in the remainder of this work.
4.3. Defining a design space of TDA explanations
| Theme A: Data quality issues | ||||
| Theme B: Different notions of interesting training samples | ||||
| ID | Theme | Code | Description | Participants |
| #1 | A | Distribution shifts | If the difference in data distribution between training and testing data is too large, the model does not learn. This can also happen over time. | P3, P4 |
| #2 | A | Data unavailability | Missing data entries or missing labels are common problems, particularly for tabular data. | P2, P3, P7, P8, P9 |
| #3 | A & B | Wrong labels | Annotation is an expensive task subject to human error. mislabelled data is hence a possible reason for model error. | P2, P4 |
| #4 | A | Data scaling | Wrong data scaling in tabular data (e.g. using meters instead of centimetres) can lead models to learn wrong correlations. | P3 |
| #5 | A | Historical data | Outdated data that is not cleaned out of the training dataset can contaminate the training. | P9 |
| #6 | B | Outlier data points | Data outliers, i.e. mislabelled data or “unusual data points” (P3) can aid for model debugging. | P3, P8 |
| #7 | B | Decision boundary samples | Samples at decision boundary are interesting to understand the model behaviour. | P7 |
| #8 | B | High-confidence samples | Samples that the model is sure about tend to represent what the model has learnt and help in understanding the model behaviour. | P7 |
| #9 | B | Similar training samples | Training samples that are similar in features to the test sample and how the model behaves on these is helpful for understanding model behaviour. | (P6), P7, P8 |
| #10 | B | Group of samples | Inspecting multiple training samples is generally more interesting to estimate model behaviour. | P8 |
The figure depicts a three-dimensional coordinate system with categorical axes. The x-axis is named “Action defining relevance” and has the categories: remove, change label, reweight and recommend data collection. The y-axis is named “Metric” and has the categories: Loss, classification probability and ground-truth probability. The z-axis is named “Number of training samples” and has the categories: Explain f(x) with 1 training sample, and explain f(x) with N training samples.
The interview analysis reveals model developers as a key audience for TDA explanations, with model debugging as a promising use case. Before addressing RQ2 – What do users need from TDA explanations?, we translate these insights into a design space of TDA explanations in model debugging. The space shall illustrate how explanatory information about training samples and model output could look like, forming the basis for the study addressing RQ2.
Interview participants provided rich information about what they usually look out for in the training data when debugging a model. We coded this information into the following themes: Data quality issues, and different notions of interesting training samples. We present the descriptions of the corresponding codes in Table 2 and use them to identify different axes of information that could be represented by TDA explanations. Specifically, we find that different data artefacts are debugged using different actions, with users interested in various metrics and considering different numbers of training samples. From these observations, we define a three-dimensional design space (see Figure 3). In the following, we detail the reasoning behind each axis and connect it to the codes in Table 2 via the ID column.
4.3.1. Axis 1: Action defining relevance
The first axis corresponds to the actions participants propose for different types of data artefacts: Removing data, changing the label, reweighting data and collecting more data (x-axis of Figure 3). Participants look at data to remove when it is clearly malicious data and likely hurts model learning, for instance mislabelled (#3) or outdated, historical data (#5). Additionally, we find that participants look for data that they wish to change, for example filling in or imputing missing data (#2), or correcting wrong data scales (#4) and labels (#3). Especially when data is hard to obtain (e.g. in the medical domain), removing data is not an option and correcting data is preferred (P6). Another action is the reweighting of training samples and collecting more data related to the error. These actions emphasise specific samples in the training process to encourage the model to learn them. This is particularly relevant to learning outliers (#6) and handling distribution shifts (#1).
Removal, change and addition of data have different effects on a model and therefore represent different quantifications of attribution: For instance, removal as in traditional TDA (Equation 1 in § 3) or by addition of data. Essentially, these examples map to the questions: How would the model behaviour change if one were to (a) exclude a training sample? or (b) add more training samples like the test sample to the dataset? By understanding preferences in this axis, we aim to identify what is needed for data attribution to be actionable and reflect the debugging process.
4.3.2. Axis 2: Metric
The second axis corresponds to the sample-wise metrics participants used when referring to model behaviour: the loss, the probability of the true class label, and the probability of the predicted class label (y-axis of Figure 3). The interviews show not all participants think of model behaviour in terms of sample loss, but also in classification probability. While high loss can be an indicator for wrong or outlier samples (#3, #6), participants voiced their interest in samples where the model has high or low classification probability (#7, #8). Such samples provide context of the global model behaviour to the model developer, and can therefore be helpful for their work.
The metric can be seen as a measure by which participants primarily perceive model behaviour in the debugging process. By understanding metric preferences, we aim to identify the metric that helps model developers most in debugging, testing whether the loss as in traditional TDA is the most actionable measure.
4.3.3. Axis 3: Number of training samples
The third axis corresponds to the number of training samples participants need to make sense of a model error through the training data. We identify two main options: A single sample and a group of samples (z-axis of Figure 3). We found that model developers usually do not inspect just one training data sample but think more globally as they usually refer to training data and samples in plural form (#10). This aspect is addressed in TDA research by works studying group attribution (Koh et al., 2019; Basu et al., 2020; Park et al., 2023). By understanding preferences in this axis, we aim to identify whether individual attribution like in traditional TDA or rather training data groups need to be in focus of TDA research.
4.4. Survey study
Building on the findings of the interview study, we aim to study the needs of model developers and understand what kind of information is useful, addressing RQ2 – What do users need from TDA explanations? We design an online survey that explores the design space of TDA explanations (see Figure 3) using scenario-based design (Carroll, 1995). We piloted the survey to ensure clear phrasing and compatibility across devices. Data was collected from June – August 2024.
4.4.1. Participants
We recruit participants with prior experience in developing machine learning models to closely align with potential real users. To recruit expert participants, we used snowball sampling alongside social media ads on X and LinkedIn, reaching out to contacts and asking them to share the study link. We expect thematic saturation in the qualitative analysis and meaningful statistics in the quantitative analysis at around 30 replies. Of 34 responses, we excluded three for low-quality answers (short and non-sensical), leaving 31 participants (nine female). 18 participants work in industry while 14 work in academia, with most having 1-5 years of machine learning experience. Four do not work with deep models. Participants come from diverse fields, including biology, health, logistics, and more. While the data types used by participants are diverse too, the majority of participants work with image or tabular data, as used in our study (see Figure 4). The participants were compensated with 25€ via Tremendous111https://www.tremendous.com/.
This figure shows five histograms about the answer distributions for the preliminary demographic questions of the survey study. From the top right corner to the bottom left corner, the histograms show the gender distribution (22 male vs. 9 female), whether people work in industry or academia (13 academia, 17 industry, 1 in both), years of experience distribution (1 ¡1 year, 15 1-3 years, 12 3-5 years, 3 ¿5years), the distribution of data types that participants work with (19 Image, 16 tabular, 14 time series, 15 text, 8 video, 1 biological, 2 graphs, 1 audio and 1 robot data), the distribution of experiences with deep learning models (27 yes vs. 4 no).
4.4.2. Scenario-based design
The main part of the study is constructed with scenario-based design (Carroll, 1995) to base the analysis on a tangible usage scenario. We create two scenarios of model debugging in an interactive mock-up of a model development suite (see Figure 5). The choice of data types (image and tabular data) is based on the dominant modalities of interview participants. The study puts participants in two imaginary scenarios where they are model developers in a company that builds (1) a bird classification app, and (2) a credit scoring app. The company recently acquired a data-centric tool to help model developers debug their models by understanding errors and identifying training data relevant to those errors. The tool is customisable to adapt to the developer’s preferences. The customisation choices, based on the TDA design space (see Figure 3) are:
Action defining relevance: (a) Removal, (b) Label change (to the next most likely class), (c) Upweighting (implemented with a factor of 10). We omit new data collection because it is difficult to predict what kind of data participants would like to add. Instead, we take upweighting as an approximation of adding new data.
Metric: (a) Loss, (b) Ground truth class probability, (c) Predicted class probability.
Number of training samples: Participants can choose between the 1 – 10 training samples to inspect. We limit the number to 10 to explore preferences for more training samples while keeping computation efforts low.
In both scenarios, participants are presented with a specific misclassification as the model error (left side of Figure 5). This error is randomised across participants to prevent bias from a specific sample. Participants are asked to choose which type of training data-related information (right side of Figure 5) is most useful to understand the reasons for the error. This way, we aim to elicit participants’ intuitions about the most actionable information for debugging the model error. Participants are encouraged to explore different settings and submit their desired customisation to complete the task. The data for these scenarios was generated using real machine learning models and openly available datasets. We detail the data generation process in Appendix C.
The figure shows a screenshot of the bird classification scenario interface of the survey study. The page is divided into two parts: On the left is the interface with the header “Customize your own development interface”, and on the right are the customization controls as an expanded menu. The controls have the instructions: “Choose a setting below and click Apply to see the effects.” Below are three drop down lists, one for Action defining relevance, one for Metric of interest and the last one for the number of training samples. The settings are set to “Removal”, “Loss” and “1” respectively. Below the dropdown lists is a grey button with the word Apply on it. The interface itself is divided into two panels as well: on the left, indicated with a “1” is a panel with the title “Test error analysis”. Below, the test accuracy is given as well as a description of the Error (“Wren is misclassified as Sparrow.”). Underneath the error description is the image of the wren that is misclassified with information about the prediction, loss, predictive probability and ground truth probability. The second panel, indicated with a “2” is titled “Training data analysis”. It shows both the individual effect (how does the model behaviour on the erroneous test sample change?) of the setting chosen in the customisation menu as well as the global effect (how does the overall model behaviour change?). Below the effect is a table with the most relevant training samples according to the chosen setting, with their ground truth and predicted labels. The dynamic information is highlighted by an orange box with arrows pointing from the customisation menu to the dynamic panel.
4.4.3. Survey structure
The survey structure is shown in Figure 6. Participants are first briefed on the study’s objective and data processing. After providing consent and confirming they meet inclusion criteria, they are asked preliminary demographic questions for context to the later analysis. In the remaining study, we aim to extract the participant’s preferences through a usage scenario (§ 4.4.2). The participants are presented with the scenario description and prompted with “Imagine you are a model developer in a company […]” to immerse them in the use case. Then, we ask the participants follow-up questions to understand the reasoning behind their choices through free-text questions. These free-text questions are about understanding the participants’ intuitions of why the model may have made an error, and how they would fix it. After completing one use case, the participants are presented with the other use case, where the order is randomised across users to prevent an order bias.
The figure shows a flowchart of the survey. The survey starts with a question about prior model developing experience to assess whether inclusion criteria are met. Then, the next state of preliminary questions follow to record context to the data. This includes demographic questions and questions about prior model development experiences. The third state is the tutorial for the interactive application, after which the quantitative data collection is started with either the bird classification (image) use case or the credit scoring (tabular) use case. In this stage, participants can explore the interface and are asked to submit their final choices. The next stage is the qualitative data collection, where participants are asked to explain their choices and proposed actions in free text. Then, an arrow indicates that the other use case will be presented before the end of the survey.
4.4.4. Data analysis
We analyse the survey study both quantitatively and qualitatively. For the quantitative analysis, we aim to understand whether a specific notion of TDA is preferred across participants because previous TDA studies have been centred around a single notion of TDA (§ 3). Since the traditional TDA setting is already a part of the design space, studying preferences gives an answer to whether participants really preferred this particular scenario. We further assess if participants prefer a specific scenario —- if not, it is a strong signal that the research should be diversified into multiple possible scenarios serving different user needs.
To this end, we compute the distribution of the participants’ choices (action, metric and number of samples). Furthermore, we test whether the collected data indicates a statistically significant preference for a specific setting. We perform a test as the data is categorical (Pearson, 1900). For the number of training samples axis, we perform Fisher’s exact test as the choice frequencies are low (Edwards, 2005). Specifically, our hypotheses are:
In other words, we test the goodness of fit of our observations against a uniform distribution across categories of an axis. Since we conduct multiple tests on the same data, we apply a Bonferroni correction (Weisstein, 2004) and define statistical significance at . Additionally, we compute each axis category’s entropy and normalised entropy (normalised by with as the number of categories in an axis, e.g. three for the metric axis) to get an indicator of the diversity of preferences (Alsakran et al., 2014).
If the entropy is high, the participant’s preferences are highly diverse, meaning there is no agreement on a certain preferred notion of the relevance of training data.
The qualitative analysis aims to understand the reasons behind participant choices and get deeper insight into user preferences. The free-text responses in the follow-up questions are analysed using inductive thematic coding (Braun and Clarke, 2012) by two coders. Per question, we analyse the common themes among participants to extract the reasons behind user preferences and needs. The analysis consists of several coding workshops and is depicted in Figure 7. We measure interrater agreement by the percentage of overlapping themes. The final agreement is 97.1%.
The figure shows the progression of the thematic analysis of the survey data in a timeline-like schematic consisting of 8 parts: The initial coding workshop, where the data of five participants was analysed, resulted in 21 themes. The next stage is independent coding of eight more participants, followed by the first intermediate coding workshop that resulted in 29 themes with an IRR of 75.8%. This is repeated and the next intermediate coding workshop resulted in 31 themes with an IRR of 84.3%. Then, four more participants are independently analysed before the final coding workshop that resulted in 32 themes and and IRR of 97.1%. Lastly, the stage of case-by-case discussions of themes indicates that single participants who answered the survey after the main analysis are discussed on a case-by-case basis.
4.5. Survey findings
We conduct the survey study to answer RQ2 – What do users need from TDA explanations? The survey data analysis indicates a clear connection between users’ intuition of what causes a model error and their TDA preferences. We elaborate further in the following.
4.5.1. Quantitative results
The quantitative analysis builds on the customisation choices of the debugging interface (see Quantitative data collection in Figure 6). Table 3 shows the frequency of selections and the associated p-values. We test if there exists a setting of action, metric and number of training samples which is consistently chosen and particularly useful to model developers. We are especially interested in whether there is a clear preference for removal, loss and one training sample as these correspond to the traditional notion of TDA introduced in § 3. For the action axis, there is no statistically significant preference for any choice and the distribution is rather uniform across choices. While the results may show a higher frequency of participants choosing loss as a metric, we observe no statistically significant preference. Hence, we fail to reject for the action and metric axes and conclude that there is no dominant preference for a particular TDA setting. For the number of samples axis, however, we find a statistically significant preference for a group of 10 samples.
The entropy analysis supports this conclusion (see Table 4). As a measure of diversity, normalised entropy scores range from 0.588 to 0.997, showing a highly diverse (¿0.5) distribution of choices. This is especially true for the Action defining relevance and Metric axes with high scores (¿0.8). There is less diversity in the number of training samples but a clear preference for choosing groups of samples, confirming Ilyas et al.’s (Ilyas et al., 2022) intuitions for studying group attribution.
In conclusion, the quantitative analysis reveals that the preferred TDA explanations represent highly diverse information with a clear need for group attribution.
| Image | Tabular | Both | ||||
| Choice | % | % | % | |||
| Action | ||||||
| Removal | 33.33% | 1 | 41.97% | 1 | 37.70% | 1 |
| Label Change | 30.00% | 1 | 16.13% | 1 | 22.95% | 1 |
| Upweighting | 36.67% | 1 | 41.94% | 1 | 39.34% | 1 |
| Metric | ||||||
| Loss | 60.00% | 1 | 35.48% | 1 | 47.54% | 1 |
| Pred. cls. prob. | 10.00% | 1 | 41.94% | 1 | 26.23% | 1 |
| GT class prob. | 30.00% | 1 | 22.58% | 1 | 26.23% | 1 |
| Image | Tabular | Both | ||||
| Choice | % | % | % | |||
| Num. of samples | ||||||
| 1 | 0.00% | 1 | 0.00% | 1 | 0.00% | 1 |
| 2 | 0.00% | 1 | 0.00% | 1 | 0.00% | 1 |
| 3 | 3.33% | 1 | 9.68% | 1 | 6.56% | 1 |
| 4 | 6.67% | 1 | 3.23% | 1 | 4.92% | 1 |
| 5 | 13.33% | 1 | 19.35% | 1 | 16.39% | 1 |
| 6 | 3.33% | 1 | 3.23% | 1 | 3.28% | 1 |
| 7 | 0.00% | 1 | 3.23% | 1 | 1.64% | 1 |
| 8 | 3.33% | 1 | 0.00% | 1 | 1.64% | 1 |
| 9 | 3.33% | 1 | 0.00% | 1 | 1.64% | 1 |
| 10 | 66.67% | ¡0.001 | 61.29% | 0.002 | 63.93% | ¡0.001 |
| Image | Tabular | Both | ||||
| Axis | ||||||
| Action | 1.095 | 0.997 | 1.023 | 0.931 | 1.073 | 0.976 |
| Metric | 0.898 | 0.817 | 1.068 | 0.972 | 1.056 | 0.961 |
| Num. samples | 1.173 | 0.603 | 1.176 | 0.657 | 1.224 | 0.588 |
4.5.2. Qualitative results
The figure shows the themes found in the coding analysis of the survey study, organised into four theme areas. The first theme area is called “Action depends on user’s hypotheses” and includes the themes on different hypotheses that participants have about the root cause of the model error (mislabelled data, incomplete training, spurious correlations, similar class features, low dataset diversity, class overlap, dataset imbalance and small dataset size) as well as themes related to the proposed bugfixing actions users proposed (Remove noisy data, changing label to correct, fix labelling pipeline for future data, training process engineering, model architecture engineering, feature engineering, upweight data to focus on specific features, feature attribution inspection, manual inspection of datapoints, collect specific new data, synthesize/augment data, resample/reweigh data and collect any new data). The second theme area is called “Individual metric preferences” and includes themes on preferred metrics (probabilities are intuitive and show the decision boundary, ground truth probability is the target, loss is most informative in the multi-class case) and one theme related to preferred actions (Label change together with a probability metric show the decision boundary). The third theme area is titled “Group attribution preferences” with themes Group for overview of the error, Balance betw. variety and conciseness. The last theme area is titled “Reliability of explanations” and includes themes related to the preferred number of training samples (group size enough for reliable attribution, small group to prevent reasoning change), preferred metric (Ground truth probability is the target) and preferred action (Removal shows causal effect).
The inductive thematic coding analysis of the free-text answers resulted in four thematic areas and theme types covering 30 themes (see Figure 8): (1) Action depends on the user’s hypothesis, (2) Individual metric preferences, (3) Group attribution preferences, and (4) Reliability of explanations. For the theme types, we find themes that correspond to the user’s hypothesis about the error and observe that these are often coded together with a corresponding proposed action. However, preferred metrics and preferred number of training samples do not correspond to the user’s hypothesis. Additionally, we identify four themes that do not correspond to any category but give interesting insights and context to the analysis. The full description of each theme is in Appendix D.
Action depends on user’s hypotheses.
The thematic analysis complements our quantitative findings: User preferences in model debugging are highly individual. The analysis provides a reason: Individual preferences are often linked to the user’s understanding and intuitions, in other words, their hypothesis about an error. The user’s hypothesis determines their course of action. Hypotheses are highly individual because users formulate them based on their expertise and past experiences. For example, we identified the theme of low dataset diversity as a hypothesis participants have about a model error: “[Images] […] in the training set are not diverse enough”, “[One class] was not representative enough”. Corresponding to this hypothesis, participants often suggested adding new data or applying data augmentations to diversify the training set, which we identified as collect specific new data and synthesise/augment data themes.
Individual metric preferences.
For individual metric preferences, we identify a link with the task at hand. In the case of binary classification, participants showed a preference for probability metrics, as it is easier to understand than the loss and allows an inspection of the decision boundary (“The loss is not very intuitive”). In multi-class classification, more participants preferred the loss as it considers all possible classes (“[The loss] contains information on all classes”). However, as the quantitative analysis shows, these differences in preference are not statistically significant. Therefore, the metric preferences are likely individual to the user and their habits.
Group attribution preferences.
The quantitative study shows a clear preference for attribution to groups of training samples as opposed to individual training samples. The qualitative analysis finds that this is due to the user’s need for having a good overview of the model error. Inspecting a group of training samples provides a better overview than a single sample, and allows the user to understand whether their hypothesis about the error is valid or not (Group for overview of the error). A second theme highlights another angle: While variety in the group is desired for a good overview of the error (e.g. group consisting of samples with diverse features), the group should be concise enough to “keep the analysis manageable” (Balance betw. variety and conciseness).
Reliability of explanations.
Our analysis finds that some themes are related to the overall concept of explanation reliability, where particularly the choice of how many training samples to inspect is associated with reliability. Participants stated that group sizes should be large enough for the attribution scores to be reliable. At the same time, he group size should be small enough to ensure reliable attribution scores, as changing the dataset could change the model’s reasoning (Small group to prevent reasoning change). We interpret this need for reliability to reflect in the motivation to uncover causal effects (“By removing […] I can determine if [the training samples] were causing confusion within the model”).
Unintended themes.
We identified four unintended themes outside the study design that provide additional context. We noticed that participants focused either on the model error (Focus on model error) or the final accuracy (Focus on final accuracy). Therefore, some of the participants explored all choice options in the survey and submitted the setting with the largest performance gain. This indicates that the end goal of performance improvement in the task is highly valued. Some participants mentioned that the study interface is unfamiliar (Unfamiliar study interface), which can be confusing. We considered this point in the analysis through discussions at the coding workshops. However, it also indicates that participants were unfamiliar with TDA explanations overall as they are not used in practice yet. The last theme highlights the importance of grouping erroneous test samples. Participants noted that a single erroneous test sample may represent a broader issue, and understanding the link between training data and groups of test samples with similar errors could enhance debugging.
5. Discussion
This work presents a needfinding study with AI practitioners for training data attribution (TDA) explanations to inform user-focused TDA research in a top-down approach. We summarise our findings to provide answers to the initial research questions (§ 5.1), identify research gaps that need to be addressed for more user-focused TDA (§ 5.2) and discuss limitations (§ 5.3).
5.1. Answers to the research questions
5.1.1. RQ1: What do users need TDA explanations for?
To answer this question, we conducted the interview study (§ 4.1). Our findings suggest that TDA as an explanation method has the potential to serve similar high-level purposes as other explanation techniques, namely enabling users to understand model behaviour, especially unexpected model behaviour. By offering insights into model behaviour, TDA can assist end-users in calibrating trust by providing context for model decisions and support model developers in debugging errors by identifying root causes in the training data. This data-centric approach is particularly valuable in model development, leading us to identify model developers as the primary audience for TDA explanations and model debugging as the key use case.
5.1.2. RQ2: What do users need from TDA explanations?
To answer this question, we designed a survey study that extracts user preferences for the information represented by TDA (§ 4.4). Our study reveals that users need different types of information to understand model behaviour, depending on their hypotheses of why the model errs during development. This necessitates TDA explanations that are flexible and adaptable to individual needs. These needs vary based on the application domain and the user’s level of expertise. Experienced users, may form hypotheses about model errors based on their past experiences. In contrast, less experienced users may form hypotheses based on what they learnt previously about machine learning models. Users have a preference for group attribution, as groups provide more information than individual samples. Additionally, users need TDA explanations to be reliable considering how data-centric actions might affect the model as a whole. Reliability entails two main aspects: First, the information needs to be reliable for it to be actionable and effective for the task (i.e., model debugging). Second, TDA explanations should also account for any broader changes in the model’s behaviour that may affect the attribution to training samples.
5.1.3. RQ3: To what extent do existing TDA approaches align with user needs?
To answer this question, we reflect on existing TDA approaches in light of the identified needs: TDA explanations need to be adaptable to the user and provide reliable and actionable insight while maintaining consistency with the overall model behaviour. From these needs, we identify different TDA tasks that users need, next to the overarching need for reliability (see Table 5).
| Attribution type | Action | Metric | Times chosen | ||
| Influence attribution (e.g., (Koh and Liang, 2017; Pruthi et al., 2020; Guo et al., 2021; Schioppa et al., 2022; Grosse et al., 2023; Park et al., 2023)) | Removal | Sample loss | 1 | 1 | 0 |
| Group influence attribution (e.g., (Park et al., 2023; Basu et al., 2020; Koh et al., 2019; Lin et al., 2024)) | Removal | Sample loss | ¿1 | 1 | 11 |
| Attributing errors (e.g. (Jain et al., 2024)) | Any | Any | ¿1 | ¿1 | 4 |
| Lacking data attribution | Addition | Any | ¿1 | ¿1 | 26 |
| Decision boundary attribution | Label change | Any | ¿1 | ¿1 | 14 |
Traditional TDA measures the influence of a training sample on model predictions by analysing the impact of its removal through leave-one-out retraining (Hampel, 1974). This method essentially evaluates model predictions in the context of a counterfactual change in the training data. It has been studied across different tasks, such as identifying mislabelled data, understanding adversarial vulnerabilities, and detecting domain mismatches (Koh and Liang, 2017). Mislabel identification, in particular, is a common evaluation strategy employed in several studies (Pruthi et al., 2020; Park et al., 2023; Schioppa et al., 2022). Consequently, traditional TDA partially addresses the need for adaptability to different use cases. However, since TDA quantifies the relevance of a training sample by its removal, it is particularly suited for scenarios where data elimination is necessary, such as malicious samples. It remains an open question whether alternative actions, such as changing labels or upsampling a data region, might provide improved explanations in terms of actionability. Our study indicates that users may have diverse ideas regarding the root causes of errors, and traditional TDA only addresses one of these perspectives.
Regarding the need for reliable, actionable, and effective information, existing TDA methods have been criticised for their low reliability in deep learning (Nguyen et al., 2023; Basu et al., 2021; Bae et al., 2022). In addition, our study introduces a different aspect of reliability that only becomes apparent when considering the full context of the use case (as we assessed users within the context of a specific task). When a dataset is modified, the overall model behaviour can be affected and must be considered. For TDA, this means that when modifying the dataset to addressing a particular error, users require explanations that not only suggest an effective modification but also ensure the model’s performance remains rather consistent.
Our analysis underscores the importance of incorporating user perspectives to identify needs critical to the real-world application of this technology. Traditional TDA explanations only partially meet these needs, and further interdisciplinary efforts are required to develop methods with improved actionability.
5.2. Overlooked research topics for user-focused training data attribution
This work adds to a collection of human-centred explainable AI (HCXAI) works calling to centre XAI research around the user (e.g., (Brennen, 2020; Kaur et al., 2020; Rutjes et al., 2019; Ehsan et al., 2024; Mei et al., 2023; Lin et al., 2024)). By focusing on TDA explanations specifically, we identify several research areas that are necessary to study to advance TDA research towards a focus on users.
Mental models and TDA: Our study supports previous work (Rutjes et al., 2019; Kaur et al., 2020) that the most useful explanations depend on the user’s mental model of why the model errs: We find that what information is actionable to users depends in their hypothesis for reasons of model behaviour (Theme area: Action depends on user’s hypothesis). Therefore, future research should explore how TDA influences and is influenced by mental models, as well as how TDA can be adapted to reflect mental models (e.g. by defining relevance through the action of upweighting as opposed to removal), potentially leading to more aligned and actionable explanations that are well-defined.
User-centred group attribution: The survey findings reveal a clear need for attributing model behaviour to groups of training samples, as it provides more informative insights (Theme area: Group attribution preference). While existing approaches (Basu et al., 2020; Park et al., 2023; Koh et al., 2019) address this, our findings highlight the need for creating groupings that make sense to users (e.g., based on ground truth or predicted class), an aspect not yet widely explored in TDA research.
Holistic understanding of model errors: Our thematic analysis of survey responses highlights the importance of understanding the model error itself before attempting to fix it (Theme: Grouping erroneous test samples). If the root cause is a spurious correlation, the error likely extends beyond the current test sample and may not fully represent the issue. We recommend future research focus on understanding error types rather than isolated instances (e.g., slice discovery (Wang et al., 2024a)) and on attributing model behaviour on groups of test samples to training data (e.g., (Jain et al., 2024)).
Reliability of TDA: TDA quantifies the impact of training samples on model errors by predicting changes in model behaviour, serving as an explanation to the user. For the explanations to be informative, the attribution must be reliable. It is essential to understand how overall model behaviour will be affected as mentioned by our participants (Theme area: Reliability of explanations). We recommend future research to extend existing work on the fragility, reliability, and stability of TDA (Basu et al., 2021; Nguyen et al., 2023; Epifano et al., 2023; Bae et al., 2022) to address these needs.
TDA and feature attribution: Combining feature attribution with TDA could create TDA explanations familiar to model developers who already use feature attribution methods like SHAP (Lundberg and Lee, 2017). While our participants indicated they frequently use feature attribution, there is a known risk of misinterpretation and overreliance (Kaur et al., 2020; Kim et al., 2022). Providing feature attribution explanations with context, such as relevant training data identified through TDA, could help in checking whether similar features in training and test data might mislead the model.
5.3. Limitations
This study has several limitations. First, the use of purposive and snowball sampling techniques in participant recruitment may have introduced selection bias. Consequently, the sample may not be fully representative of the broader population of professionals working with machine learning or model developers in general. Second, the relatively small sample size in our studies limits the robustness of the quantitative analyses, so the results should be interpreted with caution and lead us to focus more strongly on the qualitative analysis. Third, we simulated the use of TDA explanations in a model debugging interface which are lab-like conditions for the survey study. Hence, the needfinding analysis is not rooted in the real workflows and is bound to our study conditions, including the interface design.
6. Conclusion
This paper proposes a top-down approach to conducting explainable AI (XAI) research which is driven by user needs as opposed to techno-centric solutions. We illustrate this approach with the emerging subfield of training data attribution (TDA) and present a needfinding study with AI practitioners. The study consists of two stages: First, we conduct semi-structured interviews with professionals who already work with AI systems (N=10) to gain an impression of common use cases and challenges. We find that TDA could especially benefit model developers, as their workflows are strongly data-centric. We derive a design space of TDA explanatory information that represents what relevant data for model debugging could look like. Second, we develop an interactive interface to extract user preferences in the design space and integrate it into a survey study (N=31) targeted at model developers. The mixed-methods analysis shows that the explanatory information needed from TDA for model debugging is highly dependent not only on the use case but also on the user and their intuitions. Finally, we highlight several research directions in user-focused TDA that are largely overlooked in the current landscape.
Acknowledgements.
We first and foremost thank all study participants who gave us their time and input. We also thank Albert Catalán, Nikita Kister, and Arnas Uselis for participating in the pilot study and helping us improve the study design. We thank Dustin Theobald for sharing his survey website codebase which we used as a basis. We thank Kristina Kapanova for helping us host the survey securely, and everyone who helped spread the survey study. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Elisa Nguyen. This work was supported by the Tübingen AI Center.References
- (1)
- Adebayo et al. (2018) Julius Adebayo, Justin Gilmer, Michael Muelly, Ian Goodfellow, Moritz Hardt, and Been Kim. 2018. Sanity checks for saliency maps. Advances in neural information processing systems 31 (2018).
- Akyürek et al. (2022) Ekin Akyürek, Tolga Bolukbasi, Frederick Liu, Binbin Xiong, Ian Tenney, Jacob Andreas, and Kelvin Guu. 2022. Towards Tracing Knowledge in Language Models Back to the Training Data. Findings of the Association for Computational Linguistics: EMNLP 2022 (2022), 2429–2446.
- Alsakran et al. (2014) Jamal Alsakran, Xiaoke Huang, Ye Zhao, Jing Yang, and Karl Fast. 2014. Using entropy-related measures in categorical data visualization. In 2014 IEEE Pacific Visualization Symposium. IEEE, 81–88.
- Bae et al. (2024) Juhan Bae, Wu Lin, Jonathan Lorraine, and Roger Grosse. 2024. Training Data Attribution via Approximate Unrolled Differentation. arXiv preprint arXiv:2405.12186 (2024).
- Bae et al. (2022) Juhan Bae, Nathan Ng, Alston Lo, Marzyeh Ghassemi, and Roger B Grosse. 2022. If Influence Functions are the Answer, Then What is the Question?. In Advances in Neural Information Processing Systems, S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, and A. Oh (Eds.), Vol. 35. Curran Associates, Inc., 17953–17967. https://proceedings.neurips.cc/paper_files/paper/2022/file/7234e0c36fdbcb23e7bd56b68838999b-Paper-Conference.pdf
- Basu et al. (2021) Samyadeep Basu, Phil Pope, and Soheil Feizi. 2021. Influence Functions in Deep Learning Are Fragile. In International Conference on Learning Representations. https://openreview.net/forum?id=xHKVVHGDOEk
- Basu et al. (2020) Samyadeep Basu, Xuchen You, and Soheil Feizi. 2020. On second-order group influence functions for black-box predictions. In International Conference on Machine Learning. PMLR, 715–724.
- Bilodeau et al. (2024) Blair Bilodeau, Natasha Jaques, Pang Wei Koh, and Been Kim. 2024. Impossibility theorems for feature attribution. Proceedings of the National Academy of Sciences 121, 2 (2024), e2304406120.
- Braun and Clarke (2012) Virginia Braun and Victoria Clarke. 2012. Thematic analysis. American Psychological Association.
- Brennen (2020) Andrea Brennen. 2020. What Do People Really Want When They Say They Want ”Explainable AI?” We Asked 60 Stakeholders.. In Extended Abstracts of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI EA ’20). Association for Computing Machinery, New York, NY, USA, 1–7. https://doi.org/10.1145/3334480.3383047
- Brunet et al. (2019) Marc-Etienne Brunet, Colleen Alkalay-Houlihan, Ashton Anderson, and Richard Zemel. 2019. Understanding the origins of bias in word embeddings. In International conference on machine learning. PMLR, 803–811.
- Carroll (1995) John M. Carroll (Ed.). 1995. Scenario-based design: envisioning work and technology in system development. John Wiley & Sons, Inc., USA.
- Charpiat et al. (2019) Guillaume Charpiat, Nicolas Girard, Loris Felardos, and Yuliya Tarabalka. 2019. Input Similarity from the Neural Network Perspective. In Advances in Neural Information Processing Systems, H. Wallach, H. Larochelle, A. Beygelzimer, F. d'Alché-Buc, E. Fox, and R. Garnett (Eds.), Vol. 32. Curran Associates, Inc. https://proceedings.neurips.cc/paper_files/paper/2019/file/c61f571dbd2fb949d3fe5ae1608dd48b-Paper.pdf
- Choe et al. (2024) Sang Keun Choe, Hwijeen Ahn, Juhan Bae, Kewen Zhao, Minsoo Kang, Youngseog Chung, Adithya Pratapa, Willie Neiswanger, Emma Strubell, Teruko Mitamura, et al. 2024. What is Your Data Worth to GPT? LLM-Scale Data Valuation with Influence Functions. arXiv preprint arXiv:2405.13954 (2024).
- Doshi-Velez and Kim (2017) Finale Doshi-Velez and Been Kim. 2017. Towards a rigorous science of interpretable machine learning. arXiv preprint arXiv:1702.08608 (2017).
- Dwivedi et al. (2023) Rudresh Dwivedi, Devam Dave, Het Naik, Smiti Singhal, Rana Omer, Pankesh Patel, Bin Qian, Zhenyu Wen, Tejal Shah, Graham Morgan, et al. 2023. Explainable AI (XAI): Core ideas, techniques, and solutions. Comput. Surveys 55, 9 (2023), 1–33.
- Edwards (2005) Anthony WF Edwards. 2005. RA Fischer, statistical methods for research workers, (1925). In Landmark writings in western mathematics 1640-1940. Elsevier, 856–870.
- Ehsan et al. (2021) Upol Ehsan, Q Vera Liao, Michael Muller, Mark O Riedl, and Justin D Weisz. 2021. Expanding explainability: Towards social transparency in ai systems. In Proceedings of the 2021 CHI conference on human factors in computing systems. 1–19.
- Ehsan et al. (2024) Upol Ehsan, Samir Passi, Q. Vera Liao, Larry Chan, I-Hsiang Lee, Michael Muller, and Mark O Riedl. 2024. The Who in XAI: How AI Background Shapes Perceptions of AI Explanations. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 316, 32 pages. https://doi.org/10.1145/3613904.3642474
- Ehsan and Riedl (2020) Upol Ehsan and Mark O. Riedl. 2020. Human-Centered Explainable AI: Towards a Reflective Sociotechnical Approach. In HCI International 2020 - Late Breaking Papers: Multimodality and Intelligence, Constantine Stephanidis, Masaaki Kurosu, Helmut Degen, and Lauren Reinerman-Jones (Eds.). Springer International Publishing, Cham, 449–466.
- Epifano et al. (2023) Jacob Epifano, Ravichandran Ramachandran, Aaron J. Masino, and Ghulam Rasool. 2023. Revisiting the Fragility of Influence Functions. Neural networks : the official journal of the International Neural Network Society 162 (2023), 581–588.
- Feldman and Zhang (2020) Vitaly Feldman and Chiyuan Zhang. 2020. What neural networks memorize and why: Discovering the long tail via influence estimation. Advances in Neural Information Processing Systems 33 (2020), 2881–2891.
- Fok and Weld (2023) Raymond Fok and Daniel S Weld. 2023. In search of verifiability: Explanations rarely enable complementary performance in AI-advised decision making. AI Magazine (2023).
- Grosse et al. (2023) Roger Grosse, Juhan Bae, Cem Anil, Nelson Elhage, Alex Tamkin, Amirhossein Tajdini, Benoit Steiner, Dustin Li, Esin Durmus, Ethan Perez, et al. 2023. Studying large language model generalization with influence functions. arXiv preprint arXiv:2308.03296 (2023).
- Guest et al. (2023) Greg Guest, Emily E. Namey, and Marilyn L. Mitchell. 2023. Collecting Qualitative Data: A Field Manual for Applied Research. SAGE Publications, Ltd, 55 City Road. https://doi.org/10.4135/9781506374680
- Guidotti et al. (2018) Riccardo Guidotti, Anna Monreale, Salvatore Ruggieri, Franco Turini, Fosca Giannotti, and Dino Pedreschi. 2018. A Survey of Methods for Explaining Black Box Models. ACM Comput. Surv. 51, 5, Article 93 (2018), 42 pages. https://doi.org/10.1145/3236009
- Guo et al. (2021) Han Guo, Nazneen Rajani, Peter Hase, Mohit Bansal, and Caiming Xiong. 2021. FastIF: Scalable Influence Functions for Efficient Model Interpretation and Debugging. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics, Online and Punta Cana, Dominican Republic, 10333–10350. https://doi.org/10.18653/v1/2021.emnlp-main.808
- Hadash et al. (2022) Sophia Hadash, Martijn C. Willemsen, Chris Snijders, and Wijnand A. IJsselsteijn. 2022. Improving understandability of feature contributions in model-agnostic explainable AI tools. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (New Orleans, LA, USA) (CHI ’22). Association for Computing Machinery, New York, NY, USA, Article 487, 9 pages. https://doi.org/10.1145/3491102.3517650
- Hammoudeh and Lowd (2024) Zayd Hammoudeh and Daniel Lowd. 2024. Training data influence analysis and estimation: A survey. Machine Learning 113, 5 (2024), 2351–2403.
- Hampel (1974) Frank R. Hampel. 1974. The Influence Curve and its Role in Robust Estimation. J. Amer. Statist. Assoc. 69, 346 (1974), 383–393. https://doi.org/10.1080/01621459.1974.10482962
- He et al. (2016) Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun. 2016. Deep residual learning for image recognition. In Proceedings of the IEEE conference on computer vision and pattern recognition. 770–778.
- Hofmann (1994) Hans Hofmann. 1994. Statlog (German Credit Data). UCI Machine Learning Repository. DOI: https://doi.org/10.24432/C5NC77.
- Ilyas et al. (2022) Andrew Ilyas, Sung Min Park, Logan Engstrom, Guillaume Leclerc, and Aleksander Madry. 2022. Datamodels: Predicting Predictions from Training Data. In Proceedings of the 39th International Conference on Machine Learning.
- Jain et al. (2024) Saachi Jain, Kimia Hamidieh, Kristian Georgiev, Andrew Ilyas, Marzyeh Ghassemi, and Aleksander Madry. 2024. Data Debiasing with Datamodels (D3M): Improving Subgroup Robustness via Data Selection. arXiv preprint arXiv:2406.16846 (2024).
- Jin et al. (2023) Weina Jin, Jianyu Fan, Diane Gromala, Philippe Pasquier, and Ghassan Hamarneh. 2023. Invisible users: Uncovering end-users’ requirements for explainable ai via explanation forms and goals. arXiv preprint arXiv:2302.06609 (2023).
- Jin et al. (2024) Wei Jin, Haohan Wang, Daochen Zha, Qiaoyu Tan, Yao Ma, Sharon Li, and Su-In Lee. 2024. DCAI: Data-centric Artificial Intelligence. In Companion Proceedings of the ACM Web Conference 2024 (Singapore, Singapore) (WWW ’24). Association for Computing Machinery, New York, NY, USA, 1482–1485. https://doi.org/10.1145/3589335.3641297
- Juneja et al. (2024) Prerna Juneja, Wenjuan Zhang, Alison Marie Smith-Renner, Hemank Lamba, Joel Tetreault, and Alex Jaimes. 2024. Dissecting users’ needs for search result explanations. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 841, 17 pages. https://doi.org/10.1145/3613904.3642059
- Kaur et al. (2020) Harmanpreet Kaur, Harsha Nori, Samuel Jenkins, Rich Caruana, Hanna Wallach, and Jennifer Wortman Vaughan. 2020. Interpreting Interpretability: Understanding Data Scientists’ Use of Interpretability Tools for Machine Learning. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–14. https://doi.org/10.1145/3313831.3376219
- Kim et al. (2022) Sunnie SY Kim, Nicole Meister, Vikram V Ramaswamy, Ruth Fong, and Olga Russakovsky. 2022. HIVE: Evaluating the human interpretability of visual explanations. In European Conference on Computer Vision. Springer, 280–298.
- Kim et al. (2023) Sunnie S. Y. Kim, Elizabeth Anne Watkins, Olga Russakovsky, Ruth Fong, and Andrés Monroy-Hernández. 2023. ”Help Me Help the AI”: Understanding How Explainability Can Support Human-AI Interaction. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 250, 17 pages. https://doi.org/10.1145/3544548.3581001
- Koh and Liang (2017) Pang Wei Koh and Percy Liang. 2017. Understanding Black-box Predictions via Influence Functions. In Proceedings of the 34th International Conference on Machine Learning (Proceedings of Machine Learning Research, Vol. 70), Doina Precup and Yee Whye Teh (Eds.). PMLR, 1885–1894. https://proceedings.mlr.press/v70/koh17a.html
- Koh et al. (2019) Pang Wei W Koh, Kai-Siang Ang, Hubert Teo, and Percy S Liang. 2019. On the accuracy of influence functions for measuring group effects. Advances in neural information processing systems 32 (2019).
- Kwon et al. (2023) Yongchan Kwon, Eric Wu, Kevin Wu, and James Zou. 2023. Datainf: Efficiently estimating data influence in lora-tuned llms and diffusion models. arXiv preprint arXiv:2310.00902 (2023).
- Lakkaraju et al. (2022) Himabindu Lakkaraju, Dylan Slack, Yuxin Chen, Chenhao Tan, and Sameer Singh. 2022. Rethinking explainability as a dialogue: A practitioner’s perspective. arXiv preprint arXiv:2202.01875 (2022).
- Liao et al. (2020) Q. Vera Liao, Daniel Gruen, and Sarah Miller. 2020. Questioning the AI: Informing Design Practices for Explainable AI User Experiences. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3313831.3376590
- Lin et al. (2024) Chris Lin, Mingyu Lu, Chanwoo Kim, and Su-In Lee. 2024. Efficient Shapley Values for Attributing Global Properties of Diffusion Models to Data Group. arXiv preprint arXiv:2407.03153 (2024).
- Lundberg and Lee (2017) Scott M Lundberg and Su-In Lee. 2017. A unified approach to interpreting model predictions. Advances in neural information processing systems 30 (2017).
- Mei et al. (2023) Alex Mei, Michael Saxon, Shiyu Chang, Zachary C Lipton, and William Yang Wang. 2023. Users are the north star for ai transparency. arXiv preprint arXiv:2303.05500 (2023).
- Miller (2019) Tim Miller. 2019. Explanation in artificial intelligence: Insights from the social sciences. Artificial intelligence 267 (2019), 1–38.
- Nauta et al. (2023) Meike Nauta, Jan Trienes, Shreyasi Pathak, Elisa Nguyen, Michelle Peters, Yasmin Schmitt, Jörg Schlötterer, Maurice van Keulen, and Christin Seifert. 2023. From Anecdotal Evidence to Quantitative Evaluation Methods: A Systematic Review on Evaluating Explainable AI. ACM Comput. Surv. 55, 13s, Article 295 (jul 2023), 42 pages. https://doi.org/10.1145/3583558
- Naveed et al. (2023) Humza Naveed, Asad Ullah Khan, Shi Qiu, Muhammad Saqib, Saeed Anwar, Muhammad Usman, Naveed Akhtar, Nick Barnes, and Ajmal Mian. 2023. A comprehensive overview of large language models. arXiv preprint arXiv:2307.06435 (2023).
- Nguyen et al. (2023) Elisa Nguyen, Minjoon Seo, and Seong Joon Oh. 2023. A Bayesian Approach To Analysing Training Data Attribution in Deep Learning. In Proceedings of the 2023 Conference on Neural Information Processing Systems.
- OpenAI (2022) OpenAI. 2022. ChatGPT-3.5. https://chat.openai.com
- Park et al. (2023) Sung Min Park, Kristian Georgiev, Andrew Ilyas, Guillaume Leclerc, and Aleksander Madry. 2023. TRAK: Attributing Model Behavior at Scale. In International Conference on Machine Learning (ICML).
- Parliament (2023) European Parliament. 2023. AI Act, Annex III.
- Paszke et al. (2019) Adam Paszke, Sam Gross, Francisco Massa, Adam Lerer, James Bradbury, Gregory Chanan, Trevor Killeen, Zeming Lin, Natalia Gimelshein, Luca Antiga, Alban Desmaison, Andreas Kopf, Edward Yang, Zachary DeVito, Martin Raison, Alykhan Tejani, Sasank Chilamkurthy, Benoit Steiner, Lu Fang, Junjie Bai, and Soumith Chintala. 2019. PyTorch: An Imperative Style, High-Performance Deep Learning Library. In Advances in Neural Information Processing Systems 32. Curran Associates, Inc., 8024–8035. http://papers.neurips.cc/paper/9015-pytorch-an-imperative-style-high-performance-deep-learning-library.pdf
- Patnaik and Becker (1999) Dev Patnaik and Robert Becker. 1999. Needfinding: The Why and How of Uncovering People’s Needs. Design Management Journal (Former Series) 10, 2 (1999), 37–43. https://doi.org/10.1111/j.1948-7169.1999.tb00250.x arXiv:https://onlinelibrary.wiley.com/doi/pdf/10.1111/j.1948-7169.1999.tb00250.x
- Pearson (1900) Karl Pearson. 1900. X. On the criterion that a given system of deviations from the probable in the case of a correlated system of variables is such that it can be reasonably supposed to have arisen from random sampling. The London, Edinburgh, and Dublin Philosophical Magazine and Journal of Science 50, 302 (1900), 157–175.
- Pedregosa et al. (2011) F. Pedregosa, G. Varoquaux, A. Gramfort, V. Michel, B. Thirion, O. Grisel, M. Blondel, P. Prettenhofer, R. Weiss, V. Dubourg, J. Vanderplas, A. Passos, D. Cournapeau, M. Brucher, M. Perrot, and E. Duchesnay. 2011. Scikit-learn: Machine Learning in Python. Journal of Machine Learning Research 12 (2011), 2825–2830.
- Pezeshkpour et al. (2022) Pouya Pezeshkpour, Sarthak Jain, Sameer Singh, and Byron Wallace. 2022. Combining Feature and Instance Attribution to Detect Artifacts. In Findings of the Association for Computational Linguistics: ACL 2022. Association for Computational Linguistics, Dublin, Ireland, 1934–1946. https://doi.org/10.18653/v1/2022.findings-acl.153
- Pruthi et al. (2020) Garima Pruthi, Frederick Liu, Satyen Kale, and Mukund Sundararajan. 2020. Estimating Training Data Influence by Tracing Gradient Descent. In Advances in Neural Information Processing Systems, H. Larochelle, M. Ranzato, R. Hadsell, M.F. Balcan, and H. Lin (Eds.), Vol. 33. Curran Associates, Inc., 19920–19930. https://proceedings.neurips.cc/paper_files/paper/2020/file/e6385d39ec9394f2f3a354d9d2b88eec-Paper.pdf
- Radford et al. (2023) Alec Radford, Jong Wook Kim, Tao Xu, Greg Brockman, Christine Mcleavey, and Ilya Sutskever. 2023. Robust Speech Recognition via Large-Scale Weak Supervision. In Proceedings of the 40th International Conference on Machine Learning (Proceedings of Machine Learning Research, Vol. 202), Andreas Krause, Emma Brunskill, Kyunghyun Cho, Barbara Engelhardt, Sivan Sabato, and Jonathan Scarlett (Eds.). PMLR, 28492–28518. https://proceedings.mlr.press/v202/radford23a.html
- Ribeiro et al. (2016) Marco Tulio Ribeiro, Sameer Singh, and Carlos Guestrin. 2016. ” Why should i trust you?” Explaining the predictions of any classifier. In Proceedings of the 22nd ACM SIGKDD international conference on knowledge discovery and data mining. 1135–1144.
- Rong et al. (2023) Yao Rong, Tobias Leemann, Thai-Trang Nguyen, Lisa Fiedler, Peizhu Qian, Vaibhav Unhelkar, Tina Seidel, Gjergji Kasneci, and Enkelejda Kasneci. 2023. Towards Human-Centered Explainable AI: A Survey of User Studies for Model Explanations. IEEE Trans. Pattern Anal. Mach. Intell. 46, 4 (nov 2023), 2104–2122. https://doi.org/10.1109/TPAMI.2023.3331846
- Rutjes et al. (2019) Heleen Rutjes, Martijn Willemsen, and Wijnand IJsselsteijn. 2019. Considerations on explainable AI and users’ mental models. In CHI 2019 Workshop: Where is the human? Bridging the gap between AI and HCI. Association for Computing Machinery, Inc.
- Schioppa et al. (2022) Andrea Schioppa, Polina Zablotskaia, David Vilar, and Artem Sokolov. 2022. Scaling Up Influence Functions. Proceedings of the AAAI Conference on Artificial Intelligence 36, 8 (Jun. 2022), 8179–8186. https://doi.org/10.1609/aaai.v36i8.20791
- Shen et al. (2023) Hua Shen, Chieh-Yang Huang, Tongshuang Wu, and Ting-Hao Kenneth Huang. 2023. ConvXAI: Delivering Heterogeneous AI Explanations via Conversations to Support Human-AI Scientific Writing. In Companion Publication of the 2023 Conference on Computer Supported Cooperative Work and Social Computing (Minneapolis, MN, USA) (CSCW ’23 Companion). Association for Computing Machinery, New York, NY, USA, 384–387. https://doi.org/10.1145/3584931.3607492
- Simonyan et al. (2013) Karen Simonyan, Andrea Vedaldi, and Andrew Zisserman. 2013. Deep inside convolutional networks: Visualising image classification models and saliency maps. arXiv preprint arXiv:1312.6034 (2013).
- Smilkov et al. (2017) Daniel Smilkov, Nikhil Thorat, Been Kim, Fernanda Viégas, and Martin Wattenberg. 2017. Smoothgrad: removing noise by adding noise. arXiv preprint arXiv:1706.03825 (2017).
- Sundararajan et al. (2017) Mukund Sundararajan, Ankur Taly, and Qiqi Yan. 2017. Axiomatic attribution for deep networks. In International conference on machine learning. PMLR, 3319–3328.
- Taesiri et al. (2022) Mohammad Reza Taesiri, Giang Nguyen, and Anh Nguyen. 2022. Visual correspondence-based explanations improve AI robustness and human-AI team accuracy. Advances in Neural Information Processing Systems 35 (2022), 34287–34301.
- Teso et al. (2021) Stefano Teso, Andrea Bontempelli, Fausto Giunchiglia, and Andrea Passerini. 2021. Interactive Label Cleaning with Example-based Explanations. In Advances in Neural Information Processing Systems, A. Beygelzimer, Y. Dauphin, P. Liang, and J. Wortman Vaughan (Eds.). https://openreview.net/forum?id=T6m9bNI7C__
- Turri et al. (2024) V. Turri, K. Morrison, K. Robinson, C. Abidi, J. Forlizzi, A. Perer, and R. Dzombak. 2024. Transparency in the Wild: Navigating Transparency in a Deployed AI System to Broaden Need-Finding Approaches. ACM Conference on Fairness, Accountability, and Transparency (ACM FAccT) (2024).
- von Thienen et al. (2019) Julia P. A. von Thienen, William J. Clancey, and Christoph Meinel. 2019. Theoretical Foundations of Design Thinking. Springer International Publishing, Cham, 13–38. https://doi.org/10.1007/978-3-319-97082-0_2
- Wah et al. (2011) C. Wah, S. Branson, P. Welinder, P. Perona, and S. Belongie. 2011. The Caltech-UCSD Birds-200-2011 Dataset. Technical Report CNS-TR-2011-001. California Institute of Technology.
- Wang et al. (2019) Danding Wang, Qian Yang, Ashraf Abdul, and Brian Y Lim. 2019. Designing theory-driven user-centric explainable AI. In Proceedings of the 2019 CHI conference on human factors in computing systems. 1–15.
- Wang et al. (2024a) Fulton Wang, Julius Adebayo, Sarah Tan, Diego Garcia-Olano, and Narine Kokhlikyan. 2024a. Error discovery by clustering influence embeddings. In Proceedings of the 37th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’23). Curran Associates Inc., Red Hook, NY, USA, Article 1809, 13 pages.
- Wang et al. (2024b) Sheng-Yu Wang, Aaron Hertzmann, Alexei A Efros, Jun-Yan Zhu, and Richard Zhang. 2024b. Data Attribution for Text-to-Image Models by Unlearning Synthesized Images. arXiv preprint arXiv:2406.09408 (2024).
- Weisstein (2004) Eric W Weisstein. 2004. Bonferroni correction. https://mathworld. wolfram. com/ (2004).
- Zheng et al. (2023) Xiaosen Zheng, Tianyu Pang, Chao Du, Jing Jiang, and Min Lin. 2023. Intriguing properties of data attribution on diffusion models. arXiv preprint arXiv:2311.00500 (2023).
Appendix A Inclusion criteria: High-risk application areas
We refer to the definition of high-risk application areas according to Annex III of the European Union’s AI Act (Parliament, 2023). For readability, we include an overview:
AI applications in products that require a specific level of safety:
Toys,
Aviation,
Cars,
Medical devices,
Lifts.
Biometric identification and categorisation of natural persons.
Management and operation of critical infrastructure.
Education and vocational training.
Employment, worker management and access to self-employment.
Access to and enjoyment of essential private services and public services and benefits.
Law enforcement.
Migration, asylum and border control management.
Assistance in legal interpretation and application of the law.
Appendix B Interview study themes
The thematic coding analysis of the survey study (N=10) yielded 29 themes corresponding to six theme areas and four unexpected themes (§ 4.2). In the following, we further elaborate on these themes.
B.1. Theme area: Role of ML systems
The theme area “Role of ML systems” groups three themes that represent what kind of role ML systems take in our participant’s work:
Work assistant: ML systems act as work assistants as they take workload off the participant. Chatbots generally fill the role of a work assistant that “[takes work off of the participant’s hands and makes their work easier]” (P10) and is “available around the clock” (P10). Participants use chatbot systems to improve their writing in English (P1, P11), to search for information where previously they would ”[ask] Google” (P1), or to ideate research ideas (P11). Moreover, P10 trusts their company-internal chatbot enough to redirect simple employee questions to the chatbot.
Decision support: As decision support systems, ML systems deliver information that acts as a basis for decisions taken by end-users, e.g. diagnosis support (P6). This is the main role taken by the ML systems that the developer participants work on (“[We] make predictions about where we should do our [pharmaceutical] studies and our development.” (P8), “try to accommodate the policymakers so that they’re able to decide [where more health services are needed]” (P9)). P3 uses ML systems to identify and explain the contributing factors to product issues: “If we can predict it, we can also have an idea what are the factors mostly creating this phenomenon.”
Automation: In some cases, ML systems are deployed for the automation of some tasks, e.g. autonomous driving (P2) or automated processing of camera recordings (P4).
B.2. Theme area: Workflow with ML systems
The theme area “Workflow with ML systems” groups four themes that emerged in the interviews related to standard or often-occuring workflows:
Collaboration with domain experts: Developers and system users collaborate closely for building and evaluating ML models, where bugs are reported to the developers by the end-users (P2, P3, P6, P4, P9). This collaboration is especially strong if the ML system is developed in-house, and less strong if the ML system is a product the company purchased.
Standard model choices: Developers rarely build their own model architectures but opt for standard ones (“[If] it’s already an established problem, you’re probably not going to do better than an algorithm that’s already been laid out to solve that problem for you.” (P8)).
Data-focused debugging process: The development work in practice is centred around data. Developer participants mention that most often, data artefacts are the reason for bugs in the model (P2, P3, P4, P7, P8, P9).
Data quality checks: Data quality checks are a set part of the data preprocessing pipeline (P2, P3, P4, P8). P2 and P9 reported that they first assess data quality before inspecting the model in debugging. This includes looking for any possible data artefacts, especially missing data.
B.3. Theme area: Role of training data
The theme area “Role of training data” groups five themes stemming from participant’s responses when talking about the role of training data in the interviews:
Most important variable in a model: Model developer participants consistently view training data as the most important variable in a model, e.g. “[we] […] believe that […] the models can only be as good as the […] data that you feed in.” (P4). Hence, the debugging process is focused around data (Theme: Data-focused debugging process).
Interesting but no capacity: This theme applies to system user participants, in particular P6. They mention that they would be curious about training data to understand whether the model learned from high-quality data but “[it] is a luxury that [requires time]”, highlighting the practical constraint of time pressure.
Not interesting: This theme applies to system user participants. They were unaware of which training data may have been used to train the models (P1, P10) and mentioned that this does not matter too much for their work as they cannot change it anyhow.
Increasing training data as a bug fix: A reoccuring theme in our interviews is that training data is flexible – the training set can be expanded. This is often done in practice when encountering a bug, e.g. P2 explained that collecting more data is a common way to overcome model shortcomings in autonomous driving.
Training data artefacts as main bug causes: In line with previous themes highlighting the importance of training data to the development process, participants explicitly mention that most often, training data artefacts are the cause of bad model performance. Hence, data artefacts are also the first thing that developers look our for when encountering a bug.
B.4. Theme area: Challenges in working with ML systems
The theme area “Challenges in working with ML systems” groups six themes related to the main pain points that participants mentioned
Trust calibration: Our findings agree with Kim et al. (2023): It is unclear how much and when a system can be trusted. P1 sometimes finds themselves in a dilemma in which they wish to learn something from the chatbot, but are unable to calibrate their trust in the response due to missing knowledge: “I don’t know everything regarding this topic. I [don’t even] know what he’s replying to me.” (P1).
Validation and Evaluation: P2 mentions that the validation itself is a challenge due to multiple requirements that the ML system should fulfil. In practice, the predictive performance of a model matters, along with other objectives such as inference speed, whether the model is bias-free, and more. In addition, A data-centric aspect of evaluation being challenging is the difficulty of evaluation when labels are absent.
Data quality issues: Data quality issues are a major challenge when working with ML systems. In Table 2, we detail the quality issues mentioned by the pariticipants.
Resource constraints: For developer participants, memory and compute constraints are relevant in the development and evaluation process of their models, especially to P2 and P4 as they work with ML systems on the edge.
Stochastic nature of ML models: The stochastic nature of ML models poses a specific challenge to deep models in the development process and debugging tasks: “[The] same data set, same model, you train multiple times, you can get [different] results.” (P2). It is therefore difficult to understand if a bug was truly fixed or if it is a consequence of randomness.
Misuse due to lacking know-how: P10 mentions inadequate know-how in ML system usage as a challenge: “the employees often don’t manage to ask the chatbot the right questions”.
B.5. Theme area: Use of (feature attribution) XAI in practice
The theme area “Use of (feature attribution) XAI in practice” groups six themes related to how XAI is already being used by the participants.
XAI is not used: Some system user participants were unaware of XAI since explanations are not a part of the ML tools they use (P1, P10, P11).
XAI is little used: This theme is related to the system users. P6 reports that XAI tools they used in radiology images so far (i.e. heatmaps) do not deliver a full answer to the why question, as counterfactual information is missing: “[If] I just get an overall highlight in these basal lung regions and the prediction that is atelectasis, I still don’t know why this is atelectasis but not pleural effusion or consolidation.” Hence, they only use it when the information is useful.
Model development: For developer participants, XAI is most commonly used as a tool for model development (P2, P3, P8). As such, XAI tools (within our participants, these are exclusively feature attribution explanations, i.e., SHAP (Lundberg and Lee, 2017) or visualisation of attention) offer explanations for per-example debugging of e.g. wrong predictions (“we work with hands-on examples that we then […] relate the input features that take a poor or a good prediction [to].” - P8).
Understand real-world phenomena: P3 described the use of XAI as a tool to understand phenomena represented by the ML model: “[Building] the model, the whole purpose is to get some explainability. Because […] we know that [a problem is] happening and predicting doesn’t really add value. But if we can predict it, we can also have an idea what are the factors mostly creating this phenomenon.”
Get customer buy-in: P8 states that XAI tools are useful in getting customer buy-in and convincing the customers of the model’s decision suggestion: “Will [the customers] be happy with [high accuracy]? Probably. Will they be convinced? Will they be 100% confident that that’s the way they want to do it? Probably not. So it’s all about the buy-in that you get, and SHAP definitely helps.”
Sanity check for model reasoning: One particular case of using XAI for model development is as a sanity check for model reasoning (“If it’s more like an ideation project, then I would say that the explainability analysis [is done] to see […] what is guiding the prediction as well as […] how are my predictions on the test set explained? And […] does it align with the actual value that I have?” - P8).
B.6. Theme area: User perspectives on TDA
The theme area “User perspectives on TDA” groups four themes that emerged in the participant’s answers when discussing TDA.
Usefulness of XAI is use-case dependent: Some participants find XAI only useful in certain use-cases and under certain conditions such as time availability, intuitiveness of the explanation, and the possibility of taking action based on the explanation. When explanations are used for communication with business departments, they do not deliver useful information when the underlying model is not high-performing yet (P3, P8).
XAI is useful: While participants voiced certain conditions for the usefulness of XAI, the overall majority agreed that explanations are useful in their work (P1, P2, P3, P9, P11).
TDA complementary to feature attribution: Participants are overall positive towards TDA explanations. They mentioned that TDA could be complementary to feature attribution explanations and provide context, as feature attribution information is rather local (e.g., comparing feature attribution maps between training and test samples that are highly related to each other).
Different notions of interesting training samples: Within our participants, we find that there are several different notions of what an interesting training sample to a model output could be among the participants. We detail the codes in this theme in Table 2 of the main paper.
Appendix C Data generation for the survey study
The second probe of our study presents an interactive mock-up of a model development suite where participants can customise what kind of data-centric information is shown. This data is pre-computed using real models and openly available datasets for an image classification and a tabular data classification use case. In the following, we describe the data generation process.
C.1. Use case: Bird classification app
The setting of the image classification use case is about a company that builds a bird classification app. This bird classification app setting has been used in previous HCXAI studies, e.g. Kim et al. (2023). To generate the data for this use case, we finetune the pretrained ResNet18 (He et al., 2016) checkpoint in PyTorch (Paszke et al., 2019) with the CUB dataset (Wah et al., 2011).
Data preprocessing.
The CUB dataset (Wah et al., 2011) is a multi-class image classification with 200 classes, and 5994 images in training and 5794 images in the test set. While the dataset includes fine-grained annotations, we only utilise the target labels in the data generation process. We subsample CUB to include only 10 classes for our application to ensure that participants can learn and identify the different bird species without prior knowledge. We refer to the selected classes as their species family (e.g. Grasshopper sparrow Sparrow) for brevity. The training set includes 300 images, and the test set 277 images.
Model training.
We finetune a ResNet18 model pretrained on ImageNet (available on the torchvision hub) with the CUB training subset for 10 epochs using the AdamW optimizer, a learning rate of 0.001 with weight decay factor 0.005 for the cross-entropy objective. Since the training process of deep models is stochastic which affects any retraining-based effects (Nguyen et al., 2023), we record the last three model checkpoints for a model ensemble. The final model achieves a test accuracy of 0.57 which is suitable for the model debugging use cases in this work.
C.2. Use case: Loan approvement recommendation app
The setting of the scenario of the tabular data use case is the model development department of a bank that implements a machine learning model to give recommendations for loan approvals of their clients. To build this scenario, we train a logistic regression model on the German credit dataset (Hofmann, 1994) which is openly available on the UCI dataset repository and is licensed by CC BY 4.0.
Data preprocessing.
The German credit dataset (Hofmann, 1994) is a multivariate dataset with 20 features and 1000 entries. It is used for binary classification and includes labels of loan approval and declines. We split it into a train and test set using an 80-20 train-test split. As we wish to display the whole data in the interface mock-up, we subselect the following features for readability:
The status of existing checking account feature describes how much money is in the applicant’s checking account and is a categorical variable. Possible values are: (a) less than $0, (b) less than $200, (c) more than $200, and (d) no checking account.
The duration describes the loan’s duration in months and is a numerical variable.
The credit history describes the applicant’s history with loans from this and other banks as a categorical variable. Possible values are: (a) no credits taken, (b) all credits at this bank paid back duly, (c) existing credits paid back duly till now, (d) delay in paying off in the past, and (e) critical account.
The credit amount describes the size of the loan in $ and is a numerical variable.
The installment rate in percentage of the applicant’s disposable income describes how large the installment payments of the loan are. It is a numerical feature.
The age is a numerical feature and describes the age of the loan applicant.
Model training.
We trained a logistic regression model with stochastic gradient descent using the scikit-learn library in Python (Pedregosa et al., 2011). In detail, the model was trained in 50 epochs with L2 regularisation weighted by and the default adaptive learning rate of the library. Recording three checkpoints along the training trajectory at 30, 40 and 50 epochs results in the model ensemble we use for computing the data. The final model accuracy is , leaving room for improvement and debugging.
C.3. Model error selection
In the survey probe, the participant is asked to analyse a model error, i.e. a common misclassification. We select the errors to show by first understanding which type of misclassification is most common, i.e. which class A is often classified as another class B. To determine the test errors, we classify the test set using the model ensemble, group test samples according to true and predicted labels and select the errors represented by the largest groups. We select three error types for the multi-class problem in the bird classification use case. In the binary tabular classification use case, only two errors are possible. In these error groups, we randomly choose three erroneous samples each, resulting in nine and six erroneous test samples for the bird classification and loan approval recommendation use cases, respectively. The test samples are randomised across participants by drawing from a uniform distribution to prevent bias in the analysis stemming from a specific error or test sample.
C.4. Relevant training sample selection
In the mock-up interface used in the survey probe, we display the effect of different training data-based actions on the model error at hand and the global model behaviour. We precomputed these values using the models presented in the above. We elaborate on the procedure of precomputation below.
Determining the training samples.
The interface shows training samples most relevant to the test sample group based on different actions, i.e. removal, label change and upweighting, measured in different metrics and for different group sizes. To determine which training samples to show, we would have to precompute all possible combinations of action, metric and group size. As this is computationally prohibitive and rigorous correctness is not required for our research questions, we select the relevant training sample groups to the erroneous test samples based on individual relevance. In other words, we first iteratively remove/change the label/upweigh each training sample and retrain the models (details below). Then, we record the differences in loss/true class probability/predicted class probability for each erroneous test sample example and compute mean, variance and p-values as in (Nguyen et al., 2023). After filtering the training samples that exhibit a statistically significant effect (), we sort the training sample means from most positive to most negative effect which differs across metrics (e.g. an increase in loss is negative, while an increase in true class probability is positive). The top 10 training samples in each setting are the relevant training samples we use in the interface.
Action implementation.
Implementing and computing the actions requires elaboration:
Removal:The implementation of removal is rather straightforward – we exclude the training sample from the dataset and retrain using the same training pipeline.
Label change: Changing the label of a training sample makes sense when another label may fit the sample better. Hence, we flip the label in the case of binary classification. In the case of more than two classes, we consider two cases: (1) The training sample was misclassified. Assuming that the model learns all classes well, a misclassification could occur because the predicted label describes the data better. Therefore, in these cases, the new label is the predicted label. (2) The training sample was correctly classified. We would only want to change the label if it was wrongly labelled. Assuming that the model generally learned the classes well, we changed the label to the predicted label with the second highest probability.
Upweighting: We implement upweighting with a weight factor of 10.
Computing final effects.
In the interface, we provide simulated effects of certain data-centric actions on the model errors (e.g. removing certain training datapoints leads to an increase in loss). Individual effects do not necessarily add up to the group effect (Basu et al., 2020; Koh et al., 2019). Hence, we compute the final effects using the actual removal/label change/upweighting of training data groups. This computation is feasible since the groups and the number of times we have to retrain the models are limited. We compute and show the standard deviation of removing/changing the label/upweighting across the model checkpoints recorded.
Appendix D Survey study themes
The thematic coding analysis of the survey study (N=31) yielded 34 themes corresponding to four theme areas and four unexpected themes (§ 4.5). In the following, we further elaborate on these themes.
D.1. Theme area: Action depends on user’s hypothesis
The theme area “Action depends on user’s hypothesis” groups 21 themes linked to the reasons for participant’s choices in the action axis. We find that the proposed debugging actions of participants are highly dependent on their hypotheses of the model error’s root cause.
[Hypothesis] Mislabeled data: The model makes an error because of mislabeled data in the training set that is highly related to the erroneous test sample.
[Hypothesis] Incomplete training: The model makes an error because it has not finished training, or was trained using non-optimal hyperparameters and training procedures.
[Hypothesis] Spurious correlation: The model makes an error due to a spurious correlation being present in the test image, that the model learnt from the training data.
[Hypothesis] Similar class features: The model makes an error because the erroneous test sample is similar in features to another class in the dataset.
[Hypothesis] Low dataset diversity: The model makes an error because it was not able to generalise from training on a dataset with low diversity in the samples (i.e., samples of a class are all similar, not showing the breadth of the class).
[Hypothesis] Class overlap: The model makes an error because the class of the erroneous test sample is highly overlapping in terms of class-relevant features with another class and distinctive features may not be present in the data.
[Hypothesis] Dataset imbalance: The model makes an error because it was not able to generalise from an imbalanced dataset (i.e., class imbalance).
[Hypothesis] Small dataset size: The model makes an error because it was not able to generalise from a small dataset, instead it overfit it.
[Action] Remove noisy data: To overcome the error, noisy data such as mislabeled data, is removed from the dataset and the model is retrained.
[Action] Changing label to correct: To overcome the error, incorrectly labelled training data is corrected and the model is retrained.
[Action] Fix labelling pipeline for future data: To overcome the error, the labelling and annotation pipeline is inspected to find a root cause for labelling errors.
[Action] Training process engineering: To overcome the error, hyperparameters and training process protocols are adjusted and tuned before retraining.
[Action] Model architecture engineering: To overcome the error, the model architecture is adjusted (e.g. addition of another layer, choosing a different foundational model).
[Action] Feature engineering: To overcome the error, the feature extractors of the data preprocessing pipeline are inspected. This theme was mainly found for the tabular data use case.
[Action] Upweight to focus on specific features: To overcome the error, certain training data samples should be upweighted so that the model learns the features present in these samples better.
[Action] Feature attribution inspection: To overcome the error, the misleading features must be understood first. Feature attribution could be a useful tool.
[Action] Manual inspection: To overcome the error, the training data is manually inspected to find any data artefacts. Participants mentioned that they already do this as part of their model development pipeline.
[Action] Collect specific new data: To overcome the error, new data related to the erroneous training sample should be collected and added to the training dataset. This can improve dataset diversity and help the model generalise better.
[Action] Synthesise/augment data: To overcome the error, new data samples should be synthesised or existing data samples should be augmented further, targeting the parts of training data that are underrepresented. This theme was mentioned often under the assumption that adding completely new data is not possible.
[Action] Resample/reweight data: To overcome the error, the existing training dataset should be resampled, potentially with a specific weighting to overcome class imbalances.
[Action] Collect any new data: To overcome the error, the model needs to learn from more data. Hence, expanding the dataset overall will likely help the model performance.
D.2. Theme area: Individual metric preferences
The theme area “Individual metric preferences” groups four themes and provides insight into why certain metrics are preferred.
[Action/Metric] Label change and probability shows decision boundary: Probabilities are the preferred metric, especially in combination with the label change action, as it gives the user an intuition about the decision boundary between the predicted and ground-truth class.
[Metric] Probabilities are intuitive and show decision boundaries: Probabilities are the preferred metric because they are an intuitive metric bounded between 0-100. Also, they give the user an intuition about the decision boundary between the predicted and ground-truth class.
[Metric] Ground truth probability is the target: The probability of the ground truth class of the test sample is preferred because this is the target metric to optimise when debugging this erroneous test sample.
[Metric] Loss is most informative in the multi-class case: The loss is preferred because it entails information across many classes and is a direct indicator of how well the model is performing overall. This is particularly interesting in multi-class classification cases.
D.3. Theme area: Reliability of explanations
The theme area “Reliability of explanations” groups three themes related to the overarching need for reliability of the attribution scores.
[Action] Removal shows causal effect: The removal action is preferred because represents the causal effect of the inclusion of a specific training sample. This causal effect is assumed to be meaningful and therefore reliable.
[Number of samples] Group size enough for reliable attribution: A group (i.e., more than one training sample) is preferred for training data attribution because considering the effect of a group is more reliable than of single training samples. The group size should be sufficiently large for the attribution score to be meaningful and hence, reliable.
[Number of samples] Small group to prevent reasoning change: While a group of training samples is preferred, the size of the group should be sufficiently small so that the underlying model stays rather consistent given the proposed action (e.g., if the attribution corresponds to the effect of removing samples, this will change the data distribution which affects the model’s learning, too). If the underlying model were to change too much, the attribution scores could be arbitrary and it is not certain whether the model behaviour will stay similar, making the sample-wise debugging process difficult.
D.4. Theme area: Group attribution preferences
The theme area “Group attribution preferences” groups two themes that elaborate on desired sizes of training data groups as explanations. In contrast to the other axes of action and metric, there is a clear preference towards training data group attribution, emphasising a need for methods that serve this task.
[Number of samples] Group for an overview of the error: A group of training samples is preferred in the explanation of the model error to get a better overview of the error, as it is expected that the samples of the group with the highest attribution are related to the model error. One sample is too little for the user to understand the error and propose an action forward based on the explanation.
[Number of samples] Balance between variety and conciseness: While a group of training samples is preferred, the group size should be large enough to show a variety of different training samples related to the model error, while being small enough to be concise and not overwhelm the user.
D.5. Unexpected themes
In addition to the previous theme areas, we found three additional themes which were unexpected. They provide context to the overall analysis, show two predominant mindsets of developers and point to the research direction of grouping erroneous test samples, which has largely been overlooked.
Focus on the model error at hand: The main focus of the user during the debugging process lies in understanding the model error at hand, and identifying a way to fix the error. This is a local focus.
Focus on the overall model performance: The main focus of the user during the debugging process is to optimise the overall model performance, i.e., having the overall model accuracy in mind. This is a global focus.
Grouping erroneous test samples: In the debugging process, users seek to understand the reasons behind the model error. A single test sample is unlikely to be the only instance of erroneous model behaviour, e.g. spurious correlations. This theme describes the need for presenting the user with multiple test instances that the model errs on similarly so that the user will have a better understanding of the error which will enable them to fix it.
Study interface unfamiliar: In the comments, we occasionally found remarks that the participants were unsure about whether they had correctly fulfilled the task, indicating the overall infamiliarity with sample-wise debugging with TDA explanations.

### v2

@@
 Abstract.
-While Explainable AI (XAI) aims to make AI understandable and useful to humans, it has been criticised for relying too much on formalism and solutionism, focusing more on mathematical soundness than user needs.
-We propose an alternative to this bottom-up approach inspired by design thinking: the XAI research community should adopt a top-down, user-focused perspective to ensure user relevance.
-We illustrate this with a relatively young subfield of XAI, Training Data Attribution (TDA).
-With the surge in TDA research and growing competition, the field risks repeating the same patterns of solutionism.
-We conducted a needfinding study with a diverse group of AI practitioners to identify potential user needs related to TDA.
-Through interviews (N=10) and a systematic survey (N=31), we uncovered new TDA tasks that are currently largely overlooked.
-We invite the TDA and XAI communities to consider these novel tasks and improve the user relevance of their research outcomes.
+Explainable AI (XAI) aims to make AI systems more transparent, yet many practices emphasise mathematical rigour over practical user needs.
+We propose an alternative to this model-centric approach by following a design thinking process for the emerging XAI field of training data attribution (TDA), which risks repeating solutionist patterns seen in other subfields.
+However, because TDA is in its early stages, there is a valuable opportunity to shape its direction through user-centred practices.
+We engage directly with machine learning developers via a needfinding interview study (N=6) and a scenario-based interactive user study (N=31) to ground explanations in real workflows.
+Our exploration of the TDA design space reveals novel tasks for data-centric explanations useful to developers, such as grouping training samples behind specific model behaviours or identifying undersampled data.
+We invite the TDA, XAI, and HCI communities to engage with these tasks to strengthen their research’s practical relevance and human impact.
 1. Introduction
-Understanding AI model behaviour is critical for using AI models in practice, especially in high-stakes domains such as medicine, law, or finance (Parliament, 2023).
-End-users require the understanding of model behaviours to make informed decisions, while developers need this to build reliable models with limited biases (Dwivedi et al., 2023).
-Explainable AI (XAI) aims to provide insights into AI systems’ functioning for humans.
-It has produced a variety of explanation methods that address different model architectures and data modalities (Guidotti et al., 2018; Dwivedi et al., 2023).
-Despite the progress, the XAI community has faced criticism for its predominant techno-centric focus on solutions as opposed to the problem of helping users understand model behaviour (Ehsan et al., 2021).
-Historically, XAI research progressed bottom-up: starting with methods providing qualitative observations, then structuring the field with quantitative evaluations, and eventually considering application in real-world environments including users (see Figure 1).
-Consequently, XAI was mainly driven by “[solutionism (seeking technical solutions) and] formalism (seeking abstract, mathematical solutions)” (Ehsan et al., 2021, p.2).
-An example of this bottom-up approach is the development of the subfield of feature attribution.
-Feature attribution explanations quantify each input feature’s contribution to the model prediction, highlighting key inputs.
-One of the first feature attribution methods in the context of deep learning was presented in 2013 introducing the visualisation of neural network input gradients as saliency map explanations and providing 15 qualitative examples as evidence that saliency maps explain (Simonyan et al., 2013).
-Several other methods followed, either building on this work and proposing changes to satisfy specific formal axioms (e.g., (Smilkov et al., 2017; Sundararajan et al., 2017)), or proposing model-agnostic methods in a quest for a one-fits-all solution (e.g., (Lundberg and Lee, 2017; Ribeiro et al., 2016)).
-Around 2017, XAI reached an inflection point where researchers started to re-assess the practicality of explanation methods and proposed standard evaluation protocols to measure how well explanation methods achieve their goal of explaining (Doshi-Velez and Kim, 2017; Adebayo et al., 2018; Bilodeau et al., 2024).
-Subsequently, recommendations from the social science view of explanations emphasise that explanations are part of a social process and shifted the focus of the field in 2019 (Miller, 2019).
-XAI was criticised for its strong solutionism with an increasing amount of works studying XAI as a sociotechnical system focusing on the user (Wang et al., 2019; Ehsan et al., 2021; Brennen, 2020; Rutjes et al., 2019; Kim et al., 2022, 2023; Fok and Weld, 2023; Kaur et al., 2020; Mei et al., 2023; Ehsan et al., 2024; Lakkaraju et al., 2022) and naming this subfield human-centred XAI (HCXAI) in 2020 (Ehsan and Riedl, 2020).
-Schematic showing the bottom-up and the proposed top-down research approaches.
-There are three boxes vertically stacked.
-The bottom box reads “Methods: Studies on new algorithms.
-XAI example: Saliency maps (Simonyan et al., 2013)”.
-The middle box reads “Evaluations: Studies on protocols for comparing methods.
-XAI example: Sanity checks for saliency maps (Adebayo et al., 2017)”.
-The top box reads “Applications: Studies on methods in user-facing contexts.
-XAI example: Human-centred explainable AI (Ehsan and Riedl, 2020)”.
-There is an arrow pointing upwards to the left of the boxes, annotated with “Bottom-Up Approach -¿ Driven by techno-centric solutions” and an arrow pointing downwards to the right of the boxes, annotated with “Proposed Top-Down Approach -¿ Driven by user needs”.
-Despite the efforts to rectify the solutionism and formalism in XAI communities, we observe repeating patterns in sub-fields of XAI.
-For example, such patterns are emerging in the relatively young field of training data attribution explanations (TDA), where the model behaviour is explained based on the training data (Hammoudeh and Lowd, 2024).
-TDA was introduced to understand deep learning models in 2017 (Koh and Liang, 2017) and has produced several technical approaches to provide TDA explanations since then (Charpiat et al., 2019; Guo et al., 2021; Pruthi et al., 2020; Schioppa et al., 2022; Park et al., 2023; Bae et al., 2024).
-This development demonstrates a strong focus on seeking technical solutions in the TDA field similar to how feature attribution research started.
-Currently, TDA research is gaining momentum due to studies showing that TDA can be applied to address issues like data valuation, debiasing, memorisation and copyright in foundational models (Feldman and Zhang, 2020; Grosse et al., 2023; Choe et al., 2024; Jain et al., 2024; Zheng et al., 2023; Kwon et al., 2023; Lin et al., 2024; Wang et al., 2024b).
-Yet, the current TDA landscapes shows solutionism and formalism, with few user-focused studies, leading to a possible rift between TDA methods and user needs.
-In this work, we propose future research directions for the TDA community that are entirely sourced and motivated by the potential users of TDA.
-We strongly believe that the actual use cases and user needs should inform and guide XAI research, echoing the call for HCXAI (Ehsan and Riedl, 2020).
+Explaining and understanding AI model behaviour is critical for the use and development of AI models in practice, especially in high-stakes domains such as medicine, law, or finance (Parliament, 2023).
+However, despite significant progress in Explainable AI (XAI) techniques (Guidotti et al., 2018; Dwivedi et al., 2023; Nauta et al., 2023), the community has faced criticism for its predominant techno-centric focus on solutions as opposed to the sociotechnical problem of helping users understand model behaviour (Ehsan et al., 2021; Williams, 2021; Wolf, 2019).
+We highlight that historically, XAI research has followed a bottom-up trajectory where methods and frameworks have built on each other: beginning with technical methods that are seemingly useful, then justifying practicality through quantitative evaluation frameworks, and ultimately extending to real-world applications through field studies and user studies (Figure 1).
+An example of this bottom-up trajectory is the development of the subfield of feature attribution.
+Feature attribution explanations quantify each input feature’s contribution to the model prediction, identifying the most influential inputs.
+One of the first feature attribution methods in the context of deep learning was presented in 2013, introducing the visualisation of neural network input gradients as saliency map explanations (Simonyan et al., 2013).
+The paper provides 15 qualitative examples as evidence that saliency maps explain.
+Several other methods followed, either proposing changes to satisfy specific formal axioms (e.g., (Smilkov et al., 2017; Sundararajan et al., 2017)) or proposing more widely applicable model-agnostic methods (e.g., (Lundberg and Lee, 2017; Ribeiro et al., 2016)).
+Around 2017, XAI reached an inflection point where researchers began reassessing the practicality of explanation methods and introduced standardized evaluation protocols to systematically measure their effectiveness in achieving their explanatory goals (Doshi-Velez and Kim, 2017; Adebayo et al., 2018; Bilodeau et al., 2024).
+Soon after, recommendations from the social science view of explanations emphasised that explanations are part of a social process, which shifted the field’s focus to explanation effectiveness in real-world decision-making (Miller, 2019).
+Increasing amount of works studying XAI started to focus on the users and XAI as part of a sociotechnical system (Wang et al., 2019; Ehsan et al., 2021; Brennen, 2020; Rutjes et al., 2019; Kim et al., 2022, 2023; Fok and Weld, 2023; Kaur et al., 2020; Mei et al., 2023; Ehsan et al., 2024; Lakkaraju et al., 2022), some criticising the “[solutionism (seeking technical solutions) and] formalism (seeking abstract, mathematical solutions) (Ehsan et al., 2021, p.1)” of early approaches.
+In 2020, this subfield was coined as human-centred XAI (HCXAI) (Ehsan and Riedl, 2020).
+Despite the efforts to rectify the solutionism and formalism in XAI communities, we observe repeating patterns in the relatively young field called training data attribution (TDA) explanations, where the model behaviour is explained based on the influence of the training data (Hammoudeh and Lowd, 2024) (Figure 2).
+TDA was first introduced to understand deep learning models in 2017 (Koh and Liang, 2017), and has produced several technical explanation approaches since then (Charpiat et al., 2019; Guo et al., 2021; Pruthi et al., 2020; Schioppa et al., 2022; Park et al., 2023; Bae et al., 2024; Ilyas and Engstrom, 2025).
+The field’s development demonstrates a strong focus on seeking technical solutions similar to the trajectory of feature attribution research.
+We argue that TDA research is at a critical juncture, gaining momentum for its ability to address key issues such as data valuation (Ghorbani and Zou, 2019; Choe et al., 2024), debiasing (Jain et al., 2024), memorisation and copyright in foundational models (Feldman and Zhang, 2020; Zheng et al., 2023; Lin et al., 2024; Akyürek et al., 2022) and understanding model behaviour and errors (Grosse et al., 2023; Wang et al., 2024).
+Yet, it remains dominated by solutionism and formalism, with limited user-focused studies, potentially creating a rift between TDA methods and users’ practical needs.
+In this study, we bring users to the centre of our investigation on TDA as explanations, shifting the focus from a technology-centric perspective to a user-centric one.
+This approach enables researchers to develop practical solutions with better usability.
+We propose future research directions for the TDA and HCI community that are entirely sourced and motivated by potential users of TDA.
+To guide a systematic exploration of user needs, we focus on users who traditionally have access and agency over the training data, model developers, as TDA explains model behaviour by identifying how each training sample is relevant to a model outcome.
+This study thus aims at contributing to the study of HCXAI (Ehsan and Riedl, 2020) and, more specifically, human-centred model development tools (Kaur et al., 2020; Cabrera et al., 2023).
 We break down our research question into three parts:
-RQ1: What do users need TDA explanations for?
-RQ2: What do users need from TDA explanations?
-RQ3: To what extent do existing TDA approaches align with user needs?
-We conduct a two-stage needfinding study with potential users of TDA technology to answer our research questions.
-Through a qualitative interview study with AI practitioners (N=10), we identify model developers as a key user group for TDA explanations as opposed to model users.
-We derive a design space for TDA explanations representing different TDA tasks (e.g., identifying the training samples with the largest effect on model behaviour when removed in contrast to when given a larger weight during training).
-We study which tasks are needed in model development in a subsequent mixed-methods survey study (N=31) with model developers.
-By considering the user and their usage scenarios, we identify tasks relevant to users but largely overlooked by current TDA research.
-We invite the research community to address these gaps and shift the focus of TDA research from solutionism to user-serving questions, i.e., a top-down approach.
+RQ1: What data-related explanation needs, both explicit and latent, do model developers have?
+RQ2: To what extent do existing TDA approaches align with model developers’ needs?
+RQ3: What kind of information produced by TDA will be useful for model developers?
+We conducted a two-stage needfinding study with model developers to address our research questions.
+Through a qualitative interview study (N=6), we explore potential usage scenarios of TDA explanations.
+We note that while TDA has the potential to support many data-centric tasks for model developers, it was not yet widely known or used in practice.
+Based on the interview findings, we derive a design space for TDA explanations representing different TDA tasks (e.g., identifying the training samples with the largest effect on model behaviour when removed in contrast to when given a larger weight during training).
+In a subsequent scenario-based interactive user study (N=31), we study which specific data-centric tasks are needed to support testing users’ hypotheses.
+We identify tasks relevant to potential TDA users, but are largely overlooked by current TDA research.
+We invite the TDA and broader XAI research communities to address these gaps and reorient TDA research away from solutionism toward user-driven inquiry – adopting a top-down, user-centred perspective.
 We advocate for adopting this approach in other areas of XAI and AI as a whole to avoid focusing on technically elegant solutions that may overlook user needs.
-Our contributions are:
+To summarise, our work makes the following contributions:
 We propose a top-down approach to XAI research driven by user needs and demonstrate this approach on the emerging TDA subfield.
-Through a needfinding study, we identify potential use cases, where model developers emerged as key users of TDA.
-We find that they require TDA explanations to be (1) flexible and adapt to the user’s intuitions in debugging and (2) reliable and stable.
-Generally, we observe a preference for training data group attribution over individual attribution.
-As a result of studying the user needs, we identify tasks and research directions for user-focused TDA that are currently understudied.
+From the use cases we found, we identified research directions for user-focused TDA that are currently understudied and invite the community to study these (§ 5.2).
+We outline our approach into a high-level framework for needs-based research that connects user-oriented inquiry with method development in the early phases of technology research (§ 5.3).
 2. Related Work
-Our work proposes a top-down approach to XAI research, grounded in user needs, which positions it in human-centered explainable AI (HCXAI).
-We apply this approach to the relatively young subfield of training data attribution (TDA) to provide early insights into users and their needs.
-To this end, we utilise methods from needfinding.
-In the following, we relate our work to HCXAI, TDA and needfinding.
-2.1. Human-centred explainable AI
-The aim of explainable AI (XAI) research is to study methods that explain the behaviour of AI models (Guidotti et al., 2018).
+Our work takes a user-centered approach to XAI research, building from empirical insights on training data attribution (TDA) use cases, which aligns it with human-centered explainable AI (HCXAI).
+To this end, we draw on needfinding methods, a user-centered research approach that helps uncover latent needs and usage scenarios, to inform the design and development of TDA tools.
+The following sections connect our work to existing literature in HCXAI and TDA.
+2.1. HCXAI for Model Development
+The aim of explainable AI (XAI) research is to study methods that explain the decisions and behaviour of AI models commonly referred to as “black boxes” (Guidotti et al., 2018).
 The field has developed various explanation methods over the years to explain different types of tasks and models.
-Yet, explanations are only meaningful in front of an audience.
-This human aspect was often neglected in XAI studies (Rong et al., 2023; Nauta et al., 2023).
-To overcome this gap, Human-centred XAI (HCXAI) emphasises the human in the explanation process and focuses on user perspectives in XAI (Ehsan and Riedl, 2020).
-In HCXAI, explanations of black-box models are studied holistically as sociotechnical systems centred around the user (Ehsan et al., 2024).
+Yet, AI explanations are only meaningful if they are interpretable by the intended users in a way that informs or supports their decision-making.
+This human aspect was often neglected in early XAI studies (Rong et al., 2023; Nauta et al., 2023).
+In the Human-centred XAI (HCXAI) field, explanations of black-box models are studied holistically as sociotechnical systems centred around the human who is interested in the explanation (Ehsan et al., 2024).
 HCXAI explores how to design AI technology and explanations that focus on user understanding and practicality (Ehsan and Riedl, 2020).
@@
 human-AI-collaboration (Kim et al., 2023; Shen et al., 2023; Taesiri et al., 2022; Lakkaraju et al., 2022), design guidelines (Liao et al., 2020; Hadash et al., 2022; Wang et al., 2019), evaluation of human understanding of explanations (Rong et al., 2023; Kim et al., 2022; Hadash et al., 2022; Rutjes et al., 2019) and understanding user needs and requirements (Juneja et al., 2024; Kaur et al., 2020; Brennen, 2020; Jin et al., 2023).
-Our work contributes to the HCXAI field by proposing an early consideration of user needs to inform ongoing research in XAI methods.
-We focus particularly on TDA explanations which have not been explored widely in HCXAI before.
-2.2. Training data attribution
+However, much XAI work has focused on system end-users who are affected by an AI system (Juneja et al., 2024; Jin et al., 2023; Kim et al., 2023; Shen et al., 2023; Taesiri et al., 2022; Lakkaraju et al., 2022) and rather few explored the need for model developer who can modify the model architecture, parameters, or training process (Kaur et al., 2020; Hong et al., 2020).
+Our work contributes to the HCXAI field by proposing an early consideration of user needs, especially the model developers’ needs, to inform ongoing research in XAI methods.
+We focus particularly on TDA explanations for model developers because they have agency over the data, which have not been explored widely in HCXAI before.
+2.2. TDA as Explanation
 Training data attribution (TDA) explains model predictions by pointing to training data samples relevant to the model predictions (Hammoudeh and Lowd, 2024).
 TDA differs from the mainstream XAI approach known as feature attribution (FA) where the model predictions are attributed to the features of an input given to the model, disregarding the impact of training data.
-Interest in TDA has recently accelerated with the paradigm shift towards data-centric AI (DCAI) (Jin et al., 2024), where the importance of using the right training data is emphasised over other factors like model architecture and optimisation algorithms.
-Under the context of DCAI, TDA has been utilised to enhance data quality by cleaning faulty data labels (Teso et al., 2021) and detecting model biases (Pezeshkpour et al., 2022; Brunet et al., 2019; Wang et al., 2024a; Jain et al., 2024).
+Interest in TDA has recently accelerated with the paradigm shift towards data-centric AI (DCAI) (Jin et al., 2024), where the importance of using the right training data is emphasised over other factors like model architecture and optimisation algorithms (Sambasivan et al., 2021).
+Under the context of DCAI, TDA has been used to enhance data quality by cleaning faulty data labels (Teso et al., 2021) and detecting model biases (Pezeshkpour et al., 2022; Brunet et al., 2019; Wang et al., 2024; Jain et al., 2024), as well as answering questions around data valuation or memorisation and copyright issues in foundational models in recent years (Choe et al., 2024; Zheng et al., 2023; Grosse et al., 2023; Lin et al., 2024).
+As such, TDA offers the potential of providing actionable explanations for users with dataset access, like model developers.
 The recent rise of foundational models, such as large language models (LLMs) and diffusion-based image-generation models, further contributed to the interest in TDA, as it has proven effective for studying the inner mechanisms as well as understanding the associated copyright and privacy leakage issues therein (Feldman and Zhang, 2020; Choe et al., 2024; Zheng et al., 2023; Grosse et al., 2023; Lin et al., 2024).
-With the accelerated growth, we observe again a general focus on technical and mathematical solutions in the TDA community (Koh and Liang, 2017; Charpiat et al., 2019; Guo et al., 2021; Schioppa et al., 2022; Pruthi et al., 2020; Park et al., 2023; Bae et al., 2024; Choe et al., 2024; Kwon et al., 2023), without a deep understanding of user needs, even in studies applying TDA to a downstream tasks (Feldman and Zhang, 2020; Jain et al., 2024; Zheng et al., 2023; Akyürek et al., 2022).
+With the accelerated growth, we observe again a general focus on technical and mathematical solutions in the TDA community (Koh and Liang, 2017; Charpiat et al., 2019; Guo et al., 2021; Schioppa et al., 2022; Pruthi et al., 2020; Park et al., 2023; Bae et al., 2024; Choe et al., 2024; Kwon et al., 2023; Ilyas et al., 2022; Wang et al., 2025a, b; Ilyas and Engstrom, 2025), without a deep understanding of user needs, even in studies applying TDA to a downstream tasks (Feldman and Zhang, 2020; Jain et al., 2024; Zheng et al., 2023; Akyürek et al., 2022; Li et al., 2024).
 Our work addresses this gap by studying practical needs and providing concrete recommendations for the field.
-2.3. Needfinding
-Needfinding is an exploratory and qualitative approach to user research to understand users’ needs beyond what simply asking users may tell (Patnaik and Becker, 1999).
-It aims at studying the true user needs, regardless of solutions, which may not be apparent to the users themselves.
-Needfinding stems from the field of design thinking (von Thienen et al., 2019), which is an iterative design approach driven by user needs.
-Needfinding is an iterative process involving users where the subsequent studies extract greater information about their problems and needs through field studies, interviews, and lab studies.
-In XAI, needfinding has been used to study general needs such as user expectations for XAI (Brennen, 2020) or end-user transparency needs in AI decision-support systems (Turri et al., 2024; Lin et al., 2024) by using semi-structured interviews, scenario-based study designs and qualitative analyses of case-study related text files.
-Needfinding has also been applied to more narrow questions, like extracting user needs for web-search explanations through surveys and semi-structured interviews (Juneja et al., 2024) or understanding explainability needs in human-AI collaborations (Kim et al., 2023).
-Our work extends this overall research direction by studying the user needs of TDA explanations through a two-stage study involving semi-structured interviews and a systematic survey study.
-3. Training data attribution
-Our work studies user needs for training data attribution (TDA) explanations.
-This section provides background on TDA.
-TDA is links specific model behaviour to the training data, treating the data as the root cause for learned behaviours (Hammoudeh and Lowd, 2024).
-By providing training data relevant to what the model has learnt, TDA gives insights into the model.
-TDA has been applied to enhancing data quality and reducing model biases, as well as answering questions around data valuation, memorisation and copyright issues in foundational models in the recent years (Choe et al., 2024; Zheng et al., 2023; Grosse et al., 2023; Lin et al., 2024), attracting unprecedented attention.
+3. Background: What is TDA?
+This section provides background on training data attribution (TDA), describing what it is and why it is an interesting explanation approach.
+In theory, TDA links specific model behaviour to the training data, treating the data as the root cause of learned behaviours (Hammoudeh and Lowd, 2024).
+TDA offers insights into the model by identifying the training data most influential to what the model has learned.
 Formal definition.
-The TDA community has focused on a particular instance of attributing a model prediction on a single test sample on a single training sample.
-More formally, let us consider a training sample and a test sample , where indicates the input and indicates the true answer the model is supposed to predict.
-We define the attribution score for and as the change in the correctness of the model prediction on , measured via the model loss , before and after removing the sample from the training procedure (Hampel, 1974):
+Let us consider a training sample and a test sample , where indicates the input and indicates the true answer the model is supposed to predict.
+An attribution method assigns a score to a training data point , as the change in the correctness of the model prediction on , indicating its importance measured via the model loss , before and after removing the sample from the training procedure (Hampel, 1974):
 Here, we indicate the model trained with all training samples as and the one trained without as .
 Main focus of TDA research.
-Computing directly by re-training a model without each training sample is computationally prohibitive.
-To identify the most influential training sample, the re-training has to be repeated as many times as the number of training samples.
-The main focus of the TDA community has been on efficient approximations of Equation 1.
-In 2017, Koh & Liang (Koh and Liang, 2017) introduced a gradient-based approximation called influence functions (IF).
+Much of the research in the TDA community has centred on efficiently approximating Equation 1 because computing directly by re-training a model without each training sample is computationally prohibitive.
+To identify the most influential training sample, the re-training has to be repeated as often as the number of training samples.
+In 2017, Koh & Liang (Koh and Liang, 2017) introduced a gradient-based approximation called influence functions (IF) for deep models.
 Since IF was still computationally prohibitive due to the need to compute and invert a massive Hessian matrix, subsequent works in 2020 and 2021 have focused on speeding up the algorithm by trading off precision and computational costs (Pruthi et al., 2020; Charpiat et al., 2019; Schioppa et al., 2022; Guo et al., 2021).
-After the advent of ChatGPT in late 2022 (OpenAI, 2022) and the boom in large language models in 2023 (Naveed et al., 2023), the community has developed further enhancements to apply the TDA methods on billion-scale LLMs (Bae et al., 2024; Choe et al., 2024; Grosse et al., 2023).
+After the advent of ChatGPT in late 2022 (OpenAI, 2022) and the boom in large language models in 2023 (Naveed et al., 2023), the community has developed further enhancements to apply the TDA methods on billion-scale LLMs (Bae et al., 2024; Choe et al., 2024; Grosse et al., 2023; Park et al., 2023).
+For a survey on TDA methods, we refer the reader to Hammoudeh and Lowd (2024).
 Critique.
 We observe hints of solutionism and formalism in the development of TDA research, as the majority of the community effort is dedicated to addressing computational inefficiencies of the method itself, without questioning the practical relevance and user values of TDA as defined in Equation 1.
-This particular formulation of TDA has not gone without criticism within the community: a prior work (Nguyen et al., 2023) discussed the statistical insignificance of removing a single training sample in modern neural network training and (Ilyas et al., 2022) recognised the limitation and argued for examining the removal of multiple training samples.
-Some TDA works have explored possible applications that are potentially interesting to users such as fixing mislabelled data (Koh and Liang, 2017) or identifying brittle classes (Ilyas et al., 2022).
-However, these applications are still not driven by user needs.
-There is no known work in the TDA domain involving actual users to identify their needs to assess the relevance of the current TDA tasks and identify research directions grounded in user needs that the community should focus on.
-4. Understanding user needs of TDA explanations through needfinding study
-Our work studies user needs of training data attribution (TDA) explanations to facilitate user need-driven TDA research.
-To this end, we conduct a two-stage needfinding study.
-First, we identify use cases for TDA explanations using semi-structured interviews.
-The experimental setup is described in § 4.1.
-§ 4.2 presents the interview findings.
-We derive a design space representing different types of information that may be needed from TDA in § 4.3.
-Second, we conduct a systematic survey study targeted at extracting user needs from this space (§ 4.4).
+This particular formulation of TDA has not gone without criticism within the community: a prior work (Nguyen et al., 2023) discussed the statistical insignificance of removing a single training sample in modern neural network training which affects the fragility of approximation methods (Basu et al., 2021; Bae et al., 2022) and (Ilyas et al., 2022) recognised the limitation and argued for examining the removal of multiple training samples.
+Some TDA works have explored possible applications that are potentially interesting to developers such as fixing mislabelled data (Koh and Liang, 2017) or identifying brittle classes (Ilyas et al., 2022).
+However, these applications are not driven by user needs and rather targeted toward the evaluation of TDA methods, potentially missing application scenarios and not addressing user needs.
+Potential.
+TDA provides data-centric information that users can make use of to gain a deeper understanding of their training datasets.
+Such information may be particularly interesting if the user has agency over the training dataset, meaning that they are in a position to adjust it.
+Hence, TDA can be especially useful to model developers when a high-quality training dataset is of crucial importance (Sambasivan et al., 2021).
+Developers often make sense of their models through manually inspecting training data (Cabrera et al., 2023).
+There are tools that enable model developers to explore datasets by offering a visual interface (Hohman et al., 2020; Piorkowski et al., 2023), but they often do not inform about the connection between model output and training data.
+As a data-centric approach to gaining such insight, TDA could support and enhance model developers’ data-centric work.
+Currently, there is no known work in the TDA domain involving actual users to identify their needs to assess the relevance of the current TDA tasks and identify research directions grounded in user needs that the community should focus on.
+4. Needfinding: Asking Model Developers What They Really Need
+Our work investigates the user needs surrounding training data attribution (TDA) explanations to support a user-centered development for machine learning (ML) developers and practitioners.
+To this end, we conducted a two-stage needfinding study: first, we ran an interview study with ML practitioners (N=6) to identify the design space of TDA explanations that represent useful data-centric information for model development.
+Second, we build on this space through a scenario-based interactive user study (N=31) aimed at eliciting concrete user needs and expectations.
 IRB approval was obtained from one of the authors’ institutions.
-4.1. Interview study
-In the interview study, we conduct semi-structured interviews with users of machine learning (ML) applications outside of ML research.
-This serves as an exploratory study to gain an impression of potential usage scenarios and user needs for TDA RQ1 – What do users need TDA explanations for?
+4.1. Interview Study Procedure
+We conducted semi-structured interviews with machine learning (ML) developers who work outside academic research.
+The goal of the interview was to explore developers’ current and anticipated practices around training data to surface latent needs that may inform the design of TDA explanation tools.
 4.1.1. Participants
-To get a realistic impression of user needs for TDA explanations in practice, we target participants who work with ML applications outside of academia.
-Our inclusion criteria are: Participants should (1) have at least one month of experience in working with ML systems and (2) work in a high-risk application area according to the EU AI Act (Parliament, 2023) (e.g., health care, law enforcement, complete list in Appendix A).
-This criterion serves to find users who are likely to use explanations, as these areas are subject to further regulations (Guidotti et al., 2018; Doshi-Velez and Kim, 2017).
+To get a realistic impression of user needs for TDA explanations in practice, we target participants who develop ML applications outside of academia.
+Our inclusion criteria were: Participants should (1) have at least one year of experience in developing ML systems and (2) work in a high-risk application area according to the EU AI Act (Parliament, 2023) (e.g., health care, cars, law enforcement.
+Complete list in Appendix A).
+The first criterion ensures that participants have hands-on experience in developing machine learning models which their statements can be rooted in.
+The latter criterion serves to find participants who are likely to use explanations, as these areas are subject to further regulations (Guidotti et al., 2018; Doshi-Velez and Kim, 2017).
 Recruiting participants poses a challenge, especially in high-risk application areas.
-Hence, we use purposive sampling (Guest et al., 2023) and approach participants from the authors’ network.
-We recruit 10 participants from various domains and degrees of experience (see Table 1).
-| ID | Location | Domain | Type | Job experience / with ML | Type of ML |
-| P1 | DE | HR | User | 3 years / 1 months | Chatbot |
-| P2 | US | AV | Developer | 2 years / 7 years | Prediction model |
-| P3 | NL | TC | Developer | 3 years / 5 years | Prediction model |
-| P4 | FI | CV | Developer | 4 years / 6 years | Prediction model |
-| P6 | CH | Health | User | 2 years / 2 years | Prediction model |
-| P7 | NL | Health | Developer | 1 year / 3 years | Prediction model |
-| P8 | BE | Health | Developer | 2 years / 6 years | Prediction model |
-| P9 | PK | Health | Developer | 5 years / 2 years | Prediction model |
-| P10 | DE | HR | User | 3 years / 1 year | Chatbot |
-| P11 | DE | Health | User | 10 years / 6 years | Clustering, Chatbot |
+Hence, we used purposive sampling (Guest et al., 2023) and approached participants from the authors’ network 111As the interview study is of exploratory nature, we stopped recruitment for the interviews after three months.
+We note that potential participants in legal and finance domains were unable to take part in the study due to non-disclosure constraints..
+We recruited six participants from various domains and degrees of experience as summarised in Table 1.
+| ID | Location | Domain | Job experience / with ML |
+| I1 | US | Autonomous vehicles | 2 years / 7 years |
+| I2 | NL | Telecommunications | 3 years / 5 years |
+| I3 | FI | Computer vision for automation | 4 years / 6 years |
+| I4 | NL | Health | 1 year / 3 years |
+| I5 | BE | Health | 2 years / 6 years |
+| I6 | PK | Health | 5 years / 2 years |
 4.1.2. Interview process
-The interviews were conducted during June – September 2023, either in person or remotely through video call using Zoom 111https://zoom.us/.
-All interviews are one-on-one conversations in English, except for P10 in German.
-Participants were first briefed about the purpose of the study and data processing.
+The interviews were conducted in person or remotely through video call using Zoom222https://zoom.us/.
+All interviews were one-on-one conversations.
+The participants were first informed about the purpose of the study and data processing.
 Upon receiving informed consent, we began the interview recording.
-In general, the interviews lasted between 30 and 60 minutes.
-Each interview addresses the following topics to answer RQ1 - What do users need TDA explanations for?:
-Job-related information.
-Perspectives may vary between different domains, levels of seniority and experience with the ML tool.
-This information provides user context and will help in differentiating user groups.
-Interviewee’s workflow with ML systems.
-By asking about the workflow with the ML tool, we wish to understand the patterns of usage and challenges participants encounter.
-We collect this information to identify challenges that could potentially addressed by TDA.
-Perspectives on training data.
-Since we investigate TDA explanations, we explicitly ask participants about the role training data plays in their tasks.
-This indicates what type of data-centric information is relevant to the participant’s work and for which tasks.
+The interviews lasted between 30 and 60 minutes.
+Each interview addressed the following topics:
+Job Roles and Responsibilities.
+Perspectives may vary between different domains, levels of seniority, and experience.
+This information provides context about the interview responses for subsequent analysis.
+ML Model Development Workflow.
+By asking about interviewees’ development workflows, we aim to uncover diverse development styles and the challenges they face, both explicit and latent.
+This helps surface underlying needs that may not be readily articulated, but could inform opportunities for supporting these workflows through TDA tools.
+Uses and Relevance of Training Data.
+We explicitly asked participants about the role that training data plays in their day-to-day tasks.
+Their responses highlight the types of data-centric information they find most relevant, and the specific tasks or decision points where such information becomes critical.
 Perspectives on XAI and TDA.
-We address the participant’s perspectives on XAI and particularly on TDA to understand current explanation usage scenarios and explore participant’s opinions about TDA, a data-centric approach to XAI.
+We address the participant’s perspectives on XAI and particularly on TDA to understand current explanation usage scenarios in model development and explore participant’s opinions about TDA, a data-centric approach to XAI.
 4.1.3. Data analysis
-We analyse the interview data and extract relevant information for answering RQ1.
-The interviews are transcribed automatically using Whisper (Radford et al., 2023) and cleaned up manually by the authors.
-The transcript is then pseudonymised.
-We translated P10’s German transcript to English using DeepL 222https://www.deepl.com/translator.
-We then analyse the transcripts through an inductive thematic analysis (Braun and Clarke, 2012) by two coders, where the codes are directly annotated in the transcript.
-The analysis is iterative: The interview transcript of P1 is first jointly analysed in an initial coding workshop that resulted in an initial set of themes.
-Afterwards, the coders independently code five transcripts, expanding on the themes found in the initial analysis.
-During an intermediate coding workshop, agreements and disagreements between the coder’s themes are discussed.
-The workshop resulted in a merged definition of the themes that are used for the remaining transcripts.
-At the intermediate coding workshop, the interrater agreement is 77.3% measured by the percentage of agreement participants coded to themes.
-The final coding workshop takes place after both coders reviewed the remaining transcripts.
-The final interrater agreement is 80.3%.
-4.2. Interview findings
-The figure depicts the resulting themes from the coding analysis of the interview transcripts.
-There are 28 themes in total, spread over six theme areas.
-The theme areas are larger boxes with rounded edges, encompassing the themes denoted in square boxes (intended to mimic sticky notes).
-Themes related to training data and TDA are marked with a star.
-The themes areas with themes in parentheses are: Role of ML systems (Work assistant, decision support, automation), workflow with ML systems (Collab with domain experts, standard model choices, data-focused debugging process*, data-quality checks*), Role of training data (Most important variable in a model*, interesting but no capacity*, not interesting*, increasing training data as a bug fix*, training data artifacts as main bug causes*), challenges in working with ML systems (trust calibration, resource constraints, misuse due to lacking know-how, validation and evaluation*, data quality issues*, stochastic nature of ML models*), use of feature attribution XAI in practice (XAI is not used, XAI is little used, understand real-world phenomena, get customer buy-in, sanity check for model reasoning, model development*), user perspectives on data-centric XAI (usefulness of XAI is use-case dependent*, XAI is useful*, TDA complementary to feature attribution*, different notions of interesting training samples*).
-The interview study aims to provide insights into the potential usefulness of TDA, answering RQ1 - What do users need TDA explanations for?
-The analysis yields six theme areas, which are common across all participants (see light grey boxes in Figure 2): the role of the ML system as an application, the workflow of participants with ML systems, the role of training data for participant’s work, the main challenges in working with ML systems, the use and role of explanation tools, and perspectives and opinions about data-centric explanations like TDA.
-We elaborate on each area below and provide a full overview in Appendix B:
-4.2.1. Role of ML systems.
-Our participants work with different types of ML models across several domains and in different positions (see Table 1).
-We find that ML systems are used as work assistants (P1,P10,P11), decision support systems (P3,5,6,7,8,9), and for automation in systems like autonomous vehicles or machinery (P2, P4).
-These roles provide context for our analysis, as the needs and usage scenarios of XAI may differ based on the level of user interaction.
-For example, P4 develops automated computer vision models for animal population monitoring with minimal human involvement, while P10 uses a chatbot to handle smaller tasks like answering employee questions.
-Our analysis is set against the backdrop of these use cases.
-4.2.2. Workflow with ML systems.
-As ML systems take different roles, the workflows differ.
-We notice that ML systems are not an essential part of the workflow for system users, as opposed to developers.
-System users only integrate the ML system in their work when it delivers helpful suggestions and report bugs or ignore it otherwise (P6, P10).
-We recognise a separation of domain knowledge (collaboration with domain experts): “Because personally, I cannot know if the model is doing the correct thing […] business have to tell me” (P3).
-For developers whose job revolves around the ML system, we identify themes that are relevant for understanding usage scenarios of XAI and TDA.
-We find that the development workflow is heavily centred around data: “[What] drives your model is your data.
-[…]” (P8) (data-focused debugging process, data quality checks).
-Often, standard model architectures are employed with standard hyperparameters (standard model choices) and the main work lies in data curation (P2, P4, P3, P7, P8, P9).
-This theme area shows that there are two distinct user groups, system users who utilise ML systems and model developers who build ML systems.
-They have different workflows and potential XAI use cases.
-4.2.3. Role of training data.
-The training data plays different roles to system users and model developers.
-While it is the most important variable in a model for developers (P3, P4, P8), training data is generally not interesting to system users (P1, P10).
-System users often do not have capacity to interact much with the ML system to verify model behaviour, even if they an interest in the underlying mechanisms exists, because they are preoccupied with their own tasks (interesting but no capacity) (P6).
-However, for developers, training data is a lever for model performance (increasing training data as a bug fix, training data artifacts as main bug causes) (P2, P4, P3, P7, P8, P9, P11).
-These insights underline that XAI needs to be intuitive, and TDA is particularly suited for model development and less for system users.
-4.2.4. Challenges in working with ML systems.
-In the interviews, we ask participants about challenges in their ML workflow to explore if XAI can address them.
-We identified six themes in this theme area.
-We confirm Kim et al.’s (Kim et al., 2023) observation that system users’ main challenge for working with ML systems is trust calibration (P1, P6, P8, P11).
-In addition, we find that system users fear they may misuse the ML system due to lacking know-how (P1, P10, P11), highlighting the need for user training to make effective use of ML systems in practice.
-For model developers, we identify several challenges: Data quality issues are often the root cause of model malfunction (data quality issues) (P2, P4, P3, P7, P8, P9, P11), in turn impacting model validation and evaluation.
-For instance, participants encounter difficulties due to missing data or absent labels (P2, P4, P3, P7, P8, P9).
-Additionally, resource constraints (P2, P4, P11) and the stochastic nature of ML models (P2) are pain points for the development and evaluation of ML systems.
-Our analysis shows that the challenges faced by system users require sociotechnical consideration, whereas challenges faced by model developers are more technical with data quality and evaluation being most pressing.
-We believe that the latter is a suitable usage scenario for TDA.
-4.2.5. Use of XAI in practice.
-This theme area groups six themes about the use or lack of XAI in practice.
-Our analysis reveals that XAI is usually not a standard part of an ML system in practice (XAI is not used, XAI is little used) (P1, P4, P6, P7, P9, P10, P11).
-In fact, the uses of XAI, in particular feature attribution, among our participants are limited to model development (P2, P3, P8).
-XAI offers explanations for per-example debugging or act as a sanity check for model reasoning (P8, P9).
-Unexpectedly, developer participants also use XAI to understand real-world phenomena modelled by their ML system (P3) and as a communication means to get customer buy-in for their ML systems (P8).
+The interviews were transcribed using Whisper (Radford et al., 2023) and translated using DeepL333https://www.deepl.com/translator if necessary.
+Then they were manually cleaned up and pseudonymised by the authors.
+The transcripts were analyzed through an inductive thematic analysis (Braun and Clarke, 2012) by two coders, where the codes were directly annotated in the transcript.
+The analysis was iterative: One interview transcript from a pilot interview was first jointly analysed in an initial coding workshop that resulted in an initial set of themes.
+Afterwards, the coders independently code three transcripts at a time, expanding on the themes found in the initial analysis.
+During an intermediate coding workshop, agreements and disagreements between the coder’s themes were discussed.
+The workshop resulted in a merged definition of the themes that were used for the remaining transcripts.
+At the intermediate coding workshop, the inter-rater agreement measured by Krippendorff’s (Krippendorff, 2011) was computed with (Marzi et al., 2024).
+The final coding workshop took place after both coders have reviewed the remaining transcripts.
+The final inter-rater agreement was .
+4.2. Interview Study Findings
+The analysis resulted in themes within three key meta-themes: the main challenges in working with ML systems, whether and how explanation tools are used, and perspectives and opinions about data-centric explanations like TDA.
+The meta-themes encompass themes and are depicted in Figure 3.
+We elaborate on each meta-theme below:
+4.2.1. Challenges in ML development
+In the interviews, we asked participants about challenges in their development workflow to explore whether XAI can address them.
+We identified three TDA-related themes under this meta-theme with several challenges: Data quality issues are often the root cause of model malfunction (Theme: data quality) (I1, I3, I4, I5, I6): “the models can only be as good as the […] data that you feed in.” (I3).
+For instance, participants encountered difficulties developing models due to missing data or absent labels (I1, I3, I4, I5, I6).
+Our interviews support previous observations that a significant part of model development is data work (Sambasivan et al., 2021; Jin et al., 2024).
+Besides data quality, our interviews revealed that the evaluation and quantification of ML systems’ performance poses a challenge in the development pipeline as well (Theme: Evaluation).
+In practice, developers need to balance different desiderata and objectives from their businesses and stakeholders.
+For example, the priority of predictive accuracy versus efficiency varies according to the particular business scenario and application.
+However, a large challenge in evaluation is also linked to data quality since evaluation datasets in practice are changing across time and tasks, so that missing data and unreliable data labels are a common issue.
+Also, the stochastic nature of ML models (I1) are pain points for the development and evaluation of ML systems since bug fixes may not work in the way they were intended, requiring further analysis (Theme: Stochasticity).
+Due to this stochasticity, participants perceive additional data collection as a more reliable bug fix: “[If] you have a neural network system, if you think you’ve solved the problem, you’re never 100% guaranteed […] In practice, […] whenever we had an issue […] we would first like get a lot more data, label a lot more data and add those new examples to the training data, train a new model on these.” (I1).
+4.2.2. Use of (feature attribution) XAI
+This meta-theme groups two themes about the use or lack of XAI in practice.
+Our analysis revealed that, in most cases, when developers claimed to have used XAI, they were referring to feature attribution-based techniques (I1, I3, I5) (Theme: XAI-supported development).
+We found that feature attribution-based XAI offers explanations for per-example debugging or acts as a sanity check for model reasoning (I5, I6).
+Participants also use XAI to understand real-world phenomena modelled by their ML system: “The valuable part is that this model that can predict […] can also show us why […] when we use some explainability techniques.” (I2) and as a communication means with their business counterparts or to get customer buy-in for their ML systems: “[We] output [the explanation] to the business basically.
+That’s what they work with.” (I2), “[It’s] all about the buy-in that you get, and SHAP definitely helps.” (I5) (Theme: Reasoning checks).
 While explanations have different purposes, we note that participants use XAI tools mainly as an out-of-the-box functionality (e.g., the SHAP library (Lundberg and Lee, 2017)).
 We find that implementation thresholds must be low for the adoption of XAI in practice.
-4.2.6. User perspectives on TDA
-None of the participants were familiar with TDA, highlighting a gap between research and practice.
-Our analysis, based on discussions about the concept of TDA, identifies four themes reflecting participants’ views on XAI and TDA.
-Overall, participants find XAI useful for debugging and communication, though they noted its usefulness is use-case dependent and not always guaranteed.
-They expect TDA not to be an exception (P3, P6, P8, P11), but are optimistic about its potential, especially for model development.
-For instance, P2 mentioned that TDA could save time by identifying faulty training data.
-One participant suggested combining TDA with feature attribution for debugging (P3).
-We observed that participants have different notions of interesting training samples, with some focused on finding faulty data (P2, P11) and others on identifying samples similar to a specific one (P6, P7, P8).
-In conclusion, we identify that TDA could be valuable for model debugging, leading us to focus on this user group and use case in the remainder of this work.
+4.2.3. Perspectives on TDA
+None of the participants were familiar with TDA, highlighting a gap between research and practice and emphasising that TDA technology is in its early stages.
+Our analysis, based on discussions about the concept of TDA, identifies three themes reflecting participants’ views on XAI and TDA.
+Overall, participants find XAI useful for debugging and communication, though they agree its usefulness is use-case-dependent and not always guaranteed.
+They expect TDA not to be an exception (I2, I5) but are optimistic about its potential, especially as data is seen as the most important variable in model development (I1, I2, I3, I5): “[What] drives your model is your data.
+[…]” (I5) (Theme: Data-focused debugging).
+In line with this theme, all participants also agreed that the model debugging process is largely data-focused, with training data being viewed as a lever for model performance: “It’s often the most, it’s often the biggest lever that we have in solving a problem.” (I1) (Theme: Data-focused debugging).
+Participants were curious whether TDA could be a potential tool for understanding bugs and enabling them to curate their datasets effectively.
+When comparing TDA to other XAI approaches like feature attribution (FA), participants point out that the TDA and FA target insights about model behaviour at different granularities (I2, I5, I6) and could be complimentary (Theme: Complimentary to FA).
+For instance, I1 mentioned that TDA could save time by identifying faulty training data while I2 pointed out that TDA and FA could be used together to identify the relationship between the faulty training data and learned features.
+We observed that participants have different notions of interesting training samples, with some focused on finding faulty data (I1) and others on identifying samples similar to a specific one (I5, I6) (Theme: Intuitions on data explanations).
+In conclusion, our interview study highlights that TDA has significant potential to address persistent data-centric challenges in model development, particularly in handling training data quality and data-focused debugging.
+However, participants revealed divergent mental models of what TDA entails and what they consider as a relevant sample to identify.
+This suggests a conceptual gap that may hinder effective adoption of TDA-based data-centric explanations.
+This motivated us to further investigate which specific forms of data-centric information practitioners actually need and value during model development.
+In the second stage of our needfinding study, we explore these information needs to better align TDA-driven tools with real-world development workflows.
 4.3. Defining a design space of TDA explanations
-| Theme A: Data quality issues | ||||
-| Theme B: Different notions of interesting training samples | ||||
-| ID | Theme | Code | Description | Participants |
-| #1 | A | Distribution shifts | If the difference in data distribution between training and testing data is too large, the model does not learn. This can also happen over time. | P3, P4 |
-| #2 | A | Data unavailability | Missing data entries or missing labels are common problems, particularly for tabular data. | P2, P3, P7, P8, P9 |
-| #3 | A & B | Wrong labels | Annotation is an expensive task subject to human error. mislabelled data is hence a possible reason for model error. | P2, P4 |
-| #4 | A | Data scaling | Wrong data scaling in tabular data (e.g. using meters instead of centimetres) can lead models to learn wrong correlations. | P3 |
-| #5 | A | Historical data | Outdated data that is not cleaned out of the training dataset can contaminate the training. | P9 |
-| #6 | B | Outlier data points | Data outliers, i.e. mislabelled data or “unusual data points” (P3) can aid for model debugging. | P3, P8 |
-| #7 | B | Decision boundary samples | Samples at decision boundary are interesting to understand the model behaviour. | P7 |
-| #8 | B | High-confidence samples | Samples that the model is sure about tend to represent what the model has learnt and help in understanding the model behaviour. | P7 |
-| #9 | B | Similar training samples | Training samples that are similar in features to the test sample and how the model behaves on these is helpful for understanding model behaviour. | (P6), P7, P8 |
-| #10 | B | Group of samples | Inspecting multiple training samples is generally more interesting to estimate model behaviour. | P8 |
-The figure depicts a three-dimensional coordinate system with categorical axes.
-The x-axis is named “Action defining relevance” and has the categories: remove, change label, reweight and recommend data collection.
-The y-axis is named “Metric” and has the categories: Loss, classification probability and ground-truth probability.
-The z-axis is named “Number of training samples” and has the categories: Explain f(x) with 1 training sample, and explain f(x) with N training samples.
-The interview analysis reveals model developers as a key audience for TDA explanations, with model debugging as a promising use case.
-Before addressing RQ2 – What do users need from TDA explanations?, we translate these insights into a design space of TDA explanations in model debugging.
-The space shall illustrate how explanatory information about training samples and model output could look like, forming the basis for the study addressing RQ2.
-Interview participants provided rich information about what they usually look out for in the training data when debugging a model.
-We coded this information into the following themes: Data quality issues, and different notions of interesting training samples.
-We present the descriptions of the corresponding codes in Table 2 and use them to identify different axes of information that could be represented by TDA explanations.
-Specifically, we find that different data artefacts are debugged using different actions, with users interested in various metrics and considering different numbers of training samples.
+Drawing from our interview findings, we explore the design space of TDA explanations in model debugging.
+Interview participants provided rich information about what they usually look for in the training data when debugging a model, and for which purpose.
+Even though the objective and desired approach are similar, i.e., improving the model performance given a particular bug and data-centred debugging, respectively, developers tend to seek different information first.
+Specifically, we found that different data artifacts are investigated using different actions, with users interested in various metrics and considering different numbers of training samples depending on the context.
 From these observations, we define a three-dimensional design space (see Figure 4).
-In the following, we detail the reasoning behind each axis and connect it to the codes in Table 2 via the ID column.
+In the following, we detail how each axis was driven from the themes and codes found in the interview study (Figure 3).
 4.3.1. Axis 1: Action defining relevance
-The first axis corresponds to the actions participants propose for different types of data artefacts: Removing data, changing the label, reweighting data and collecting more data (x-axis of Figure 3).
-Participants look at data to remove when it is clearly malicious data and likely hurts model learning, for instance mislabelled (#3) or outdated, historical data (#5).
-Additionally, we find that participants look for data that they wish to change, for example filling in or imputing missing data (#2), or correcting wrong data scales (#4) and labels (#3).
+The first axis corresponds to the actions participants propose for different types of data artefacts: Removing data, changing the label, reweighting data and collecting more data (Figure 4).
+This axis is inspired by the participants’ statements about the data-centric nature of development work (Theme: Data-focused debugging).
+Dataset curation and manipulation is a common activity in development work and has been described as the “[largest lever]” (I1) for model performance.
+Different actions were mentioned or are implied in the interview study, which are linked to different hypotheses about the cause:
+Participants look at data to remove when they believe that the model may have learnt the behaviour of interest from malicious data, for instance, mislabelled or outdated, historical data (Theme: Data-focused debugging Code: Artifacts are a common reason for errors, Theme: Data quality Codes: Wrong labels, Historical data).
+Additionally, we find that participants look for data that they wish to change, for example filling in or imputing missing data, or correcting wrong data scales and labels (Theme: Data quality Code: Data unavailability, Data scaling, Theme: Evaluation Code: Missing or unreliable data labels).
 Especially when data is hard to obtain (e.g.
 in the medical domain), removing data is not an option and correcting data is preferred (P6).
-Another action is the reweighting of training samples and collecting more data related to the error.
-These actions emphasise specific samples in the training process to encourage the model to learn them.
-This is particularly relevant to learning outliers (#6) and handling distribution shifts (#1).
-Removal, change and addition of data have different effects on a model and therefore represent different quantifications of attribution: For instance, removal as in traditional TDA (Equation 1 in § 3) or by addition of data.
-Essentially, these examples map to the questions: How would the model behaviour change if one were to (a) exclude a training sample?
-or (b) add more training samples like the test sample to the dataset?
-By understanding preferences in this axis, we aim to identify what is needed for data attribution to be actionable and reflect the debugging process.
+Another action is the reweighting of training samples and collecting more data related to the error which participants mentioned to be rather easy in some domains like autonomous driving (P1).
+These actions emphasise specific samples in the training process to encourage the model to learn them (Theme: Data-focused debugging Code: New training data is easy to acquire).
+This is particularly relevant to learning outliers and handling distribution shifts (Theme: Data quality Code: Distribution shift, Theme: Intuitions on data explanations Code: Outlier data points).
+Removal, change and addition of data have different effects on a model and therefore represent different quantifications of attribution that answer different questions.
+For instance, the removal action implies a question like “How would the model behaviour change if one were to exclude a training sample?” which corresponds to traditional TDA (Equation 1 in § 3).
+The addition of training data, however, implies the question of “How would the model behaviour change if one were to expand the dataset?” which is a largely different question.
+By studying preferences in this axis, we aim to identify what data attribution needs to represent to be actionable and reflect the debugging process.
 4.3.2. Axis 2: Metric
 The second axis corresponds to the sample-wise metrics participants used when referring to model behaviour: the loss, the probability of the true class label, and the probability of the predicted class label (y-axis of Figure 4).
-The interviews show not all participants think of model behaviour in terms of sample loss, but also in classification probability.
-While high loss can be an indicator for wrong or outlier samples (#3, #6), participants voiced their interest in samples where the model has high or low classification probability (#7, #8).
-Such samples provide context of the global model behaviour to the model developer, and can therefore be helpful for their work.
-The metric can be seen as a measure by which participants primarily perceive model behaviour in the debugging process.
-By understanding metric preferences, we aim to identify the metric that helps model developers most in debugging, testing whether the loss as in traditional TDA is the most actionable measure.
+This axis is inspired by the quantities participants talked about when explaining their intuitions about TDA (Theme: Intuitions on data explanations).
+The interviews show that not all participants think of model behaviour in terms of sample loss, but also in classification probability.
+While high loss can be an indicator for wrong or outlier samples (“[Samples with high training loss are] often […] mislabeled examples.” (I1), Theme: Intuitions on data explanations Code: Outlier data points), participants also talked about relevant samples as those where the model has high or low classification probability.
+Such samples provide context to the global model behaviour to the model developer, specifically indications about the decision boundary (“I would be interested in two kind of samples, the ones that [the model is] super confident with.
+But also the samples that are, I guess 50-50, it could be either one class [to see] why [the model] still chose class A.” - I4, Theme: Intuitions on data explanations Codes: Decision boundary samples, High-confidence samples) and can therefore be helpful for their work.
+The metric can be seen as a measure by which participants primarily perceive per-example model performance in the debugging process: The loss is indicative of the performance of the task, but also prediction probability interpreted as model confidence in a decision matters for understanding model behaviour.
+Both offer different types of information about the model.
+By understanding metric preferences, we aim to identify the metric that helps model developers most in debugging and testing whether the loss, as in traditional TDA, is the most actionable measure.
 4.3.3. Axis 3: Number of training samples
@@
 We identify two main options: A single sample and a group of samples (z-axis of Figure 4).
-We found that model developers usually do not inspect just one training data sample but think more globally as they usually refer to training data and samples in plural form (#10).
-This aspect is addressed in TDA research by works studying group attribution (Koh et al., 2019; Basu et al., 2020; Park et al., 2023).
-By understanding preferences in this axis, we aim to identify whether individual attribution like in traditional TDA or rather training data groups need to be in focus of TDA research.
-4.4. Survey study
-Building on the findings of the interview study, we aim to study the needs of model developers and understand what kind of information is useful, addressing RQ2 – What do users need from TDA explanations?
-We design an online survey that explores the design space of TDA explanations (see Figure 3) using scenario-based design (Carroll, 1995).
-We piloted the survey to ensure clear phrasing and compatibility across devices.
+This axis is inspired by the explicit answer of I5 (“I think [the most relevant data] would be like a collection of [samples].”) and the implicit indication of talking about multiple samples in the participants’ answers, recognisable through the plural form use of “data”, “samples” and “artefacts” (Theme: Intuitions on data explanations Code: Group of samples).
+We found that model developers usually do not inspect just one training data sample but think more globally in groups of relevant data.
+This aspect is addressed in TDA research by works studying group attribution (Koh et al., 2019; Basu et al., 2020; Park et al., 2023; Hu et al., 2024).
+By understanding preferences in this axis, we aim to identify whether individual attribution, like in traditional TDA, or rather training data groups, needs to be in focus of TDA research.
+4.4. Scenario-based Interactive User Study
+Building on the findings of the interview study, we designed an interactive online study that explores the design space of TDA explanations (see Figure 4) using scenario-based design (Carroll, 1995).
+We piloted the study interface to ensure clear phrasing and compatibility across devices.
 Data was collected from June – August 2024.
 4.4.1. Participants
-We recruit participants with prior experience in developing machine learning models to closely align with potential real users.
-To recruit expert participants, we used snowball sampling alongside social media ads on X and LinkedIn, reaching out to contacts and asking them to share the study link.
-We expect thematic saturation in the qualitative analysis and meaningful statistics in the quantitative analysis at around 30 replies.
+We recruited participants with prior experience in developing machine learning models to closely align with potential real users.
+Contrary to the interview study, we did not constrain our participants’ work to be in a high-risk application domain for improved reach.
+To recruit expert participants, we used snowball sampling alongside social media ads on X and LinkedIn, reaching out to personal contacts and asking them to share the study link.
+We aimed for around 30 replies to provide sufficient diversity in the answers for the qualitative analysis while also being indicative of measurable effects for a quantitative analysis.444We conducted a test of statistical significance.
+A power analysis at significance level , power and large effect size yields a sample size of 31.
 Of 34 responses, we excluded three for low-quality answers (short and non-sensical), leaving 31 participants (nine female).
@@
 While the data types used by participants are diverse too, the majority of participants work with image or tabular data, as used in our study (see Figure 5).
-The participants were compensated with 25€ via Tremendous111https://www.tremendous.com/.
-This figure shows five histograms about the answer distributions for the preliminary demographic questions of the survey study.
-From the top right corner to the bottom left corner, the histograms show the gender distribution (22 male vs.
-9 female), whether people work in industry or academia (13 academia, 17 industry, 1 in both), years of experience distribution (1 ¡1 year, 15 1-3 years, 12 3-5 years, 3 ¿5years), the distribution of data types that participants work with (19 Image, 16 tabular, 14 time series, 15 text, 8 video, 1 biological, 2 graphs, 1 audio and 1 robot data), the distribution of experiences with deep learning models (27 yes vs.
-4 no).
+The participants were compensated with 25€ (=$27.05) via Tremendous555https://www.tremendous.com/.
 4.4.2. Scenario-based design
 The main part of the study is constructed with scenario-based design (Carroll, 1995) to base the analysis on a tangible usage scenario.
-We create two scenarios of model debugging in an interactive mock-up of a model development suite (see Figure 5).
-The choice of data types (image and tabular data) is based on the dominant modalities of interview participants.
-The study puts participants in two imaginary scenarios where they are model developers in a company that builds (1) a bird classification app, and (2) a credit scoring app.
+We created two scenarios of model debugging in an interactive mock-up of a model development suite (see Figure 6).
+The choice of data types (image and tabular data) was based on the dominant modalities of interview participants.
+The study put participants in two imaginary scenarios where they are model developers in a company that builds (1) a bird classification app, and (2) a credit scoring app.
 The company recently acquired a data-centric tool to help model developers debug their models by understanding errors and identifying training data relevant to those errors.
 The tool is customisable to adapt to the developer’s preferences.
-The customisation choices, based on the TDA design space (see Figure 3) are:
+The customisation choices, based on the TDA design space (see Figure 4) were:
 Action defining relevance: (a) Removal, (b) Label change (to the next most likely class), (c) Upweighting (implemented with a factor of 10).
@@
 We limit the number to 10 to explore preferences for more training samples while keeping computation efforts low.
-In both scenarios, participants are presented with a specific misclassification as the model error (left side of Figure 5).
-This error is randomised across participants to prevent bias from a specific sample.
-Participants are asked to choose which type of training data-related information (right side of Figure 5) is most useful to understand the reasons for the error.
-This way, we aim to elicit participants’ intuitions about the most actionable information for debugging the model error.
-Participants are encouraged to explore different settings and submit their desired customisation to complete the task.
-The data for these scenarios was generated using real machine learning models and openly available datasets.
-We detail the data generation process in Appendix C.
-The figure shows a screenshot of the bird classification scenario interface of the survey study.
-The page is divided into two parts: On the left is the interface with the header “Customize your own development interface”, and on the right are the customization controls as an expanded menu.
-The controls have the instructions: “Choose a setting below and click Apply to see the effects.” Below are three drop down lists, one for Action defining relevance, one for Metric of interest and the last one for the number of training samples.
-The settings are set to “Removal”, “Loss” and “1” respectively.
-Below the dropdown lists is a grey button with the word Apply on it.
-The interface itself is divided into two panels as well: on the left, indicated with a “1” is a panel with the title “Test error analysis”.
-Below, the test accuracy is given as well as a description of the Error (“Wren is misclassified as Sparrow.”).
-Underneath the error description is the image of the wren that is misclassified with information about the prediction, loss, predictive probability and ground truth probability.
-The second panel, indicated with a “2” is titled “Training data analysis”.
-It shows both the individual effect (how does the model behaviour on the erroneous test sample change?) of the setting chosen in the customisation menu as well as the global effect (how does the overall model behaviour change?).
-Below the effect is a table with the most relevant training samples according to the chosen setting, with their ground truth and predicted labels.
-The dynamic information is highlighted by an orange box with arrows pointing from the customisation menu to the dynamic panel.
-4.4.3. Survey structure
-The survey structure is shown in Figure 6.
-Participants are first briefed on the study’s objective and data processing.
-After providing consent and confirming they meet inclusion criteria, they are asked preliminary demographic questions for context to the later analysis.
-In the remaining study, we aim to extract the participant’s preferences through a usage scenario (§ 4.4.2).
-The participants are presented with the scenario description and prompted with “Imagine you are a model developer in a company […]” to immerse them in the use case.
-Then, we ask the participants follow-up questions to understand the reasoning behind their choices through free-text questions.
-These free-text questions are about understanding the participants’ intuitions of why the model may have made an error, and how they would fix it.
-After completing one use case, the participants are presented with the other use case, where the order is randomised across users to prevent an order bias.
-The figure shows a flowchart of the survey.
-The survey starts with a question about prior model developing experience to assess whether inclusion criteria are met.
-Then, the next state of preliminary questions follow to record context to the data.
-This includes demographic questions and questions about prior model development experiences.
-The third state is the tutorial for the interactive application, after which the quantitative data collection is started with either the bird classification (image) use case or the credit scoring (tabular) use case.
-In this stage, participants can explore the interface and are asked to submit their final choices.
-The next stage is the qualitative data collection, where participants are asked to explain their choices and proposed actions in free text.
-Then, an arrow indicates that the other use case will be presented before the end of the survey.
+In both scenarios, participants were presented with a specific misclassification as the model error (left side of Figure 6).
+This error was randomised across participants to prevent bias from a specific sample.
+Participants were asked to choose which type of training data-related information (right side of Figure 6) is most useful to understand the reasons for the error.
+This way, we aimed to elicit participants’ intuitions about the most actionable information for debugging the model error.
+Participants were encouraged to explore different settings and submit their desired customisation to complete the task.
+The data for these scenarios was generated using real machine learning models and openly available datasets (Wah et al., 2011; Hofmann, 1994).
+We detail the data generation process in Appendix B.
+4.4.3. Study structure
+The structure of the scenario-based interactive study is shown in Figure 7.
+Participants were first briefed on the study’s objective and data processing.
+After providing consent and confirming they meet inclusion criteria, they were asked preliminary demographic questions for context to the later analysis.
+In the remaining study, we aimed to extract the participant’s preferences through a usage scenario (§ 4.4.2).
+The participants were presented with the scenario description that details their objective of understanding possible reasons for a presented error.
+They were prompted with “Imagine you are a model developer in a company […]” to immerse them in the use case.
+Then, we asked the participants follow-up questions to understand the reasoning behind their choices through free-text questions (e.g.“Why did you choose to look at the {chosen_metric} metric?”).
+These free-text questions were about understanding the participants’ intuitions of why the model may have made an error (“In your opinion, why did the model fail at the above classification?”), and how they would fix it (“What would you do to fix this error?”).
+After completing one use case, the participants were presented with the other use case, where the order is randomised across users to prevent an order bias.
 4.4.4. Data analysis
-We analyse the survey study both quantitatively and qualitatively.
-For the quantitative analysis, we aim to understand whether a specific notion of TDA is preferred across participants because previous TDA studies have been centred around a single notion of TDA (§ 3).
-Since the traditional TDA setting is already a part of the design space, studying preferences gives an answer to whether participants really preferred this particular scenario.
-We further assess if participants prefer a specific scenario —- if not, it is a strong signal that the research should be diversified into multiple possible scenarios serving different user needs.
-To this end, we compute the distribution of the participants’ choices (action, metric and number of samples).
-Furthermore, we test whether the collected data indicates a statistically significant preference for a specific setting.
-We perform a test as the data is categorical (Pearson, 1900).
-For the number of training samples axis, we perform Fisher’s exact test as the choice frequencies are low (Edwards, 2005).
+We analyse the scenario-based interactive study both quantitatively and qualitatively.
+The quantitative analysis aims to examine the diversity of participant preferences in TDA explanations.
+Specifically, we investigated whether the predominant form of explanation in TDA, i.e.
+the counterfactual effect of removing a sample (§ 3), aligns with the explanatory needs of model developers.
+If there is alignment, TDA research focuses on the right problem, and methods produced in this field are valuable to model developers.
+However, if this is not the case, it is a strong signal that the research should be diversified into multiple possible scenarios serving different user needs.
+To this end, we computed the distribution of the participants’ choices (action, metric and number of samples).
+We tested whether the collected data indicates a statistically significant preference for a specific setting.
+We performed a test per dimension of the design space (action defining relevance, metric, number of training samples) as the data is categorical (Pearson, 1900).
+We considered two categories for the number of training samples, the first one being a single sample and the second one being groups larger than one as defined in the design space (Figure 4).
 Specifically, our hypotheses are:
@@
 The qualitative analysis aims to understand the reasons behind participant choices and get deeper insight into user preferences.
-The free-text responses in the follow-up questions are analysed using inductive thematic coding (Braun and Clarke, 2012) by two coders.
-Per question, we analyse the common themes among participants to extract the reasons behind user preferences and needs.
-The analysis consists of several coding workshops and is depicted in Figure 7.
-We measure interrater agreement by the percentage of overlapping themes.
-The final agreement is 97.1%.
-The figure shows the progression of the thematic analysis of the survey data in a timeline-like schematic consisting of 8 parts: The initial coding workshop, where the data of five participants was analysed, resulted in 21 themes.
-The next stage is independent coding of eight more participants, followed by the first intermediate coding workshop that resulted in 29 themes with an IRR of 75.8%.
-This is repeated and the next intermediate coding workshop resulted in 31 themes with an IRR of 84.3%.
-Then, four more participants are independently analysed before the final coding workshop that resulted in 32 themes and and IRR of 97.1%.
-Lastly, the stage of case-by-case discussions of themes indicates that single participants who answered the survey after the main analysis are discussed on a case-by-case basis.
-4.5. Survey findings
-We conduct the survey study to answer RQ2 – What do users need from TDA explanations?
-The survey data analysis indicates a clear connection between users’ intuition of what causes a model error and their TDA preferences.
+The free-text responses in the follow-up questions were analysed using inductive thematic coding (Braun and Clarke, 2012) by two coders.
+Per question, we analysed the common themes among participants to extract the reasons behind user preferences and needs.
+The analysis consisted of several coding workshops and is depicted in Figure 8.
+We measure interrater agreement in Krippendorff’s (Krippendorff, 2011).
+The final agreement is .
+4.5. Study findings
+We conducted the interactive study to study model developers’ TDA needs through a scenario.
+The data analysis indicates a clear connection between developers’ intuition of what causes a model error and their TDA needs.
 We elaborate further in the following.
@@
 Table 2 shows the frequency of selections and the associated p-values.
-We test if there exists a setting of action, metric and number of training samples which is consistently chosen and particularly useful to model developers.
-We are especially interested in whether there is a clear preference for removal, loss and one training sample as these correspond to the traditional notion of TDA introduced in § 3.
-For the action axis, there is no statistically significant preference for any choice and the distribution is rather uniform across choices.
+We test whether a specific action, metric, and number of training samples are consistently chosen and particularly useful to model developers.
+We are especially interested in a clear preference for removal, loss, and one training sample as these correspond to the predominant notion of TDA introduced in § 3.
+The analysis yields the following results: For the action axis, there is no statistically significant preference for any choice, and the distribution is rather uniform across choices.
 While the results may show a higher frequency of participants choosing loss as a metric, we observe no statistically significant preference.
 Hence, we fail to reject for the action and metric axes and conclude that there is no dominant preference for a particular TDA setting.
-For the number of samples axis, however, we find a statistically significant preference for a group of 10 samples.
-The entropy analysis supports this conclusion (see Table 4).
+Instead, developers have individual preferences.
+For the number of samples axis, however, we observe that all of our participants chose ¿1 samples across both scenarios.
+More specifically, we find a statistically significant preference for a group of 10 samples, indicating a clear preference for training data groups instead of individual samples.
+The entropy analysis further supports this conclusion (see Table 3).
 As a measure of diversity, normalised entropy scores range from 0.588 to 0.997, showing a highly diverse (¿0.5) distribution of choices.
-This is especially true for the Action defining relevance and Metric axes with high scores (¿0.8).
+This is especially true for the action and metric axes with particularly high scores (¿0.8).
 There is less diversity in the number of training samples but a clear preference for choosing groups of samples, confirming Ilyas et al.’s (Ilyas et al., 2022) intuitions for studying group attribution.
@@
 | Action | ||||||
-| Removal | 33.33% | 1 | 41.97% | 1 | 37.70% | 1 |
-| Label Change | 30.00% | 1 | 16.13% | 1 | 22.95% | 1 |
-| Upweighting | 36.67% | 1 | 41.94% | 1 | 39.34% | 1 |
+| Removal | 33.3% | 1 | 42.0% | 1 | 37.7% | 1 |
+| Label Change | 30.0% | 1 | 16.1% | 1 | 23.0% | 1 |
+| Upweighting | 36.7% | 1 | 41.9% | 1 | 39.3% | 1 |
 | Metric | ||||||
-| Loss | 60.00% | 1 | 35.48% | 1 | 47.54% | 1 |
-| Pred. cls. prob. | 10.00% | 1 | 41.94% | 1 | 26.23% | 1 |
-| GT class prob. | 30.00% | 1 | 22.58% | 1 | 26.23% | 1 |
+| Loss | 60.0% | 1 | 35.5% | 1 | 47.5% | 1 |
+| Pred. cls. prob. | 10.0% | 1 | 41.9% | 1 | 26.2% | 1 |
+| GT class prob. | 30.0% | 1 | 22.6% | 1 | 26.2% | 1 |
 | Image | Tabular | Both | ||||
@@
 | Num. of samples | ||||||
-| 1 | 0.00% | 1 | 0.00% | 1 | 0.00% | 1 |
+| 1 | 0.0% | 1 | 0.0% | 1 | 0.0% | 1 |
+| ¿1 | 100.0% | 0 | 100.0% | 0 | 100.0% | 0 |
 | 2 | 0.00% | 1 | 0.00% | 1 | 0.00% | 1 |
@@
 4.5.2. Qualitative results
-The figure shows the themes found in the coding analysis of the survey study, organised into four theme areas.
-The first theme area is called “Action depends on user’s hypotheses” and includes the themes on different hypotheses that participants have about the root cause of the model error (mislabelled data, incomplete training, spurious correlations, similar class features, low dataset diversity, class overlap, dataset imbalance and small dataset size) as well as themes related to the proposed bugfixing actions users proposed (Remove noisy data, changing label to correct, fix labelling pipeline for future data, training process engineering, model architecture engineering, feature engineering, upweight data to focus on specific features, feature attribution inspection, manual inspection of datapoints, collect specific new data, synthesize/augment data, resample/reweigh data and collect any new data).
-The second theme area is called “Individual metric preferences” and includes themes on preferred metrics (probabilities are intuitive and show the decision boundary, ground truth probability is the target, loss is most informative in the multi-class case) and one theme related to preferred actions (Label change together with a probability metric show the decision boundary).
-The third theme area is titled “Group attribution preferences” with themes Group for overview of the error, Balance betw.
-variety and conciseness.
-The last theme area is titled “Reliability of explanations” and includes themes related to the preferred number of training samples (group size enough for reliable attribution, small group to prevent reasoning change), preferred metric (Ground truth probability is the target) and preferred action (Removal shows causal effect).
-The inductive thematic coding analysis of the free-text answers resulted in four thematic areas and theme types covering 30 themes (see Figure 8): (1) Action depends on the user’s hypothesis, (2) Individual metric preferences, (3) Group attribution preferences, and (4) Reliability of explanations.
-For the theme types, we find themes that correspond to the user’s hypothesis about the error and observe that these are often coded together with a corresponding proposed action.
-However, preferred metrics and preferred number of training samples do not correspond to the user’s hypothesis.
-Additionally, we identify four themes that do not correspond to any category but give interesting insights and context to the analysis.
-The full description of each theme is in Appendix D.
-Action depends on user’s hypotheses.
-The thematic analysis complements our quantitative findings: User preferences in model debugging are highly individual.
-The analysis provides a reason: Individual preferences are often linked to the user’s understanding and intuitions, in other words, their hypothesis about an error.
-The user’s hypothesis determines their course of action.
-Hypotheses are highly individual because users formulate them based on their expertise and past experiences.
-For example, we identified the theme of low dataset diversity as a hypothesis participants have about a model error: “[Images] […] in the training set are not diverse enough”, “[One class] was not representative enough”.
-Corresponding to this hypothesis, participants often suggested adding new data or applying data augmentations to diversify the training set, which we identified as collect specific new data and synthesise/augment data themes.
-Individual metric preferences.
-For individual metric preferences, we identify a link with the task at hand.
-In the case of binary classification, participants showed a preference for probability metrics, as it is easier to understand than the loss and allows an inspection of the decision boundary (“The loss is not very intuitive”).
-In multi-class classification, more participants preferred the loss as it considers all possible classes (“[The loss] contains information on all classes”).
+The inductive thematic coding analysis of the free-text answers resulted in eight themes that address the reasoning behind participants’ choices for action, metric, and number of samples preferences (see Figure 9).
+We elaborate on the themes below.
+Theme: Error hypotheses informs action
+The thematic analysis complements our quantitative findings: Developer preferences in model debugging are highly individual.
+The analysis provides a reason in line with the sensemaking framework introduced in Cabrera et al.
+(Cabrera et al., 2023): Individual preferences for proposed actions are often linked to the developer’s understanding and intuitions.
+Needs are highly individual because developers may have different mental models based on their expertise and past experiences.
+In other words, their mental models of the model’s learned behaviour affect their hypothesis about an error and the chosen action.
+The developer’s hypothesis determines their explanatory needs: An explanation needs to enable the developer to reject or confirm their hypothesis of why a model error occurs so that they can make informed decisions and actions to resolve the error.
+For example, participants who hypothesise the cause of an error to be data-related also propose data-related actions: “[Images] […] in the training set are not diverse enough”, “[One class] was not representative enough”.
+Corresponding to this hypothesis, participants often suggested adding new data or applying data augmentations to diversify the training set.
+If the hypothesis is linked to the model setup, the proposed action corresponds to the model.
+Themes about type of actions
+In terms of types of actions, we identified three main themes: Data-based actions, model-based actions, and inspection actions.
+In particular, the data-based actions are relevant to studying TDA needs (Theme: Data-based actions).
+In line with the interview study findings, we found that there are various data-based actions that participants propose: From removing malicious data to balancing the dataset through reweighting or adding new data.
+Additionally, we found that participants propose synthetically creating new data to add diverse data or fill up undersampled parts of the dataset (“address class imbalance by synthesizing data”).
+This shows that the actions participants propose are various and our collection is not likely to be exhaustive.
+Hence, useful and actionable TDA needs to facilitate a variety of actions that help developers verify their hypotheses about the model error.
+Model-based actions address changes in model architecture or hyperparameter tuning and are connected to participant hypotheses about incomplete training (“Apply regularization methods like dropout or batch normalization to improve generalization and reduce overfitting.”) (Theme: Model-based actions).
+Inspection actions are linked to first understanding a suspected bias of the model to get a clearer hypothesis about the error (Theme: Inspection actions).
+We found feature attribution explanation approaches (“Check which features the model considers important using techniques like SHAP […].”) and manual inspection-based approaches (“Typically, I would focus on the specific test sample and try to understand why the model is making a mistake by analyzing the feature values and their relationships and checking for outliers or anomalies.”).
+TDA explanations were not mentioned, but could ideally support these inspection activities.
+Themes about metrics
+We identified a link between the preferred metric and the task at hand.
+In the case of binary classification, participants showed a preference for probability metrics, as it is easier to understand than the loss and allows an inspection of the decision boundary (“The loss is not very intuitive”, Theme: Metric informative of decision boundary).
+In multi-class classification, more participants preferred the loss as it considers all possible classes (“[The loss] contains information on all classes”, Theme: Metric informative of training objective).
 However, as the quantitative analysis shows, these differences in preference are not statistically significant.
-Therefore, the metric preferences are likely individual to the user and their habits.
-Group attribution preferences.
-The quantitative study shows a clear preference for attribution to groups of training samples as opposed to individual training samples.
-The qualitative analysis finds that this is due to the user’s need for having a good overview of the model error.
-Inspecting a group of training samples provides a better overview than a single sample, and allows the user to understand whether their hypothesis about the error is valid or not (Group for overview of the error).
+Therefore, the metric preferences and actionability are likely individual to the user and their habits.
+Themes about number of training samples
+The quantitative study showed a clear preference for attribution to groups of training samples as opposed to individual training samples.
+The qualitative analysis found that this is due to the developer’s need for having a good overview of the model error.
+Inspecting a group of training samples provides a better overview than a single sample, and allows the user to understand whether their hypothesis about the error is valid or not (Theme: Group for global effect).
 A second theme highlights another angle: While variety in the group is desired for a good overview of the error (e.g.
-group consisting of samples with diverse features), the group should be concise enough to “keep the analysis manageable” (Balance betw.
-variety and conciseness).
-Reliability of explanations.
-Our analysis finds that some themes are related to the overall concept of explanation reliability, where particularly the choice of how many training samples to inspect is associated with reliability.
-Participants stated that group sizes should be large enough for the attribution scores to be reliable.
-At the same time, he group size should be small enough to ensure reliable attribution scores, as changing the dataset could change the model’s reasoning (Small group to prevent reasoning change).
-We interpret this need for reliability to reflect in the motivation to uncover causal effects (“By removing […] I can determine if [the training samples] were causing confusion within the model”).
-Unintended themes.
-We identified four unintended themes outside the study design that provide additional context.
-We noticed that participants focused either on the model error (Focus on model error) or the final accuracy (Focus on final accuracy).
-Therefore, some of the participants explored all choice options in the survey and submitted the setting with the largest performance gain.
-This indicates that the end goal of performance improvement in the task is highly valued.
-Some participants mentioned that the study interface is unfamiliar (Unfamiliar study interface), which can be confusing.
-We considered this point in the analysis through discussions at the coding workshops.
-However, it also indicates that participants were unfamiliar with TDA explanations overall as they are not used in practice yet.
-The last theme highlights the importance of grouping erroneous test samples.
-Participants noted that a single erroneous test sample may represent a broader issue, and understanding the link between training data and groups of test samples with similar errors could enhance debugging.
+group consisting of samples with diverse features), the group should be concise enough to “keep the analysis manageable” and group size should be small enough to ensure reliable attribution scores, as changing the dataset could change the model’s reasoning (Theme: Group precision).
+In conclusion, the analysis of the scenario-based interactive user study deepens our understanding of model developers’ needs for data-related explanations.
+We identify why there are diverse information needs across different developers: Since each developer may have a different mental model of their ML system, the information that is actionable to them is diverse as well and results in highly individual needs.
 5. Discussion
-This work presents a needfinding study with AI practitioners for training data attribution (TDA) explanations to inform user-focused TDA research in a top-down approach.
-We summarise our findings to provide answers to the initial research questions (§ 5.1), identify research gaps that need to be addressed for more user-focused TDA (§ 5.2) and discuss limitations (§ 5.3).
+This work presents a needfinding study for training data attribution (TDA) explanations in model development to inform user-focused TDA research in a top-down approach.
+We summarise our findings to provide answers to the research questions (§ 5.1), identify research gaps that need to be addressed for more user-focused TDA (§ 5.2).
+We discuss the study framework in a general sense as an approach to bridge method and needfinding research especially in early development phases of a technology (§ 5.3), followed by limitations (§ 5.4).
 5.1. Answers to the research questions
-5.1.1. RQ1: What do users need TDA explanations for?
-To answer this question, we conducted the interview study (§ 4.1).
-Our findings suggest that TDA as an explanation method has the potential to serve similar high-level purposes as other explanation techniques, namely enabling users to understand model behaviour, especially unexpected model behaviour.
-By offering insights into model behaviour, TDA can assist end-users in calibrating trust by providing context for model decisions and support model developers in debugging errors by identifying root causes in the training data.
-This data-centric approach is particularly valuable in model development, leading us to identify model developers as the primary audience for TDA explanations and model debugging as the key use case.
-5.1.2. RQ2: What do users need from TDA explanations?
-To answer this question, we designed a survey study that extracts user preferences for the information represented by TDA (§ 4.4).
-Our study reveals that users need different types of information to understand model behaviour, depending on their hypotheses of why the model errs during development.
+This study sets out to understand user needs for TDA explanations in model development.
+We pose three questions to understand use cases, explicit and latent needs, and understudied areas of research in user-focused TDA.
+We provide our answers based on our study below.
+5.1.1. RQ1: What data-related explanation needs, both explicit and latent, do model developers have?
+To address this question, we conducted a two-part user study comprising an interview study (§ 4.1) and a scenario-based interactive study (§ 4.4).
+We aimed to explore how data-related explanations are used in model development and to gain a deeper understanding of both explicit and latent needs.
+Explanations are needed when the model behaves unexpectedly.
+Since ML development is largely data-driven, developers often inspect unexpected model behaviour from a data-centric perspective (Interview theme: Data-focused debugging).
+We observed a strong connection between a developer’s mental model of the ML system and their corresponding explanation needs.
+In practice, developers tend to form hypotheses, often data-related, to make sense of unexpected model behaviour (e.g., a developer may hypothesise that a model behaves unexpectedly due to data artefacts).
+Explanations serve as a means to confirm or reject these hypotheses and to refine or generate new ones, and inform subsequent actions taken by the developer to improve the system (Interactive study theme: Error hypotheses inform action).
+Since each developer’s mental model is shaped by their experience and expectations, their explanation needs are also diverse.
+The needs TDA explanations must meet reflect developers’ data-centric information needs during model building.
+For example, developers seek insights to identify low-quality data when they suspect wrong labels to be the cause for a model error, or detect broader issues like distribution shifts when they believe this could have occured (Interview theme: Data quality), or recognise gaps where additional data is needed when they suspect their dataset to be insufficient for the model to learn the task (Interview theme: Data is most important).
+Hence, these findings reveal that model developers have explicit needs for explanations that help identify possible causes of model error, whether by examining the impact of removing certain training samples or by highlighting the potential value of adding new data.
+At a deeper level, developers also express latent needs: explanations must be actionable, reliable, and aligned with their mental models.
+Explanations are not only used to understand the model’s behaviour, but also to guide meaningful decisions during development.
+5.1.2. RQ2: To what extent do existing TDA approaches align with model developers’ needs?
+To answer this question, we reflect on existing TDA approaches in light of the identified explicit and latent needs: TDA explanations need to be actionable, reliable, and aligned with the developer’s mental models.
+From these needs, we identify different TDA tasks that developers need TDA explanations to fulfil, next to the overarching need for reliability (see Table 4).
+| Attribution type | Action | Metric | Times chosen | ||
+| Influence attribution (e.g., (Koh and Liang, 2017; Pruthi et al., 2020; Guo et al., 2021; Schioppa et al., 2022; Grosse et al., 2023; Park et al., 2023; Bae et al., 2024)) | Removal | Sample loss | 1 | 1 | 0 |
+| Group influence attribution (e.g., (Park et al., 2023; Basu et al., 2020; Koh et al., 2019; Lin et al., 2024)) | Removal | Sample loss | ¿1 | 1 | 11 |
+| Attributing errors | Any | Any | ¿1 | ¿1 | 4 |
+| Lacking data attribution | Addition | Any | ¿1 | ¿1 | 26 |
+| Decision boundary attribution | Label change | Any | ¿1 | ¿1 | 14 |
+Existing TDA approaches partly address explicit needs.
+The formal definition of TDA (§ 3) characterises attribution as the change in model output resulting from training with versus without a specific training sample (Hampel, 1974).
+(Koh and Liang, 2017; Grosse et al., 2023; Park et al., 2023; Bae et al., 2024; Choe et al., 2024; Kwon et al., 2023; Pruthi et al., 2020) focus on efficiently and faithfully estimating this quantity (influence attribution in Table 4) or its extension to groups of training samples (Basu et al., 2020; Park et al., 2023; Koh et al., 2019) (group influence attribution in Table 4).
+TDA has been applied to a variety of tasks, including the identification of mislabeled data, analysis of adversarial vulnerabilities, and detection of domain mismatches (Koh and Liang, 2017).
+Mislabel detection, in particular, is a widely used evaluation across several studies (Pruthi et al., 2020; Park et al., 2023; Schioppa et al., 2022), and TDA methods have demonstrated the ability to identify outliers and mislabeled samples in training data (Koh and Liang, 2017; Pruthi et al., 2020).
+These capabilities align with certain explicit developer needs, particularly those focused on identifying low-quality data to remove.
+However, since TDA quantifies a sample’s relevance through the lens of its removal, it is especially well-suited for scenarios involving data elimination, such as detecting harmful or noisy samples.
+In contrast, broader data-centric explanation needs, such as identifying distribution shifts or detecting underrepresented regions in the data, are generally not addressed by current TDA methods, as they are not typically designed with these needs in mind.
+Existing TDA work does not consider developers’ latent need for actionable information.
+Our study shows that developers have the latent need for data-based explanations to enable them to confirm or revise their mental models of model behaviour, which are diverse.
+Fulfilling this latent need is tied to the developer’s mental model and makes an explanation actionable.
+This human-centred aspect of actionability is rarely considered in existing TDA work.
+While existing TDA methods are useful for measuring the impact of removing and retraining on individual samples, because they are studied to quantify this objective, developers may require different forms of attribution to address a wider range of questions during model development (e.g., lacking data attribution in Table 4).
+Existing TDA work partly implicitly addresses developers’ need for reliability.
+Regarding the need for reliable information, existing TDA methods have been criticised for their low reliability in deep learning (Nguyen et al., 2023; Basu et al., 2021; Bae et al., 2022).
+In addition, our study introduces a different aspect of reliability that only becomes apparent when considering the full context of the use case (as we assessed participants within the context of a specific task).
+When a dataset is modified, the overall model behaviour can be affected and must be considered.
+For TDA, this means that when modifying the dataset to address a particular error, developers require explanations that not only suggest an effective modification but also ensure the model’s performance remains rather consistent.
+Our analysis underscores the importance of incorporating user perspectives to identify needs critical to the real-world application of this technology.
+Traditional TDA explanations only partially meet these needs, and further interdisciplinary efforts beyond method research are required to develop methods with improved actionability.
+5.1.3. RQ3: What kind of information produced by TDA will be useful for model developers?
+To answer this question, we examine the gap between the needs we found and the needs that existing TDA work meets (see Table 4).
+Our study reveals that developers need different types of information to understand model behaviour, depending on their mental models and hypotheses of why the model errs during development, in line with (Cabrera et al., 2023).
 This necessitates TDA explanations that are flexible and adaptable to individual needs.
-These needs vary based on the application domain and the user’s level of expertise.
-Experienced users, may form hypotheses about model errors based on their past experiences.
-In contrast, less experienced users may form hypotheses based on what they learnt previously about machine learning models.
-Users have a preference for group attribution, as groups provide more information than individual samples.
-Additionally, users need TDA explanations to be reliable considering how data-centric actions might affect the model as a whole.
+These needs vary based on the application domain and the developer’s mental models that are shaped by the developer’s experiences and expectations.
+Experienced developers may form hypotheses about model errors based on their past experiences.
+In contrast, less experienced developers may form hypotheses based on what they learnt previously about machine learning models.
+TDA that aligns with the developer’s mental models will be useful.
+Developers prefer group attribution, as groups provide more information than individual samples and are expected to paint a more conclusive picture of the model behaviour.
+Hence, information related to group attribution will be useful for model developers.
+Additionally, developers need TDA explanations to be reliable, considering how data-centric actions might affect the model as a whole.
 Reliability entails two main aspects: First, the information needs to be reliable for it to be actionable and effective for the task (i.e., model debugging).
 Second, TDA explanations should also account for any broader changes in the model’s behaviour that may affect the attribution to training samples.
-5.1.3. RQ3: To what extent do existing TDA approaches align with user needs?
-To answer this question, we reflect on existing TDA approaches in light of the identified needs: TDA explanations need to be adaptable to the user and provide reliable and actionable insight while maintaining consistency with the overall model behaviour.
-From these needs, we identify different TDA tasks that users need, next to the overarching need for reliability (see Table 5).
-| Attribution type | Action | Metric | Times chosen | ||
-| Influence attribution (e.g., (Koh and Liang, 2017; Pruthi et al., 2020; Guo et al., 2021; Schioppa et al., 2022; Grosse et al., 2023; Park et al., 2023)) | Removal | Sample loss | 1 | 1 | 0 |
-| Group influence attribution (e.g., (Park et al., 2023; Basu et al., 2020; Koh et al., 2019; Lin et al., 2024)) | Removal | Sample loss | ¿1 | 1 | 11 |
-| Attributing errors (e.g. (Jain et al., 2024)) | Any | Any | ¿1 | ¿1 | 4 |
-| Lacking data attribution | Addition | Any | ¿1 | ¿1 | 26 |
-| Decision boundary attribution | Label change | Any | ¿1 | ¿1 | 14 |
-Traditional TDA measures the influence of a training sample on model predictions by analysing the impact of its removal through leave-one-out retraining (Hampel, 1974).
-This method essentially evaluates model predictions in the context of a counterfactual change in the training data.
-It has been studied across different tasks, such as identifying mislabelled data, understanding adversarial vulnerabilities, and detecting domain mismatches (Koh and Liang, 2017).
-Mislabel identification, in particular, is a common evaluation strategy employed in several studies (Pruthi et al., 2020; Park et al., 2023; Schioppa et al., 2022).
-Consequently, traditional TDA partially addresses the need for adaptability to different use cases.
-However, since TDA quantifies the relevance of a training sample by its removal, it is particularly suited for scenarios where data elimination is necessary, such as malicious samples.
-It remains an open question whether alternative actions, such as changing labels or upsampling a data region, might provide improved explanations in terms of actionability.
-Our study indicates that users may have diverse ideas regarding the root causes of errors, and traditional TDA only addresses one of these perspectives.
-Regarding the need for reliable, actionable, and effective information, existing TDA methods have been criticised for their low reliability in deep learning (Nguyen et al., 2023; Basu et al., 2021; Bae et al., 2022).
-In addition, our study introduces a different aspect of reliability that only becomes apparent when considering the full context of the use case (as we assessed users within the context of a specific task).
-When a dataset is modified, the overall model behaviour can be affected and must be considered.
-For TDA, this means that when modifying the dataset to addressing a particular error, users require explanations that not only suggest an effective modification but also ensure the model’s performance remains rather consistent.
-Our analysis underscores the importance of incorporating user perspectives to identify needs critical to the real-world application of this technology.
-Traditional TDA explanations only partially meet these needs, and further interdisciplinary efforts are required to develop methods with improved actionability.
 5.2. Overlooked research topics for user-focused training data attribution
-This work adds to a collection of human-centred explainable AI (HCXAI) works calling to centre XAI research around the user (e.g., (Brennen, 2020; Kaur et al., 2020; Rutjes et al., 2019; Ehsan et al., 2024; Mei et al., 2023; Lin et al., 2024)).
-By focusing on TDA explanations specifically, we identify several research areas that are necessary to study to advance TDA research towards a focus on users.
-Mental models and TDA: Our study supports previous work (Rutjes et al., 2019; Kaur et al., 2020) that the most useful explanations depend on the user’s mental model of why the model errs: We find that what information is actionable to users depends in their hypothesis for reasons of model behaviour (Theme area: Action depends on user’s hypothesis).
-Therefore, future research should explore how TDA influences and is influenced by mental models, as well as how TDA can be adapted to reflect mental models (e.g.
-by defining relevance through the action of upweighting as opposed to removal), potentially leading to more aligned and actionable explanations that are well-defined.
-User-centred group attribution: The survey findings reveal a clear need for attributing model behaviour to groups of training samples, as it provides more informative insights (Theme area: Group attribution preference).
+This work adds to a collection of human-centred explainable AI (HCXAI) works calling to centre XAI research around users (e.g., (Brennen, 2020; Kaur et al., 2020; Rutjes et al., 2019; Ehsan et al., 2024; Mei et al., 2023; Lin et al., 2024)).
+We focus particularly on model developer users, as they have agency over the dataset.
+By focusing on TDA explanations specifically for model developers, we identify several research areas that are necessary to study to advance TDA research towards a focus on developers.
+Mental models and TDA: Our study supports previous work (Rutjes et al., 2019; Kaur et al., 2020) that the most useful explanations depend on the user’s mental model of why the model errs: We find that what information is actionable to users depends in their hypothesis for reasons of model behaviour (Theme: Error hypothesis informs action).
+Therefore, future research should explore how TDA influences and is influenced by mental models, as well as how TDA can be adapted to reflect mental models (e.g., by defining relevance through the action of upweighting as opposed to removal), potentially leading to more aligned and actionable explanations that are well-defined.
+User-centred group attribution: The survey findings reveal a clear need for attributing model behaviour to groups of training samples, as it provides more informative insights (Theme: Group for global effect).
 While existing approaches (Basu et al., 2020; Park et al., 2023; Koh et al., 2019) address this, our findings highlight the need for creating groupings that make sense to users (e.g., based on ground truth or predicted class), an aspect not yet widely explored in TDA research.
-Holistic understanding of model errors: Our thematic analysis of survey responses highlights the importance of understanding the model error itself before attempting to fix it (Theme: Grouping erroneous test samples).
+Holistic understanding of model errors: Our thematic analysis of survey responses highlights the importance of understanding the model error itself before attempting to fix it (Theme: Inspection actions, Group precision).
 If the root cause is a spurious correlation, the error likely extends beyond the current test sample and may not fully represent the issue.
-We recommend future research focus on understanding error types rather than isolated instances (e.g., slice discovery (Wang et al., 2024a)) and on attributing model behaviour on groups of test samples to training data (e.g., (Jain et al., 2024)).
+We recommend future research focus on understanding error types rather than isolated instances (e.g., slice discovery (Wang et al., 2024)) and on attributing model behaviour on groups of test samples to training data (e.g., (Jain et al., 2024)).
 Reliability of TDA: TDA quantifies the impact of training samples on model errors by predicting changes in model behaviour, serving as an explanation to the user.
 For the explanations to be informative, the attribution must be reliable.
-It is essential to understand how overall model behaviour will be affected as mentioned by our participants (Theme area: Reliability of explanations).
+It is essential to understand how overall model behaviour will be affected, as mentioned by our participants.
 We recommend future research to extend existing work on the fragility, reliability, and stability of TDA (Basu et al., 2021; Nguyen et al., 2023; Epifano et al., 2023; Bae et al., 2022) to address these needs.
@@
 Providing feature attribution explanations with context, such as relevant training data identified through TDA, could help in checking whether similar features in training and test data might mislead the model.
+5.3. A framework for needs-based research
+In this work, we demonstrate a three-step process to identify potentially overlooked research directions and inform needs-based research for socially situated technology like machine learning model explanations:
+Understanding the user: The first step aims to identify the potential user of a technology: Examining who may have an interest in the technology under consideration of their agency, as well as what purpose the technology serves.
+In this work, we consider XAI technology.
+Hence, potential users interact with machine learning systems and have an interest in understanding them.
+Often in XAI, the users of ML technology in high-risk application areas are selected as the main audience due to legal requirements of transparency by the EU AI Act (Parliament, 2023) for example, but interested users could extend beyond this group to any interested users.
+Since we specifically focus on TDA explanations, we were able to narrow the user group to model developers who have agency over the training data.
+Understanding their needs: The second step aims to understand the needs of users through qualitative methods like semi-structured interviews with thematic coding.
+By engaging with users, we can understand the application context of the technology, including existing workflows that the technology needs to fit in and existing challenges that the technology could potentially help with.
+Especially in the case of novel technology that is not yet used, it is important to identify potential use cases to understand what needs exist.
+In this work, we were able to analyse common use cases of TDA explanations in model development from interviews with machine learning developers (§ 4.1) and derive the design space of TDA explanations (§ 4.3).
+Validate their needs: The third step aims to validate the needs found in the previous step through a prototype or scenario-based study, so that users can experience the technology.
+This step enables analysis that is rooted in real interaction, and shall provide deeper insights and validation of the user needs.
+In this work, we designed a mock-up interface for the scenario-based interactive user study and were able to deepen our understanding of what kind of information is needed from TDA explanations (§ 4.4).
+We were able to clearly identify preferences for group attribution, as well as currently understudied research directions, such as the need for explanations of model behaviour that point to missing data.
+This process is intended to build a bridge between user research and methods research, as it allows for potential users to understand the capabilities of novel methods, and results in open questions and example use cases for method researchers.
+While this paper applies the framework to XAI, in particular TDA explanations, it is a general approach that could be applied to any technology intended for use in a sociotechnical context.
 5.4. Limitations
@@
 Second, the relatively small sample size in our studies limits the robustness of the quantitative analyses, so the results should be interpreted with caution and lead us to focus more strongly on the qualitative analysis.
-Third, we simulated the use of TDA explanations in a model debugging interface which are lab-like conditions for the survey study.
+Third, we simulated the use of TDA explanations in a model debugging interface, which are lab-like conditions for the survey study.
 Hence, the needfinding analysis is not rooted in the real workflows and is bound to our study conditions, including the interface design.
+Yet, we remark that this third limitation is somewhat systematic, as TDA explanations are not widely adopted in real workflows yet.
+At the time of the studies, our participants were not familiar with TDA explanations at all, and the first libraries for data attribution were only recently published (Bareeva et al., 2024; Deng et al., 2024).
+We therefore believe that future studies will overcome this limitation as more real-world users engage with the TDA pipeline in their work.
 6. Conclusion
@@
 We illustrate this approach with the emerging subfield of training data attribution (TDA) and present a needfinding study with AI practitioners.
-The study consists of two stages: First, we conduct semi-structured interviews with professionals who already work with AI systems (N=10) to gain an impression of common use cases and challenges.
-We find that TDA could especially benefit model developers, as their workflows are strongly data-centric.
+The study consists of two stages: First, we conduct semi-structured interviews with machine learning developers in high-risk domain areas (N=6) to gain an understanding of common use cases and challenges.
 We derive a design space of TDA explanatory information that represents what relevant data for model debugging could look like.
-Second, we develop an interactive interface to extract user preferences in the design space and integrate it into a survey study (N=31) targeted at model developers.
-The mixed-methods analysis shows that the explanatory information needed from TDA for model debugging is highly dependent not only on the use case but also on the user and their intuitions.
-Finally, we highlight several research directions in user-focused TDA that are largely overlooked in the current landscape.
+Second, we develop an interactive interface to explore developer preferences in the design space through a scenario-based interactive user study (N=31) targeted at model developers.
+The second study allows us to validate and expand the needfinding analysis.
+The mixed-methods analysis shows that the explanatory information needed from TDA for model debugging is highly dependent not only on the use case but also on the developer and their mental model and intuitions.
+Finally, we establish our study approach as a framework for needs-based research and highlight several research directions in user-focused TDA that are largely overlooked in the current landscape.
 Acknowledgements.
@@
 Advances in neural information processing systems 31 (2018).
+- Agarwal et al.
+(2017) Naman Agarwal, Brian Bullins, and Elad Hazan.
+2017.
+Second-order stochastic optimization for machine learning in linear time.
+Journal of Machine Learning Research 18, 116 (2017), 1–40.
 - Akyürek et al.
@@
 IEEE, 81–88.
+- Arnoldi (1951) Walter Edwin Arnoldi. 1951. The principle of minimized iterations in the solution of the matrix eigenvalue problem. Quarterly of applied mathematics 9, 1 (1951), 17–29.
 - Bae et al. (2024) Juhan Bae, Wu Lin, Jonathan Lorraine, and Roger Grosse. 2024. Training Data Attribution via Approximate Unrolled Differentation. arXiv preprint arXiv:2405.12186 (2024).
@@
 https://proceedings.neurips.cc/paper_files/paper/2022/file/7234e0c36fdbcb23e7bd56b68838999b-Paper-Conference.pdf
+- Bareeva et al.
+(2024) Dilyara Bareeva, Galip Ümit Yolcu, Anna Hedström, Niklas Schmolenski, Thomas Wiegand, Wojciech Samek, and Sebastian Lapuschkin.
+2024.
+Quanda: An Interpretability Toolkit for Training Data Attribution Evaluation and Beyond.
+arXiv:2410.07158 [cs.LG] https://arxiv.org/abs/2410.07158
 - Basu et al.
@@
 Association for Computing Machinery, New York, NY, USA, 1–7.
-https://doi.org/10.1145/3334480.3383047
+doi:10.1145/3334480.3383047
 - Brunet et al.
@@
 PMLR, 803–811.
+- Cabrera et al.
+(2023) Ángel Alexander Cabrera, Marco Tulio Ribeiro, Bongshin Lee, Robert Deline, Adam Perer, and Steven M.
+Drucker.
+2023.
+What Did My AI Learn?
+How Data Scientists Make Sense of Model Behavior.
+ACM Trans.
+Comput.-Hum.
+Interact.
+30, 1, Article 1 (March 2023), 27 pages.
+doi:10.1145/3542921
 - Carroll (1995) John M. Carroll (Ed.). 1995. Scenario-based design: envisioning work and technology in system development. John Wiley & Sons, Inc., USA.
@@
 arXiv preprint arXiv:2405.13954 (2024).
+- Deng et al.
+(2024) Junwei Deng, Ting-Wei Li, Shiyuan Zhang, Shixuan Liu, Yijun Pan, Hao Huang, Xinhe Wang, Pingbang Hu, Xingjian Zhang, and Jiaqi Ma.
+2024.
+dattri: A Library for Efficient Data Attribution.
+Advances in Neural Information Processing Systems 37 (2024), 136763–136781.
 - Doshi-Velez and Kim (2017) Finale Doshi-Velez and Been Kim. 2017. Towards a rigorous science of interpretable machine learning. arXiv preprint arXiv:1702.08608 (2017).
@@
 Surveys 55, 9 (2023), 1–33.
-- Edwards (2005) Anthony WF Edwards. 2005. RA Fischer, statistical methods for research workers, (1925). In Landmark writings in western mathematics 1640-1940. Elsevier, 856–870.
 - Ehsan et al.
@@
 Association for Computing Machinery, New York, NY, USA, Article 316, 32 pages.
-https://doi.org/10.1145/3613904.3642474
+doi:10.1145/3613904.3642474
 - Ehsan and Riedl (2020) Upol Ehsan and Mark O.
@@
 Springer International Publishing, Cham, 449–466.
+- Engstrom et al.
+(2025) Logan Engstrom, Andrew Ilyas, Benjamin Chen, Axel Feldmann, William Moses, and Aleksander Madry.
+2025.
+Optimizing ml training with metagradient descent.
+arXiv preprint arXiv:2503.13751 (2025).
 - Epifano et al.
@@
 - Fok and Weld (2023) Raymond Fok and Daniel S Weld. 2023. In search of verifiability: Explanations rarely enable complementary performance in AI-advised decision making. AI Magazine (2023).
+- Ghorbani and Zou (2019) Amirata Ghorbani and James Zou. 2019. Data shapley: Equitable valuation of data for machine learning. In International conference on machine learning. PMLR, 2242–2251.
 - Grosse et al.
@@
 SAGE Publications, Ltd, 55 City Road.
-https://doi.org/10.4135/9781506374680
+doi:10.4135/9781506374680
 - Guidotti et al.
@@
 51, 5, Article 93 (2018), 42 pages.
-https://doi.org/10.1145/3236009
+doi:10.1145/3236009
 - Guo et al.
@@
 Association for Computational Linguistics, Online and Punta Cana, Dominican Republic, 10333–10350.
-https://doi.org/10.18653/v1/2021.emnlp-main.808
+doi:10.18653/v1/2021.emnlp-main.808
 - Hadash et al.
@@
 Association for Computing Machinery, New York, NY, USA, Article 487, 9 pages.
-https://doi.org/10.1145/3491102.3517650
+doi:10.1145/3491102.3517650
 - Hammoudeh and Lowd (2024) Zayd Hammoudeh and Daniel Lowd. 2024. Training data influence analysis and estimation: A survey. Machine Learning 113, 5 (2024), 2351–2403.
-- Hampel (1974) Frank R. Hampel. 1974. The Influence Curve and its Role in Robust Estimation. J. Amer. Statist. Assoc. 69, 346 (1974), 383–393. https://doi.org/10.1080/01621459.1974.10482962
+- Hampel (1974) Frank R. Hampel. 1974. The Influence Curve and its Role in Robust Estimation. J. Amer. Statist. Assoc. 69, 346 (1974), 383–393. doi:10.1080/01621459.1974.10482962
+- Hara et al. (2019) Satoshi Hara, Atsushi Nitanda, and Takanori Maehara. 2019. Data cleansing for models trained with SGD. Advances in Neural Information Processing Systems 32 (2019).
 - He et al.
@@
 - Hofmann (1994) Hans Hofmann. 1994. Statlog (German Credit Data). UCI Machine Learning Repository. DOI: https://doi.org/10.24432/C5NC77.
+- Hohman et al.
+(2020) Fred Hohman, Kanit Wongsuphasawat, Mary Beth Kery, and Kayur Patel.
+2020.
+Understanding and Visualizing Data Iteration in Machine Learning.
+In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20).
+Association for Computing Machinery, New York, NY, USA, 1–13.
+doi:10.1145/3313831.3376177
+- Hong et al.
+(2020) Sungsoo Ray Hong, Jessica Hullman, and Enrico Bertini.
+2020.
+Human Factors in Model Interpretability: Industry Practices, Challenges, and Needs.
+Proc.
+ACM Hum.-Comput.
+Interact.
+4, CSCW1, Article 68 (May 2020), 26 pages.
+doi:10.1145/3392878
+- Hu et al.
+(2024) Yuzheng Hu, Pingbang Hu, Han Zhao, and Jiaqi Ma.
+2024.
+Most Influential Subset Selection: Challenges, Promises, and Beyond.
+In Advances in Neural Information Processing Systems, A.
+Globerson, L.
+Mackey, D.
+Belgrave, A.
+Fan, U.
+Paquet, J.
+Tomczak, and C.
+Zhang (Eds.), Vol.
+37.
+Curran Associates, Inc., 119778–119810.
+https://proceedings.neurips.cc/paper_files/paper/2024/file/d8684e49752e06ac5e4b554b60ad212a-Paper-Conference.pdf
+- Ilyas and Engstrom (2025) Andrew Ilyas and Logan Engstrom. 2025. MAGIC: Near-Optimal Data Attribution for Deep Learning. arXiv preprint arXiv:2504.16430 (2025).
 - Ilyas et al.
@@
 Association for Computing Machinery, New York, NY, USA, 1482–1485.
-https://doi.org/10.1145/3589335.3641297
+doi:10.1145/3589335.3641297
 - Juneja et al.
@@
 Association for Computing Machinery, New York, NY, USA, Article 841, 17 pages.
-https://doi.org/10.1145/3613904.3642059
+doi:10.1145/3613904.3642059
 - Kaur et al.
@@
 Association for Computing Machinery, New York, NY, USA, 1–14.
-https://doi.org/10.1145/3313831.3376219
+doi:10.1145/3313831.3376219
 - Kim et al.
@@
 Association for Computing Machinery, New York, NY, USA, Article 250, 17 pages.
-https://doi.org/10.1145/3544548.3581001
+doi:10.1145/3544548.3581001
 - Koh and Liang (2017) Pang Wei Koh and Percy Liang.
@@
 Advances in neural information processing systems 32 (2019).
+- Krippendorff (2011) Klaus Krippendorff. 2011. Computing Krippendorff’s alpha-reliability.
 - Kwon et al.
@@
 arXiv preprint arXiv:2202.01875 (2022).
+- Li et al.
+(2024) Wenjie Li, Jiawei Li, Christian Schroeder de Witt, Ameya Prabhu, and Amartya Sanyal.
+2024.
+Delta-Influence: Unlearning Poisons via Influence Functions.
+arXiv preprint arXiv:2411.13731 (2024).
 - Liao et al.
@@
 Association for Computing Machinery, New York, NY, USA, 1–15.
-https://doi.org/10.1145/3313831.3376590
+doi:10.1145/3313831.3376590
 - Lin et al.
@@
 - Lundberg and Lee (2017) Scott M Lundberg and Su-In Lee. 2017. A unified approach to interpreting model predictions. Advances in neural information processing systems 30 (2017).
+- Marzi et al.
+(2024) Giacomo Marzi, Marco Balzano, and Davide Marchiori.
+2024.
+K-Alpha Calculator–Krippendorff’s Alpha Calculator: A user-friendly tool for computing Krippendorff’s Alpha inter-rater reliability coefficient.
+MethodsX 12 (2024), 102545.
 - Mei et al. (2023) Alex Mei, Michael Saxon, Shiyu Chang, Zachary C Lipton, and William Yang Wang. 2023. Users are the north star for ai transparency. arXiv preprint arXiv:2303.05500 (2023).
@@
 55, 13s, Article 295 (jul 2023), 42 pages.
-https://doi.org/10.1145/3583558
+doi:10.1145/3583558
 - Naveed et al.
@@
 http://papers.neurips.cc/paper/9015-pytorch-an-imperative-style-high-performance-deep-learning-library.pdf
-- Patnaik and Becker (1999) Dev Patnaik and Robert Becker.
-1999.
-Needfinding: The Why and How of Uncovering People’s Needs.
-Design Management Journal (Former Series) 10, 2 (1999), 37–43.
-https://doi.org/10.1111/j.1948-7169.1999.tb00250.x arXiv:https://onlinelibrary.wiley.com/doi/pdf/10.1111/j.1948-7169.1999.tb00250.x
 - Pearson (1900) Karl Pearson.
@@
 Association for Computational Linguistics, Dublin, Ireland, 1934–1946.
-https://doi.org/10.18653/v1/2022.findings-acl.153
+doi:10.18653/v1/2022.findings-acl.153
+- Piorkowski et al.
+(2023) David Piorkowski, Inge Vejsbjerg, Owen Cornec, Elizabeth M.
+Daly, and Öznur Alkan.
+2023.
+AIMEE: An Exploratory Study of How Rules Support AI Developers to Explain and Edit Models.
+Proc.
+ACM Hum.-Comput.
+Interact.
+7, CSCW2, Article 255 (Oct.
+2023), 25 pages.
+doi:10.1145/3610046
 - Pruthi et al.
@@
 46, 4 (nov 2023), 2104–2122.
-https://doi.org/10.1109/TPAMI.2023.3331846
+doi:10.1109/TPAMI.2023.3331846
 - Rutjes et al.
@@
 Association for Computing Machinery, Inc.
+- Sambasivan et al.
+(2021) Nithya Sambasivan, Shivani Kapania, Hannah Highfill, Diana Akrong, Praveen Paritosh, and Lora M Aroyo.
+2021.
+“Everyone wants to do the model work, not the data work”: Data Cascades in High-Stakes AI.
+In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Yokohama¡/city¿, ¡country¿Japan¡/country¿, ¡/conf-loc¿) (CHI ’21).
+Association for Computing Machinery, New York, NY, USA, Article 39, 15 pages.
+doi:10.1145/3411764.3445518
 - Schioppa et al.
@@
 2022), 8179–8186.
-https://doi.org/10.1609/aaai.v36i8.20791
+doi:10.1609/aaai.v36i8.20791
 - Shen et al.
@@
 Association for Computing Machinery, New York, NY, USA, 384–387.
-https://doi.org/10.1145/3584931.3607492
+doi:10.1145/3584931.3607492
 - Simonyan et al.
@@
 https://openreview.net/forum?id=T6m9bNI7C__
-- Turri et al.
-(2024) V.
-Turri, K.
-Morrison, K.
-Robinson, C.
-Abidi, J.
-Forlizzi, A.
-Perer, and R.
-Dzombak.
-2024.
-Transparency in the Wild: Navigating Transparency in a Deployed AI System to Broaden Need-Finding Approaches.
-ACM Conference on Fairness, Accountability, and Transparency (ACM FAccT) (2024).
-- von Thienen et al.
-(2019) Julia P.
-A.
-von Thienen, William J.
-Clancey, and Christoph Meinel.
-2019.
-Theoretical Foundations of Design Thinking.
-Springer International Publishing, Cham, 13–38.
-https://doi.org/10.1007/978-3-319-97082-0_2
 - Wah et al. (2011) C. Wah, S. Branson, P. Welinder, P. Perona, and S. Belongie. 2011. The Caltech-UCSD Birds-200-2011 Dataset. Technical Report CNS-TR-2011-001. California Institute of Technology.
+- Wang et al.
+(2025a) Andrew Wang, Elisa Nguyen, Runshi Yang, Juhan Bae, Sheila A McIlraith, and Roger Grosse.
+2025a.
+Better Training Data Attribution via Better Inverse Hessian-Vector Products.
+arXiv preprint arXiv:2507.14740 (2025).
 - Wang et al.
@@
 - Wang et al.
-(2024a) Fulton Wang, Julius Adebayo, Sarah Tan, Diego Garcia-Olano, and Narine Kokhlikyan.
-2024a.
+(2024) Fulton Wang, Julius Adebayo, Sarah Tan, Diego Garcia-Olano, and Narine Kokhlikyan.
+2024.
 Error discovery by clustering influence embeddings.
@@
 - Wang et al.
-(2024b) Sheng-Yu Wang, Aaron Hertzmann, Alexei A Efros, Jun-Yan Zhu, and Richard Zhang.
-2024b.
-Data Attribution for Text-to-Image Models by Unlearning Synthesized Images.
-arXiv preprint arXiv:2406.09408 (2024).
+(2025b) Jiachen T Wang, Dawn Song, James Zou, Prateek Mittal, and Ruoxi Jia.
+2025b.
+Capturing the Temporal Dependence of Training Data Influence.
+In The Thirteenth International Conference on Learning Representations.
 - Weisstein (2004) Eric W Weisstein. 2004. Bonferroni correction. https://mathworld. wolfram. com/ (2004).
+- Williams (2021) Oyindamola Williams. 2021. Towards human-centred explainable ai: A systematic literature review. Master’s Thesis (2021).
+- Wolf (2019) Christine T Wolf. 2019. Explainability scenarios: towards scenario-based XAI design. In Proceedings of the 24th International Conference on Intelligent User Interfaces. 252–257.
 - Zheng et al. (2023) Xiaosen Zheng, Tianyu Pang, Chao Du, Jing Jiang, and Min Lin. 2023. Intriguing properties of data attribution on diffusion models. arXiv preprint arXiv:2311.00500 (2023).
@@
 Assistance in legal interpretation and application of the law.
-Appendix B Interview study themes
-The thematic coding analysis of the survey study (N=10) yielded 29 themes corresponding to six theme areas and four unexpected themes (§ 4.2). In the following, we further elaborate on these themes.
-B.1. Theme area: Role of ML systems
-The theme area “Role of ML systems” groups three themes that represent what kind of role ML systems take in our participant’s work:
-Work assistant: ML systems act as work assistants as they take workload off the participant.
-Chatbots generally fill the role of a work assistant that “[takes work off of the participant’s hands and makes their work easier]” (P10) and is “available around the clock” (P10).
-Participants use chatbot systems to improve their writing in English (P1, P11), to search for information where previously they would ”[ask] Google” (P1), or to ideate research ideas (P11).
-Moreover, P10 trusts their company-internal chatbot enough to redirect simple employee questions to the chatbot.
-Decision support: As decision support systems, ML systems deliver information that acts as a basis for decisions taken by end-users, e.g.
-diagnosis support (P6).
-This is the main role taken by the ML systems that the developer participants work on (“[We] make predictions about where we should do our [pharmaceutical] studies and our development.” (P8), “try to accommodate the policymakers so that they’re able to decide [where more health services are needed]” (P9)).
-P3 uses ML systems to identify and explain the contributing factors to product issues: “If we can predict it, we can also have an idea what are the factors mostly creating this phenomenon.”
-Automation: In some cases, ML systems are deployed for the automation of some tasks, e.g. autonomous driving (P2) or automated processing of camera recordings (P4).
-B.2. Theme area: Workflow with ML systems
-The theme area “Workflow with ML systems” groups four themes that emerged in the interviews related to standard or often-occuring workflows:
-Collaboration with domain experts: Developers and system users collaborate closely for building and evaluating ML models, where bugs are reported to the developers by the end-users (P2, P3, P6, P4, P9).
-This collaboration is especially strong if the ML system is developed in-house, and less strong if the ML system is a product the company purchased.
-Standard model choices: Developers rarely build their own model architectures but opt for standard ones (“[If] it’s already an established problem, you’re probably not going to do better than an algorithm that’s already been laid out to solve that problem for you.” (P8)).
-Data-focused debugging process: The development work in practice is centred around data.
-Developer participants mention that most often, data artefacts are the reason for bugs in the model (P2, P3, P4, P7, P8, P9).
-Data quality checks: Data quality checks are a set part of the data preprocessing pipeline (P2, P3, P4, P8).
-P2 and P9 reported that they first assess data quality before inspecting the model in debugging.
-This includes looking for any possible data artefacts, especially missing data.
-B.3. Theme area: Role of training data
-The theme area “Role of training data” groups five themes stemming from participant’s responses when talking about the role of training data in the interviews:
-Most important variable in a model: Model developer participants consistently view training data as the most important variable in a model, e.g.
-“[we] […] believe that […] the models can only be as good as the […] data that you feed in.” (P4).
-Hence, the debugging process is focused around data (Theme: Data-focused debugging process).
-Interesting but no capacity: This theme applies to system user participants, in particular P6.
-They mention that they would be curious about training data to understand whether the model learned from high-quality data but “[it] is a luxury that [requires time]”, highlighting the practical constraint of time pressure.
-Not interesting: This theme applies to system user participants.
-They were unaware of which training data may have been used to train the models (P1, P10) and mentioned that this does not matter too much for their work as they cannot change it anyhow.
-Increasing training data as a bug fix: A reoccuring theme in our interviews is that training data is flexible – the training set can be expanded.
-This is often done in practice when encountering a bug, e.g.
-P2 explained that collecting more data is a common way to overcome model shortcomings in autonomous driving.
-Training data artefacts as main bug causes: In line with previous themes highlighting the importance of training data to the development process, participants explicitly mention that most often, training data artefacts are the cause of bad model performance.
-Hence, data artefacts are also the first thing that developers look our for when encountering a bug.
-B.4. Theme area: Challenges in working with ML systems
-The theme area “Challenges in working with ML systems” groups six themes related to the main pain points that participants mentioned
-Trust calibration: Our findings agree with Kim et al.
-(2023): It is unclear how much and when a system can be trusted.
-P1 sometimes finds themselves in a dilemma in which they wish to learn something from the chatbot, but are unable to calibrate their trust in the response due to missing knowledge: “I don’t know everything regarding this topic.
-I [don’t even] know what he’s replying to me.” (P1).
-Validation and Evaluation: P2 mentions that the validation itself is a challenge due to multiple requirements that the ML system should fulfil.
-In practice, the predictive performance of a model matters, along with other objectives such as inference speed, whether the model is bias-free, and more.
-In addition, A data-centric aspect of evaluation being challenging is the difficulty of evaluation when labels are absent.
-Data quality issues: Data quality issues are a major challenge when working with ML systems. In Table 2, we detail the quality issues mentioned by the pariticipants.
-Resource constraints: For developer participants, memory and compute constraints are relevant in the development and evaluation process of their models, especially to P2 and P4 as they work with ML systems on the edge.
-Stochastic nature of ML models: The stochastic nature of ML models poses a specific challenge to deep models in the development process and debugging tasks: “[The] same data set, same model, you train multiple times, you can get [different] results.” (P2).
-It is therefore difficult to understand if a bug was truly fixed or if it is a consequence of randomness.
-Misuse due to lacking know-how: P10 mentions inadequate know-how in ML system usage as a challenge: “the employees often don’t manage to ask the chatbot the right questions”.
-B.5. Theme area: Use of (feature attribution) XAI in practice
-The theme area “Use of (feature attribution) XAI in practice” groups six themes related to how XAI is already being used by the participants.
-XAI is not used: Some system user participants were unaware of XAI since explanations are not a part of the ML tools they use (P1, P10, P11).
-XAI is little used: This theme is related to the system users.
-P6 reports that XAI tools they used in radiology images so far (i.e.
-heatmaps) do not deliver a full answer to the why question, as counterfactual information is missing: “[If] I just get an overall highlight in these basal lung regions and the prediction that is atelectasis, I still don’t know why this is atelectasis but not pleural effusion or consolidation.” Hence, they only use it when the information is useful.
-Model development: For developer participants, XAI is most commonly used as a tool for model development (P2, P3, P8).
-As such, XAI tools (within our participants, these are exclusively feature attribution explanations, i.e., SHAP (Lundberg and Lee, 2017) or visualisation of attention) offer explanations for per-example debugging of e.g.
-wrong predictions (“we work with hands-on examples that we then […] relate the input features that take a poor or a good prediction [to].” - P8).
-Understand real-world phenomena: P3 described the use of XAI as a tool to understand phenomena represented by the ML model: “[Building] the model, the whole purpose is to get some explainability.
-Because […] we know that [a problem is] happening and predicting doesn’t really add value.
-But if we can predict it, we can also have an idea what are the factors mostly creating this phenomenon.”
-Get customer buy-in: P8 states that XAI tools are useful in getting customer buy-in and convincing the customers of the model’s decision suggestion: “Will [the customers] be happy with [high accuracy]?
-Probably.
-Will they be convinced?
-Will they be 100% confident that that’s the way they want to do it?
-Probably not.
-So it’s all about the buy-in that you get, and SHAP definitely helps.”
-Sanity check for model reasoning: One particular case of using XAI for model development is as a sanity check for model reasoning (“If it’s more like an ideation project, then I would say that the explainability analysis [is done] to see […] what is guiding the prediction as well as […] how are my predictions on the test set explained?
-And […] does it align with the actual value that I have?” - P8).
-B.6. Theme area: User perspectives on TDA
-The theme area “User perspectives on TDA” groups four themes that emerged in the participant’s answers when discussing TDA.
-Usefulness of XAI is use-case dependent: Some participants find XAI only useful in certain use-cases and under certain conditions such as time availability, intuitiveness of the explanation, and the possibility of taking action based on the explanation.
-When explanations are used for communication with business departments, they do not deliver useful information when the underlying model is not high-performing yet (P3, P8).
-XAI is useful: While participants voiced certain conditions for the usefulness of XAI, the overall majority agreed that explanations are useful in their work (P1, P2, P3, P9, P11).
-TDA complementary to feature attribution: Participants are overall positive towards TDA explanations.
-They mentioned that TDA could be complementary to feature attribution explanations and provide context, as feature attribution information is rather local (e.g., comparing feature attribution maps between training and test samples that are highly related to each other).
-Different notions of interesting training samples: Within our participants, we find that there are several different notions of what an interesting training sample to a model output could be among the participants.
-We detail the codes in this theme in Table 2 of the main paper.
-Appendix C Data generation for the survey study
+Appendix B Data generation for the scenario-based interactive user study
 The second probe of our study presents an interactive mock-up of a model development suite where participants can customise what kind of data-centric information is shown.
@@
 In the following, we describe the data generation process.
-C.1. Use case: Bird classification app
+B.1. Use case: Bird classification app
 The setting of the image classification use case is about a company that builds a bird classification app.
@@
 The final model achieves a test accuracy of 0.57 which is suitable for the model debugging use cases in this work.
-C.2. Use case: Loan approvement recommendation app
+B.2. Use case: Loan approvement recommendation app
 The setting of the scenario of the tabular data use case is the model development department of a bank that implements a machine learning model to give recommendations for loan approvals of their clients.
@@
 The final model accuracy is , leaving room for improvement and debugging.
-C.3. Model error selection
+B.3. Model error selection
 In the survey probe, the participant is asked to analyse a model error, i.e.
@@
 The test samples are randomised across participants by drawing from a uniform distribution to prevent bias in the analysis stemming from a specific error or test sample.
-C.4. Relevant training sample selection
+B.4. Relevant training sample selection
 In the mock-up interface used in the survey probe, we display the effect of different training data-based actions on the model error at hand and the global model behaviour.
@@
 We compute and show the standard deviation of removing/changing the label/upweighting across the model checkpoints recorded.
-Appendix D Survey study themes
-The thematic coding analysis of the survey study (N=31) yielded 34 themes corresponding to four theme areas and four unexpected themes (§ 4.5). In the following, we further elaborate on these themes.
-D.1. Theme area: Action depends on user’s hypothesis
-The theme area “Action depends on user’s hypothesis” groups 21 themes linked to the reasons for participant’s choices in the action axis.
-We find that the proposed debugging actions of participants are highly dependent on their hypotheses of the model error’s root cause.
-[Hypothesis] Mislabeled data: The model makes an error because of mislabeled data in the training set that is highly related to the erroneous test sample.
-[Hypothesis] Incomplete training: The model makes an error because it has not finished training, or was trained using non-optimal hyperparameters and training procedures.
-[Hypothesis] Spurious correlation: The model makes an error due to a spurious correlation being present in the test image, that the model learnt from the training data.
-[Hypothesis] Similar class features: The model makes an error because the erroneous test sample is similar in features to another class in the dataset.
-[Hypothesis] Low dataset diversity: The model makes an error because it was not able to generalise from training on a dataset with low diversity in the samples (i.e., samples of a class are all similar, not showing the breadth of the class).
-[Hypothesis] Class overlap: The model makes an error because the class of the erroneous test sample is highly overlapping in terms of class-relevant features with another class and distinctive features may not be present in the data.
-[Hypothesis] Dataset imbalance: The model makes an error because it was not able to generalise from an imbalanced dataset (i.e., class imbalance).
-[Hypothesis] Small dataset size: The model makes an error because it was not able to generalise from a small dataset, instead it overfit it.
-[Action] Remove noisy data: To overcome the error, noisy data such as mislabeled data, is removed from the dataset and the model is retrained.
-[Action] Changing label to correct: To overcome the error, incorrectly labelled training data is corrected and the model is retrained.
-[Action] Fix labelling pipeline for future data: To overcome the error, the labelling and annotation pipeline is inspected to find a root cause for labelling errors.
-[Action] Training process engineering: To overcome the error, hyperparameters and training process protocols are adjusted and tuned before retraining.
-[Action] Model architecture engineering: To overcome the error, the model architecture is adjusted (e.g. addition of another layer, choosing a different foundational model).
-[Action] Feature engineering: To overcome the error, the feature extractors of the data preprocessing pipeline are inspected. This theme was mainly found for the tabular data use case.
-[Action] Upweight to focus on specific features: To overcome the error, certain training data samples should be upweighted so that the model learns the features present in these samples better.
-[Action] Feature attribution inspection: To overcome the error, the misleading features must be understood first. Feature attribution could be a useful tool.
-[Action] Manual inspection: To overcome the error, the training data is manually inspected to find any data artefacts.
-Participants mentioned that they already do this as part of their model development pipeline.
-[Action] Collect specific new data: To overcome the error, new data related to the erroneous training sample should be collected and added to the training dataset.
-This can improve dataset diversity and help the model generalise better.
-[Action] Synthesise/augment data: To overcome the error, new data samples should be synthesised or existing data samples should be augmented further, targeting the parts of training data that are underrepresented.
-This theme was mentioned often under the assumption that adding completely new data is not possible.
-[Action] Resample/reweight data: To overcome the error, the existing training dataset should be resampled, potentially with a specific weighting to overcome class imbalances.
-[Action] Collect any new data: To overcome the error, the model needs to learn from more data. Hence, expanding the dataset overall will likely help the model performance.
-D.2. Theme area: Individual metric preferences
-The theme area “Individual metric preferences” groups four themes and provides insight into why certain metrics are preferred.
-[Action/Metric] Label change and probability shows decision boundary: Probabilities are the preferred metric, especially in combination with the label change action, as it gives the user an intuition about the decision boundary between the predicted and ground-truth class.
-[Metric] Probabilities are intuitive and show decision boundaries: Probabilities are the preferred metric because they are an intuitive metric bounded between 0-100.
-Also, they give the user an intuition about the decision boundary between the predicted and ground-truth class.
-[Metric] Ground truth probability is the target: The probability of the ground truth class of the test sample is preferred because this is the target metric to optimise when debugging this erroneous test sample.
-[Metric] Loss is most informative in the multi-class case: The loss is preferred because it entails information across many classes and is a direct indicator of how well the model is performing overall.
-This is particularly interesting in multi-class classification cases.
-D.3. Theme area: Reliability of explanations
-The theme area “Reliability of explanations” groups three themes related to the overarching need for reliability of the attribution scores.
-[Action] Removal shows causal effect: The removal action is preferred because represents the causal effect of the inclusion of a specific training sample.
-This causal effect is assumed to be meaningful and therefore reliable.
-[Number of samples] Group size enough for reliable attribution: A group (i.e., more than one training sample) is preferred for training data attribution because considering the effect of a group is more reliable than of single training samples.
-The group size should be sufficiently large for the attribution score to be meaningful and hence, reliable.
-[Number of samples] Small group to prevent reasoning change: While a group of training samples is preferred, the size of the group should be sufficiently small so that the underlying model stays rather consistent given the proposed action (e.g., if the attribution corresponds to the effect of removing samples, this will change the data distribution which affects the model’s learning, too).
-If the underlying model were to change too much, the attribution scores could be arbitrary and it is not certain whether the model behaviour will stay similar, making the sample-wise debugging process difficult.
-D.4. Theme area: Group attribution preferences
-The theme area “Group attribution preferences” groups two themes that elaborate on desired sizes of training data groups as explanations.
-In contrast to the other axes of action and metric, there is a clear preference towards training data group attribution, emphasising a need for methods that serve this task.
-[Number of samples] Group for an overview of the error: A group of training samples is preferred in the explanation of the model error to get a better overview of the error, as it is expected that the samples of the group with the highest attribution are related to the model error.
-One sample is too little for the user to understand the error and propose an action forward based on the explanation.
-[Number of samples] Balance between variety and conciseness: While a group of training samples is preferred, the group size should be large enough to show a variety of different training samples related to the model error, while being small enough to be concise and not overwhelm the user.
-D.5. Unexpected themes
-In addition to the previous theme areas, we found three additional themes which were unexpected.
-They provide context to the overall analysis, show two predominant mindsets of developers and point to the research direction of grouping erroneous test samples, which has largely been overlooked.
-Focus on the model error at hand: The main focus of the user during the debugging process lies in understanding the model error at hand, and identifying a way to fix the error. This is a local focus.
-Focus on the overall model performance: The main focus of the user during the debugging process is to optimise the overall model performance, i.e., having the overall model accuracy in mind.
-This is a global focus.
-Grouping erroneous test samples: In the debugging process, users seek to understand the reasons behind the model error.
-A single test sample is unlikely to be the only instance of erroneous model behaviour, e.g.
-spurious correlations.
-This theme describes the need for presenting the user with multiple test instances that the model errs on similarly so that the user will have a better understanding of the error which will enable them to fix it.
-Study interface unfamiliar: In the comments, we occasionally found remarks that the participants were unsure about whether they had correctly fulfilled the task, indicating the overall infamiliarity with sample-wise debugging with TDA explanations.
+Appendix C Extended background: Training Data Attribution Methods
+We provide a high-level overview of TDA methods cited in Table 4 of the main paper and how they relate to the attribution types.
+Specifically, we elaborate on influence functions methods (§ C.1), methods that trace the training process (§ C.2), methods that unroll the training process (§ C.3) and group influence methods (§ C.4).
+C.1. Influence Function Methods
+Influence functions (IF), first introduced in robust statistics (Hampel, 1974), are one of the main types of methods in TDA.
+Instead of costly retraining without a training sample to quantify ’s influence on the model’s behaviour via Equation 1, (IF) approximate this quantity by:
+where uses a Newton-like step to estimate the effect of a small perturbation on the model’s parameters, where the perturbation is the exclusion of .
+For a full derivation, we refer to Koh and Liang (2017) or Bae et al.
+(2022).
+Intuitively, influence functions approximate the parameter landscape around the parameters , and compute the update step taken in the direction of removing , to estimate its effect on the loss at .
+IF require approximations to be applicable to deep models.
+Since equation 3 requires the inversion of the parameter Hessian (), it also requires the objective function (i.e., the loss) to be twice-differentiable and the Hessian to be positive definite, to ensure that exists.
+However, these conditions generally do not hold for deep models.
+Moreover, the Hessian is a large and costly matrix to invert, so IF algorithms usually do not explicitly estimate but implicitly through a matrix-vector product.
+They estimate the damped inverse Hessian Vector Product (iHVP) , where is the damping parameter and is the identity matrix.
+Damping improves the numerical stability of the iHVP estimation.
+In the following, we use the terms damped iHVP and iHVP interchangeably.
+IF methods generally differ in iHVP approximation.
+Different TDA methods propose different approximations of the iHVP: Koh and Liang (2017), who first presented IFs for deep models, estimate the iHVP using an iterative estimator called Linear time Stochastic Second-Order Algorithm (LiSSA) (Agarwal et al., 2017).
+Guo et al.
+(2021) also use LiSSA in their algorithm, but propose an additional speed up for the task of identifying the most influential training points for a test point by first retrieving candidate with a k-nearest neighbour search in the model’s embedding space.
+Wang et al.
+(2025a) improve the iHVP and therefore influence estimation by using the Eigenvalue-corrected Kronecker-Factored Approximate Curvature (EKFAC) (Grosse et al., 2023) as a preconditioning matrix on the iterative estimator.
+Alternatively, Schioppa et al.
+(2022) propose to speed up the approximation of the iHVP by projecting the computation into a lower-dimensional space spanned by the dominant eigenvectors of with Arnoldi iterations (Arnoldi, 1951).
+Also Park et al.
+(2023) and Choe et al.
+(2024) leverage a projection into lower dimensions to speed up the iHVP estimation, where the former uses random projections while the latter makes use of gradient structures.
+Grosse et al.
+(2023) propose to estimate the utilising Eigenvalue-corrected Kronecker Factored Approximate Curvature (EKFAC), allowing for a faster computation of the influence function since it is not an iterative method.
+Influence function research is active and has been gaining traction in recent years.
+The main focus lies on finding an accurate estimation of the loss difference on a single test sample when a single training sample is removed.
+C.2. Methods That Trace The Training Process
+In 2020, Pruthi et al.
+(2020) presented a different way to compute training data attribution that traces a training sample’s contribution to the model performance on a test sample throughout training.
+The objective definition of TDA is slightly different from the change after leave-one-out retraining (Eq.
+1):
+TDA is hence defined as the sum of the loss difference on at each training step with across the full training.
+In practice, computing this idealised notion of TDA is also prohibitive across all training samples, so Pruthi et al.
+(2020) propose to approximate equation 4 with a sum of gradient dot products across model checkpoints recorded during training:
+where is the learning rate at checkpoint .
+We note that in principle, each dot product corresponds to the formulation of influence functions (Eq.
+3) where the inverse Hessian is approximated by the identity matrix.
+TracIn (Pruthi et al., 2020) offers an easy-to-compute TDA method and objective.
+It is focused on the relationship between a single training and test sample, measured on the loss changes across training, whenever the training sample is seen during training.
+C.3. Methods That Unroll The Training Process
+In a similar yet different spirit to tracing the training process as in TracIn (Pruthi et al., 2020), another line of approaches to computing TDA are methods that are based on unrolling the training process (Hara et al., 2019; Bae et al., 2024; Wang et al., 2025b; Ilyas and Engstrom, 2025), sometimes also referred to as dynamic TDA methods (Hammoudeh and Lowd, 2024).
+Unrolling the training process results in a computational graph which can be traced backwards to compute attribution.
+Hence, these methods also consider the training procedure, but differentiate through it to approximate the change in model parameters with leave-one-out retraining (cf.
+equation 1) (Hara et al., 2019).
+Since differentiating through each training step is computationally expensive, methods often group several steps in training segments and base their approximation on a sum of influence function-like computations per segment.
+Similar to influence functions, Hara et al.
+(2019) leverage the implicit function theorem to compute the Hessian vector product, and Bae et al.
+(2024) make use of EKFAC approximations of the Hessian (Grosse et al., 2023) for the influence-like computation.
+Ilyas and Engstrom (2025) propose a different approach using metagradients (Engstrom et al., 2025) to explicitly compute the influence function for single-model training data attribution, fixing sources of training process stochasticity such as batch ordering and model initialisation.
+While this approach achieves near-optimal performance, it is very costly.
+As unrolling-based methods, like influence functions, aim at capturing the effect of leave-one-out retraining, they provide methods to consider deep model training specificities (e.g., learning rate schedules) and generally target the attribution of model behaviour measured in loss on a single test sample to a single training sample when it is removed from training.
+C.4. Group Attribution Methods
+TDA methods generally address the individual attribution of the model behaviour to a training sample .
+Another line of work studies methods for quantifying the group attribution of to a set of training samples .
+The TDA objective is analogous to individual attribution, and is generally quantified by the loss difference after leave-some-out retraining:
+Like individual attribution, computing equation 6 directly is computationally prohibitive when the objective is to identify the group of training samples with the largest attribution.
+Hence, group TDA work studies methods to approximate this quantity efficiently.
+Group attribution assuming linearity
+An approach that limits the computational cost makes the core assumption that group attribution is linear: In other words, the attribution score of a group of training samples is defined as the sum of the individual attribution scores of the training samples .
+Koh et al.
+(2019) found that interaction effects between training samples of a group can be neglected, and that the linearity assumption is sufficiently accurate for group TDA estimation.
+As a result, all individual TDA methods, like influence functions, can also be used to estimate group TDA.
+Moreover, the common evaluation metric linear datamodeling score (Park et al., 2023) used to evaluate the performance of TDA methods is based on the assumption of linearity.
+Group attribution considering interaction effects
+Group TDA methods that assume linearity have the same computational cost as individual TDA methods, but ignore any potential interaction effects between training samples of a group.
+While Koh et al.
+(2019) found the interaction effects to be minimal and therefore negligible, a different work by Basu et al.
+(2020) improves the accuracy of influence functions in capturing equation 6 by extending influence functions with a second-order term to capture interaction effects.
+Shapley value-based approaches like Data Shapley (Ghorbani and Zou, 2019; Lin et al., 2024) also consider interaction effects by default.
+Hu et al.
+(2024) propose an adaptive algorithm that iterates over the training dataset and gradually identifies the most influential subset for a given model behaviour.